### v1

# Character theoretic techniques for nonabelian partial difference sets

Abstract.
A $(v,k,\lambda,\mu)$-partial difference set (PDS) is a subset $D$ of size $k$ of a group $G$ of order $v$ such that every nonidentity element $g$ of $G$ can be expressed in either $\lambda$ or $\mu$ different ways as a product $xy^{-1}$, $x,y\in D$, depending on whether or not $g$ is in $D$. If $D$ is inverse closed and $1\notin D$, then the Cayley graph ${\rm Cay}(G,D)$ is a $(v,k,\lambda,\mu)$-strongly regular graph (SRG). PDSs have been studied extensively over the years, especially in abelian groups, where techniques from character theory have proven to be particularly effective. Recently, there has been considerable interest in studying PDSs in nonabelian groups, and the purpose of this paper is develop character theoretic techniques that apply in the nonabelian setting. We prove that analogues of character theoretic results of Ott [25] about generalized quadrangles of order $s$ also hold in the general PDS setting, and we are able to use these techniques to compute the intersection of a putative PDS with the conjugacy classes of the parent group in many instances. With these techniques, we are able to prove the nonexistence of PDSs in numerous instances and provide severe restrictions in cases when such PDSs may still exist. Furthermore, we are able to use these techniques constructively, computing several examples of PDSs in nonabelian groups not previously recognized in the literature, including an infinite family of genuinely nonabelian PDSs associated to the block-regular Steiner triple systems originally studied by Clapham [6] and related infinite families of genuinely nonabelian PDSs associated to the block-regular Steiner 222-designs first studied by Wilson [33].
1. Introduction
If $G$ is a group, then we say that $D\subset G$ is a $(v,k,\lambda,\mu)$-partial difference set (PDS) if $G$ has order $v$, $D$ has size $k$, and if the product $gh^{-1}$ for all $g,h\in D$ with $g\neq h$ represents all nonidentity elements in $D$ exactly $\lambda$ times and all nonidentity elements in $G-D$ exactly $\mu$ times. We refer to the tuple $(v,k,\lambda,\mu)$ as the parameters of $D$. In addition, we call a PDS regular if $D$ is closed under taking inverses (in which case we write $D=D^{(-1)}$) and $1\not\in D$. This definition is not particularly restrictive, for if $D$ is a PDS with $1\in D$, then $D-\{1\}$ is a PDS as well, and if $\lambda\neq\mu$, then $D=D^{(-1)}$. If we have $\lambda=\mu$ then $D$ is called a difference set, and we say that $D$ is a $(v,k,\lambda)$-difference set (DS). We additionally call $D$ a reversible $(v,k,\lambda)$-DS if $D=D^{(-1)}$.
PDSs have been the subject of great interest over the years (see, e.g., [24]), especially in abelian groups. In the abelian group setting, character theoretic methods have proven to be particularly effective. For example, if we extend an irreducible character $\chi$ for an abelian group $G$ in the natural way to the group algebra $\mathbb{C}[G]$ and identify an inverse-closed subset $D\subset G$ with $\sum_{d\in D}d\in\mathbb{C}[G]$, then $D$ is a $(v,k,\lambda,\mu)$-PDS if and only if
(1) $\chi(D)=\begin{cases}k,\text{ if $\chi$ is the trivial character,}\\ \frac{(\lambda-\mu)\pm\sqrt{\Delta}}{2},\text{ otherwise,}\end{cases}$
[…]
There has been interest recently in PDSs in nonabelian groups [9, 13, 31, 32]; for a recent survey of known results related to PDSs in nonabelian groups, see [27]. However, most character theoretic results that apply in the abelian case do not hold in the nonabelian case: for example, while linear characters evaluate to the values listed in Equation (1) above, general irreducible characters do not necessarily need to evaluate to one of these values on a PDS (see Lemma 3.2). Perhaps more importantly, the characters evaluating to one of a few specific values on an inverse-closed subset is not a sufficient condition for the subset to be a PDS in the nonabelian case (only a necessary condition); see Remark 4.8.
Up until now, the only restrictions that apply in the nonabelian setting are those of [32], which were inspired by the work of Yoshiara on groups acting regularly on the point set of a generalized quadrangle (GQ) [34]. The main method in each of [32, 34] was to constrain the value $\Phi(g)$, which counts the number of vertices (resp., points) moved to adjacent vertices (resp., points) by an automorphism $g$ in the automorphism group $G$ of the SRG (resp., GQ). Yoshiara was able to utilize Benson’s Lemma [1, Lemma 4.3], which gave modular constraints on $\Phi(g)$ for GQs. De Winter, Kamischke, and Wang [10] later proved a generalization of Benson’s Lemma for arbitrary SRGs, and Swartz and Tauscheck [32] used the generalization of De Winter, Kamischke, and Wang to push Yoshiara’s result into an analogous modular restriction on SRGs with a regular automorphism group.
Recently, Ott [25] studied $\Phi$ – which is indeed a class function on $G$ – from a character-theoretic perspective, and Ott used this to prove that GQs of order $s$ ($s+1$ points incident with each line, $s+1$ lines incident with each point), $s$ even, cannot have a group acting regularly on its point set. As it turns out, the ideas Ott used are applicable in situations far beyond GQs (and reversible difference sets); in fact, one could argue that the PDS setting is in fact where these techniques are most effective.
The purpose of this paper is to extend Ott’s results to PDSs in nonabelian groups. (Indeed, just as techniques such as Equation (1) are only truly effective in the abelian setting, the results outlined here are most effective in certain nonabelian settings.) These techniques have enabled us to rule out the existence of a PDS inside a range of groups and to provide severe restrictions on certain families of PDSs. We also develop new techniques for discovering and constructing new examples of PDSs inside of nonabelian groups. For a given group $G$ of order $v$, if there exists a parameter set $(v,k,\lambda,\mu)$ which could potentially correspond to a PDS inside $G$, then these methods use character theory to give information about the size of the intersection of this potential PDS with every conjugacy class of $G$.
We have used these methods to show that if a $(v,k,\lambda,\mu)$-PDS exists in a group $G$, where there is a minimal nontrivial normal subgroup $N$ of $G$ such that $G/N$ is abelian and $\gcd(|G/N|,\sqrt{\Delta})=1$, then the size of the intersection of the PDS with the cosets of $N$ are known, providing severe restrictions on such PDSs. We have also used our methods to construct several PDSs. For $p$ a prime, we demonstrate the existence of an infinite family of PDSs in nonabelian groups related to Steiner triple systems with parameters $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$, where $p^{d}$ is greater than 999 and congruent to $7\pmod{12}$, as well as Steiner 222-designs ${\mathrm{S}}(2,k,v)$ with parameters $(p^{d}(p^{d}-1)/(k(k-1)),k(p^{d}-k)/(k-1),(p^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$ for $p^{d}$ sufficiently large and $p^{d}\equiv k(k-1)+1\pmod{2k(k-1)}$. (These Steiner triple systems were shown to be block-regular by Clapham [6], and the Steiner 222-designs were shown to be block regular by Wilson [33]. It is immediate that these will be PDSs; however, PDS families in Steiner 222-designs do not seem to have been recognized thus far in the literature beyond a recent mention in [28].)
This paper is organized as follows. In Section 2, we outline preliminary results related to PDSs, SRGs, and character theory over local rings necessary for later sections. Section 3 is dedicated to generalizing the results of Ott to the PDS setting. In Section 4, we introduce computational methods for the case when one of $\sqrt{\Delta}$, $v$ has factors which are relatively prime to the others that are used both to rule out the existence of numerous PDSs and also discover several new examples in nonabelian groups; we also provide an example demonstrating how the techniques from this paper can be used when $\sqrt{\Delta}$ and $v$ have the same primes as factors (such as when $\Delta=v$). Finally, we include appendices in which we list parameters ruled out either theoretically or computationally (Appendices A, B) and parameters for which there exists a potentially valid $\Phi$ function (Appendix C).
2. Preliminaries
2.1. Partial difference sets
For our purposes, $G$ will always be a finite group, and $D$ will always denote a regular PDS in $G$. In this paper, we study the structure of $G$ and $D$ utilizing character theory. To do so, we often identify $D$ with the element $\sum_{d\in D}d$ in the group ring $\mathbb{C}[G]$, and we occasionally abuse notation by writing $D=\sum_{d\in D}d$. This alternative description of a PDS is useful due to the following result (see [24, Theorem 1.3]).
Lemma 2.1.
Let $D$ be a regular PDS in a group $G$. Then, in the group ring $\mathbb{C}[G]$,
$D^{2}=k1+\lambda D+\mu(G-D-1).$
In fact this condition is sufficient for $D$ to be a regular PDS.
2.2. Strongly regular graphs
As we have mentioned, the notion of a strongly regular graph is closely related to that of a PDS, so we now take the opportunity to formally introduce SRGs. For further reference, see [16, Chapter 10]. We call a graph $\Gamma$ a $(v,k,\lambda,\mu)$-SRG on $v$ vertices if $\Gamma$ is $k$-regular and if, for any pair of vertices $u,v$ in $\Gamma$, either $u,v$ have $\lambda$ neighbors in common when $uv$ is an edge, or they have $\mu$ neighbors in common when $uv$ is not an edge. Now, if $\Gamma$ is a $(v,k,\lambda,\mu)$-SRG with an adjacency matrix $A$, then $A$ has eigenvalues $k$,
$\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta}),\;\;\;\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$
[…]
$m_{i}=\frac{1}{2}\left((v-1)+(-1)^{i}\frac{2k+(v-1)(\lambda-\mu)}{\sqrt{\Delta}}\right)$
for $i=1,2$ (see [16, Chapter 10]). As we will see, the values $\theta_{1},\theta_{2}$, and $\Delta$ are of particular importance in the study of PDSs.
We set aside now two important relations for the eigenvalues, which we will use freely throughout the paper. Their proof follows immediately from the definition for $\theta_{1},\theta_{2}$ provided above.
Lemma 2.2.
Let $\Gamma$ be a $(v,k,\lambda,\mu)$-SRG with eigenvalues $k,\theta_{1},\theta_{2}$ as before. Then,
  • (i)
    $\theta_{1}\theta_{2}=\mu-k$, and
  • (ii)
    $\lambda-\mu=\theta_{1}+\theta_{2}$.
[…]
Furthermore, we say that the $(v,k,\lambda,\mu)$-SRG is imprimitive if either $\mu=0$ or $\mu=k$ and is primitive otherwise. As noted in [4, Section 1.1.3], imprimitive SRGs are either the union of complete graphs of the same size or are complete multipartite graphs. We will only consider parameter sets that correspond to primitive strongly regular graphs, and so we will always assume that $0<\mu<k$.
Finally, we remark that the existence of a $(v,k,\lambda,\mu)$-SRG is equivalent to the existence of its complement, which is itself a $(v,v-k-1,v-2k+\mu-2,v-2k+\lambda)$-SRG, and, indeed, the complement of a regular $(v,k,\lambda,\mu)$-PDS in $G-\{1\}$ is itself a $(v,v-k-1,v-2k+\mu-2,v-2k+\lambda)$-PDS in $G$. Note that one can rule out the existence of a PDS by ruling out the existence of its complement.
2.3. Some families of SRGs
We have already mentioned the particular class of finite geometries known as generalized quadrangles. We now provide some definitions and notation for generalized quadrangles, as we will need them in Proposition 3.18. A generalized quadrangle of order $(s,t)$ is an incidence geometry of points and lines satisfying the following conditions:
  • (i)
    Each point is incident with $t+1$ lines, and two points are both incident with at most one line.
  • (ii)
    Each line is incident to $s+1$ points, and any two lines are both incident to at most one point.
  • (iii)
    If $P$ is a point and $\ell$ a line not incident with $P$, there is exactly one point on $\ell$ collinear with $P$.
[…]
One particularly important family of strongly regular graphs are those related to Steiner 222-designs, which we use in Proposition 4.9, Theorem 4.11, Proposition 4.12, and Proposition 4.13. A Steiner 222-design, ${\mathrm{S}}(2,k,v)$, is a $v$-element set, along with a collection of $k$-subsets of this set, called blocks, such that any two elements are contained in exactly one block. The families of Steiner 222-designs in Proposition 4.9, Theorem 4.11, Propostion 4.12, and Proposition 4.13 are all algebraic constructions using the elements of an abelian group $G$. In these cases, the Steiner 222-design is represented by a distinguished set $\mathscr{B}$ of blocks, known as base blocks. The remaining blocks take the form $g+B$ for $B\in\mathscr{B}$ and $g\in G$.
A Steiner 222-design has a natural representation as a strongly regular graph if one identifies the blocks of the Steiner 222-design with the vertices of the graph such that two vertices are adjacent if the intersection of their corresponding blocks has one element. If $v$ is the number of elements in the Steiner 222-design, then the corresponding SRG has parameters
$\left(\frac{v(v-1)}{k(k-1)},\frac{k(v-k)}{k-1},\frac{v-1}{k-1}+(k-1)^{2}-2,k^{2}\right).$
We will speak often of one particular class of Steiner 222-designs, known as Steiner triple systems, which occur when $k=3$. For further information on Steiner triple systems, Steiner 222-designs, and Steiner systems more generally, see [7, Section I.4].
2.4. Character theory and representations over local rings
Throughout this paper, we assume that the reader is familiar with fundamental results from representation theory and character theory. For reference, the reader may consult [20, Chapter 2]. In addition, we also utilize local rings and local representations. We place here the main facts.
Local rings are constructed for particular primes to describe the “local behavior” around this prime. They are used in a variety of different settings, though for our purposes we only consider them in the field $\mathbb{Q}(\zeta)$ for $\zeta$ an $n$th root of unity.
Definition 2.3.
For a specified prime $p$, and prime ideal $\mathfrak{p}_{0}\subseteq\mathbb{Z}[\zeta]$ such that $p\in\mathfrak{p}_{0}$, define the ring $\mathfrak{R}$ by
=$\mathfrak{R}=\biggl{\{}\frac{x}{y}:x,y\in\mathbb{Z}[\zeta],y\not\in\mathfrak{p}_{0}\biggr{\}}.$
$\mathfrak{R}$ is known as the local ring at $\mathfrak{p}_{0}$.
The local ring $\mathfrak{R}$ has the following important properties, which are proven in [12, 4.1]. For a more general explanation of the subject for non-experts, see [11, 16].
Lemma 2.4.
Let $p$ be prime, and let $\mathfrak{R}$ be the local ring at $\mathfrak{p}_{0}$ such that $p\in\mathfrak{p}_{0}$, and let $G$ be an order $n$ group, then
  • (i)
    any $\mathbb{Q}(\zeta)$-representation of $G$ is similar to a $\mathfrak{R}$-representation of $G$,
  • (ii)
    $\mathfrak{R}$ is a PID, and
  • (iii)
    $\mathfrak{R}$ has a unique maximal ideal $\mathfrak{p}$ containing $p$.
We also need the following theorem of Brauer, see [20, Theorem 10.3].
Theorem 2.5.
Let $G$ have exponent $n$, and let $\zeta$ be an $n^{\text{th}}$ root of unity. Then every $\mathbb{C}$-character of $G$ is afforded by a $\mathbb{Q}(\zeta)$-representation of $G$.
We will use part (ii) of Lemma 2.4 to construct representations of $G$ over fields of prime characteristic from the ring $\mathfrak{R}$. We then relate these representations over prime characteristic to the $\mathbb{C}$-characters of $G$ using Theorem 2.5 and part (i) of Lemma 2.4. An explanation of this method is provided in the brief introduction to the proof of Theorem 3.8. For a deeper treatment, see the overview in [20, Chapter 15], or see [17] for a dedicated analysis of the subject from the perspective of Brauer characters.
3. The character approach
In this section, we generalize (and, in some cases, expand upon) Ott’s results about reversible $((1+s)(1+s^{2}),s^{2}+s+1,s+1)$-DSs [25]. Ott was able to guarantee several restrictions on reversible difference sets corresponding to generalized quadrangles of order $s$ for which $\gcd(v,\sqrt{\Delta})=1$. In Subsection 3.1, we first prove results that will hold for any $v$ and $\sqrt{\Delta}$. In Subsection 3.2, we expand Ott’s results to analyze an arbitrary regular $(v,k,\lambda,\mu)$-partial difference set for which there is at least one prime $p$ dividing $v$ which does not divide $\sqrt{\Delta}$. (Note that, in the abelian setting, $v$ and $\sqrt{\Delta}$ must have the same prime divisors [24, Theorem 6.1], so the results in Subsection 3.2 only apply in the nonabelian setting.) Largely, the results in Subsections 3.1 and 3.2 carry over mutatis mutandis from results in [25]; we include these proofs for the sake of completeness. In Subsection 3.3, we prove results about the intersection of a PDS with cosets of certain normal subgroups; like the results of Subsection 3.2, the results in Subsection 3.3 only apply in the nonabelian setting when $v$ has at least one linear character whose order is coprime to $\sqrt{\Delta}$. Finally, in Subsection 3.4, we provide some examples of applications of these results.
Throughout this section, $G$ will be a group containing a regular $(v,k,\lambda,\mu)$-PDS, $D$, with an irreducible representation $\mathfrak{F}:G\longrightarrow M_{n}(\mathbb{C})$. Any representation $\mathfrak{F}$ given by a mapping into complex matrices can be extended to a $\mathbb{C}[G]$-representation by sending $\sum_{g\in G}c_{g}g$ to $\sum_{g\in G}c_{g}\mathfrak{F}(g)$. For this reason, we do not distinguish between the functions $\mathbb{C}[G]\longrightarrow M_{n}(\mathbb{C})$ and $G\longrightarrow M_{n}(\mathbb{C})$. For an irreducible representation $\mathfrak{F}$, we denote the vector space underlying $\mathfrak{F}$ as $V_{\mathfrak{F}}$. Additionally, ${\rm Irr}(G)$ will be the set of complex irreducible characters of $G$, $\chi$ will always be a general irreducible character, $\xi$ will always be a linear character, and $\mathcal{L}$ will be the group of linear characters.
3.1. Results that hold for any $v$, $\sqrt{\Delta}$
Define the function
$\Phi(h)=|C_{G}(h)||h^{G}\cap D|,$
where $h^{G}=\{g^{-1}hg:g\in G\}$ is the conjugacy class of $h$ in $G$. This is a class function, since $h^{G}=(g^{-1}hg)^{G}$ and $|C_{G}(g^{-1}hg)|=|C_{G}(h)|$. Thus, $\Phi$ may be represented as a sum $\Phi(h)=\sum_{\chi\in{\rm Irr}(G)}c_{\chi}\chi(h)$ of $\chi\in{\rm Irr}(G)$ with coefficients $c_{\chi}$ (see [25, Section 1]).
Lemma 3.1.
For all $g\in G$, $\Phi(g)=\sum_{\chi\in{\rm Irr}(G)}\overline{\chi(D)}\chi(g)$.
Proof.
Denote the distinct conjugacy classes of $G$ by $h_{1}^{G},\dots,h_{r}^{G}$. By the orthogonality relations,
$c_{\chi}$ $=[\Phi,\chi]=\frac{1}{|G|}\sum_{h\in G}\Phi(h)\overline{\chi(h)}=\frac{1}{|G|}\sum_{h\in G}|C_{G}(h)||h^{G}\cap D|\overline{\chi(h)}=\sum_{h\in G}\frac{|C_{G}(h)|}{|G|}|h^{G}\cap D|\overline{\chi(h)}$
$=\sum_{i=1}^{r}\frac{|C_{G}(h)||H_{i}|}{|G|}|h_{i}^{G}\cap D|\overline{\chi(h_{i})}=\sum_{i=1}^{r}|h_{i}^{G}\cap D|\overline{\chi(h_{i})}=\sum_{x\in D}\overline{\chi(x)}=\overline{\chi(D)}.$
So, $\Phi(g)=\sum_{\chi\in{\rm Irr}(G)}c_{\chi}\chi(g)=\sum_{\chi\in{\rm Irr}(G)}\overline{\chi(D)}\chi(g)$ as claimed. ∎
Now, consider a nonprincipal irreducible character $\chi$ of $G$ afforded by the representation $\mathfrak{F}$; compare to [25, Lemma 2.2].
Lemma 3.2.
For $D$ a regular PDS inside $G$ and a representation $\mathfrak{F}$ affording a nonprincipal irreducible character $\chi$ with underlying vector space $V_{\mathfrak{F}}$, $\mathfrak{F}(D)$ has two eigenvalues $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ and $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$ with corresponding eigenspaces $V_{\mathfrak{F}}(\theta_{1})$ and $V_{\mathfrak{F}}(\theta_{2})$ such that $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$, and
$\chi(D)=\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}$.
In particular, if $\chi$ is a nonprincipal linear character, then $\chi(D)=\theta_{1}$ or $\chi(D)=\theta_{2}$.
Proof.
Let $\mathfrak{F}$ be an irreducible representation of $G$ affording $\chi$ with underlying vector space $V_{\mathfrak{F}}$. By Lemma 2.1, $D^{2}=(k-\mu)1+(\lambda-\mu)D+\mu G$. Then,
$\mathfrak{F}(D^{2})$ $=\mathfrak{F}\bigl{(}(k-\mu)1+(\lambda-\mu)D+\mu G\bigr{)}$
$\mathfrak{F}(D^{2})$ $=(k-\mu)\mathfrak{F}(1)+(\lambda-\mu)\mathfrak{F}(D)+\mu(\mathfrak{F}(G))$
$\mathfrak{F}(D^{2})$ $=(k-\mu)I+(\lambda-\mu)\mathfrak{F}(D).$
Note that $\mathfrak{F}(G)=0$, for $\mathfrak{F}$ is an irreducible representation. Therefore,
0 0 $=\mathfrak{F}(D^{2})-(k-\mu)I-(\lambda-\mu)\mathfrak{F}(D).$
Let $\alpha$ be an eigenvalue of $\mathfrak{F}(D)$ with corresponding eigenvector $v$. Then,
0 0 $=\Bigl{(}\mathfrak{F}(D)^{2}-(\lambda-\mu)\mathfrak{F}(D))-(k-\mu)I\Bigr{)}v$
0 0 $=\alpha^{2}+(\mu-\lambda)\alpha+(\mu-k).$
So $\alpha=\frac{1}{2}(\lambda-\mu\pm\sqrt{\Delta})$, with $\Delta=4(k-\mu)+(\lambda-\mu)^{2}$. Thus, $\alpha$ equals one of $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ or $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$ as claimed.
We have just shown that $\mathfrak{F}(D)$ is a solution to the polynomial $z^{2}-(\lambda-\mu)z+\mu-k$, where $z^{2}-(\lambda-\mu)z+\mu-k=(z-\theta_{1})(z-\theta_{2})$ by Lemma 2.2. Moreover, $\theta_{1}\neq\theta_{2}$. For if $\theta_{1}=\theta_{2}$, then we must have $\mu\geqslant k$, but this is impossible, for $\mu\leqslant k$ by definition, and if $\mu=k$, then $\lambda=\mu$ cannot hold. So, $\theta_{1}\neq\theta_{2}$. Therefore, the minimal polynomial of $\mathfrak{F}(D)$ is degree 111 or 222 with distinct roots, so $\mathfrak{F}(D)$ is diagonalizable with $\theta_{1},\theta_{2}$ on the diagonal. If $V_{\mathfrak{F}}(\theta_{i})$ is the eigenspace corresponding to the eigenvectors of $\theta_{i}$, then $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$ and $\chi(D)=\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}$. Note that if $\chi$ is a linear character, then it has one eigenspace of dimension 111, in which case $\chi(D)=\theta_{1}$ or $\chi(D)=\theta_{2}$. ∎
The following is a direct analogue of [25, Lemma 2.3].
Lemma 3.3.
Assume ${\rm Cay}(G,D)$ is not a conference graph and $\mu>0$. Let $\xi\in\mathcal{L}$ be a nonprincipal linear character of order $p$ with kernel $N$. Then, $p|(k-\xi(D))$, $|N\cap D|=\frac{k-\xi(D)}{p}+\xi(D)$, and $|Na\cap D|=\frac{k-\xi(D)}{p}$ for every $a\not\in N$.
Proof.
Now, $D\not\subseteq\ker\xi$, for $\xi$ is a nonprincipal irreducible linear character. Therefore, $\xi(D)=\theta_{i}\neq k$ by Lemma 3.2, so for some $g\in D$ we know that $\xi(g)=\omega$, for $\omega$ a $p^{\text{th}}$ root of unity. Recall from Lemma 3.1 that $\overline{\chi(D)}=\sum_{1=i}^{r}|h_{i}^{G}\cap D|\overline{\chi(h_{i})}$ for conjugacy class representatives $h_{1},\dots,h_{r}$. Then,
0 0 $=\xi(D)-\xi(D)$
$=\sum_{i=1}^{r}|h_{i}^{G}\cap D|\overline{\xi(h_{i})}-\xi(D).$
We re-express this as a sum over the cosets $Ng_{0},\dots,Ng_{p-1}$ for coset representatives $g_{0},\dots,g_{p-1}\in G$. In addition, we may order the $g_{0},\dots,g_{p-1}$ so that $\overline{\xi(Ng_{i})}=\omega^{i}$.
0 0 $=\sum_{i=0}^{p-1}|Ng_{i}\cap D|\overline{\xi(Ng_{i})}-\xi(D)$
$=\sum_{i=1}^{p-1}|Ng_{i}\cap D|\omega^{i}+(|N\cap D|-\xi(D))\omega^{0}$
So, $\omega$ is a root of the polynomial $p(x)=\sum_{i=1}^{p-1}|Ng_{i}\cap D|x^{i}+(|N\cap D|-\xi(D))$. Since $\xi(D)=\theta_{\alpha}\in\mathbb{Q}$ by assumption, then the cyclotomic polynomial $x^{p-1}+\dots+x+1$ divides $p(x)$, forcing
$|N\cap D|-\xi(D)=|Ng_{1}\cap D|=\dots=|Ng_{p-1}\cap D|=c\in\mathbb{Z}.$
Furthermore,
$k=1_{G}(D)=\sum_{j=0}^{p-1}|Ng_{j}\cap D|=c+\xi(D)+(p-1)c=\xi(D)+pc$
so $\frac{k-\xi(D)}{p}=c\in\mathbb{Z}$ and $|N\cap D|=\frac{k-\xi(D)}{p}+\xi(D)$. In particular, $p$ divides $k-\xi(D)$. ∎
We take this opportunity to prove some consequences of these results.
Lemma 3.4.
Assume $G$ contains a regular $(v,k,\lambda,\mu)$-PDS $D$.
  • (i)
    If $\chi_{1}$ and $\chi_{2}$ are two irreducible characters of $G$, $\zeta$ is a primitive $v^{\text{th}}$ root of unity, and $\sigma$ is an automorphism of $\mathbb{Q}(\zeta)$ such that $\chi_{2}=\chi_{1}^{\sigma}$, then $\chi_{2}(D)=\chi_{1}(D)$. In particular, for all $\chi\in{\rm Irr}(G)$, $\chi(D)\in\mathbb{Z}$ and hence $\overline{\chi}(D)=\chi(D)$.
  • (ii)
    If $h_{1}^{G},\dots,h_{r}^{G}$ denote the distinct conjugacy classes of $G$ and $\chi\in{\rm Irr}(G)$, then
    $\sum_{i=1}^{r}|h_{i}^{G}\cap D|\chi(h_{i})=\chi(D).$
  • (iii)
    If $K$ is a normal subgroup of $G$, then
    $\sum_{g\in K}\Phi(g)=|G|\cdot|K\cap D|.$
  • (iv)
    If $m_{i}$ is the multiplicity of the eigenvalue $\theta_{i}$, then
    $m_{i}=\sum_{\mathfrak{F}\text{ irreducible}}(\dim V_{\mathfrak{F}})(\dim V_{\mathfrak{F}}(\theta_{i})).$
Proof.
First, (i) follows directly from Lemma 3.2 (so $\chi_{1}(D)\in\mathbb{Z}$) and the fact that $\sigma|_{\mathbb{Q}}$ is the identity:
$\chi_{2}(D)=\chi_{1}^{\sigma}(D)=\chi_{1}(D)^{\sigma}=\chi_{1}(D).$
Next, for (ii), Lemma 3.1 and its proof imply that
$\sum_{i=1}^{r}|h_{i}^{G}\cap D|\chi(h_{i})=\frac{1}{|G|}\sum_{h\in G}\chi(h)\overline{|h^{G}\cap D|}=[\chi,\Phi]=\chi(D).$
For (iii), if $g_{1}^{G},\dots,g_{s}^{G}$ are the distinct conjugacy classes of $K$, then
$\sum_{g\in K}\Phi(g)$ $=\sum_{g\in G}|g^{G}\cap D||C_{G}(g)|$
$=\sum_{i=1}^{s}|g_{i}^{G}||g_{i}^{G}\cap D||C_{G}(g_{i})|$
$=|G|\cdot\sum_{i=1}^{s}|g_{i}^{G}\cap D|$
$=|G|\cdot|K\cap D|.$
Finally, since the action of $G$ on the associated SRG can be identified with the right regular representation $\rho$ of $G$, and it is well known that its corresponding character $\chi_{\rho}$ satisfies
$\chi_{\rho}=\sum_{\chi\in{\rm Irr}(G)}\chi(1)\chi.$
Lemma 3.2 implies that
$m_{i}=\sum_{\mathfrak{F}\text{ irreducible}}(\dim V_{\mathfrak{F}})(\dim V_{\mathfrak{F}}(\theta_{i}))$
for $i=1,2$, proving (iv). ∎
Remark 3.5.
  • (1)
[…]
  • (2)
    Unlike in the abelian case, while the conditions listed in Lemma 3.2 are necessary, they are not sufficient for the existence of a PDS in a nonabelian group; see Remark 4.8.
  • (3)
    The condition that $D=D^{(-1)}$ is necessary in the proofs of [25, Lemma 2.2] and Lemma 3.2. Unfortunately, most known difference sets are not reversible (and DSs are only conjectured to be reversible in very limited circumstances), so new ideas will be necessary to extend these results to general DSs.
  • (4)
[…]
3.2. Results that hold when a prime $p$ divides $v$ but not $\sqrt{\Delta}$
We now prove an analogue of [25, Corollary 2.1]. Note again that the assumption that $\mu>0$, while necessary, amounts to assuming that the associated SRG is not the union of complete subgraphs of the same size. Going forward, we will tacitly assume $\mu>0$.
Lemma 3.6.
Suppose $p$ is a prime dividing $v$ but not $\sqrt{\Delta}$. Assuming $\mu>0$, there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v_{1}=(k-\theta_{1})/\mu_{1}$, $v_{2}=(k-\theta_{2})/\mu_{2}$, $v=v_{1}v_{2}$, and $p$ divides only one of $v_{1}$, $v_{2}$.
Proof.
Recall for a primitive SRG that $v\mu=(k-\theta_{1})(k-\theta_{2})$ (see [4, Section 1.1.1]). For some factorization of $\mu$ into $\mu_{1}\mu_{2}$, we have $v=\frac{k-\theta_{1}}{\mu_{1}}\frac{k-\theta_{2}}{\mu_{2}}$ where $v_{1}=\frac{k-\theta_{1}}{\mu_{1}}$ and $v_{2}=\frac{k-\theta_{2}}{\mu_{2}}$ are integers. If $p$ divides $v_{1}$ and $v_{2}$,
$0\equiv 0-0\equiv(k-\theta_{2})-(k-\theta_{1})\equiv\theta_{1}-\theta_{2}\equiv\sqrt{\Delta}\pmod{p}$
contradicting our assumption that $p$ and $\Delta$ were relatively prime. ∎
Compare the following result to [25, Lemma 2.4].
Lemma 3.7.
Let $G$ be a group with a regular PDS $D$ and a nontrivial group of linear characters $\mathcal{L}$. Let $P\subseteq\mathcal{L}$ be a Sylow $p$-subgroup of $\mathcal{L}$, and suppose further that $p$ and $\sqrt{\Delta}$ are coprime. Then, for any $\xi^{\prime}\in P$,
$\sum_{\xi\in P}\xi(D)=k+(|P|-1)\xi^{\prime}(D)$.
In particular, all nontrivial linear characters from $P$ are equivalent over $D$.
Proof.
[…]
$S$ $=\sum_{\xi\in P}\xi(D)$
$=1_{G}(D)+\sum_{\xi\in P-\{1_{G}\}}\xi(D)$
[…]
$S$ $=1_{G}(D)+r\theta_{\beta}+(|P|-r-1)\theta_{\alpha}$
$=1_{G}(D)-\theta_{\alpha}+r\theta_{\beta}+(|P|-r)\theta_{\alpha}$
$=k-\theta_{\alpha}+r\theta_{\beta}+(|P|-r)\theta_{\alpha}$
On the other hand, if we define $N_{p}$ to be the kernel of $P$,
$S$ $=\sum_{g\in D}\sum_{\xi\in P}\xi(g)$
$=\sum_{g\in D\cap N_{p}}\sum_{\xi\in P}\xi(g)\ +\sum_{g\in D-N_{p}}\sum_{\xi\in P}\xi(g)$
$=\sum_{\xi\in P}\xi(D\cap N_{p})\ +\sum_{g\in D-N_{p}}0$
$=\sum_{\xi\in P}|D\cap N_{p}|+0$
$=|P||D\cap N_{p}|$
[…]
Theorem 3.8.
Let $\xi\in\mathcal{L}$ be a linear character of order $p^{a}$ with $p$ coprime to $\sqrt{\Delta}$, and let $\chi\in{\rm Irr}(G)$. Then, $\xi\chi(D)=\chi(D)$.
The following proof is quite technical, so we provide a broad explanation of its various moving parts. We would like to show that $\chi(D)=\xi\chi(D)$. To do this, we use Theorem 2.5 to take a $\mathbb{Q}(\zeta)$-representation $\mathfrak{F}$ affording the $\mathbb{C}$-character $\chi$, use Lemma 2.4 to assume that $\mathfrak{F}$ is a $\mathfrak{R}$-representation, and construct a homomorphism, which we denote by $\overline{\phantom{x}}$, from the $\mathfrak{R}$-module underlying $\mathfrak{F}$ to an $L$-module, which is also an $L$-vector space since $L$ is a field. Under this homomorphism, $\overline{\xi(D)\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$, since for all $x\in G$, $\overline{\xi(x)}$ is a scalar whose order divides $p^{a}$ in $L$, so $\overline{\xi(x)}$ must evaluate to 111. We then show that for any matrix $\mathfrak{F}(D)$ with $\mathbb{Q}(\zeta)$-eigenspace $V_{\mathfrak{F}}(\theta_{\alpha})$ and $\mathfrak{R}$-eigenspace $M_{\mathfrak{F}}(\theta_{\alpha})$, that
$\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})={\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha}),$
where $\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})$ is the eigenspace of the representation $\overline{\mathfrak{F}}$ induced by $\mathfrak{F}$ on the $L$-vector space. This finally allows us to conclude
$\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})\stackrel{{\scriptstyle*}}{{=}}\dim_{L}\overline{M_{\overline{\xi\mathfrak{F}}}}(\theta_{\alpha})=\dim_{\mathbb{Q}(\zeta)}V_{\xi\mathfrak{F}}(\theta_{\alpha}),$
where we apply the equality $\overline{\xi\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$ at *. By Lemma 3.2, equality of the dimensions of the eigenspaces is enough to establish $\xi\chi(D)=\chi(D)$.
Proof of Theorem 3.8.
Let $\chi$ be a $\mathbb{C}$-character in ${\rm Irr}(G)$. By Theorem 2.5, we may pick a $\mathbb{Q}(\zeta)$-representation $\mathfrak{F}$ of $G$ affording $\chi$. We denote the $\mathbb{Q}(\zeta)$-vector space underlying $\mathfrak{F}$ as $V_{\mathfrak{F}}$, and we pick a basis $B$ for this space. We define $\mathfrak{R}$ as in Definition 2.3, noting that it has a unique maximal ideal $\mathfrak{p}$ containing $p$ by Lemma 2.4. Now, $V_{\mathfrak{F}}$ contains a $\mathfrak{R}[G]$-module defined by taking all $\mathfrak{R}$-linear combinations of the vectors $\{\mathfrak{F}(g)v_{i}:v_{i}\in B,g\in G\}$. Call this $\mathfrak{R}$-module $M_{\mathfrak{F}}$. Since $\mathfrak{R}$ is a PID by Lemma 2.4, we may pick a basis $w_{1},\dots,w_{n}$ of $M_{\mathfrak{F}}$. This is also a basis of $V_{\mathfrak{F}}$, for $B\subseteq M_{\mathfrak{F}}$. Therefore, we may coordinatize the space $V_{\mathfrak{F}}$ by the basis vectors $w_{1},\dots,w_{n}$, so that all $\mathfrak{F}(g)$ have their entries in $\mathfrak{R}$. Thus, $\mathfrak{F}(D)$ is an $\mathfrak{R}$-matrix, proving part (i) of Lemma 2.4. For further reference, see especially the exposition of Feit [12, 4.1 in Chapter 4].
[…]
$\overline{\xi\mathfrak{F}(D)}\overline{w}$ $=\sum_{x\in D}\overline{\xi(x)\mathfrak{F}(x)}\overline{w}$
$=\sum_{x\in D}(1)\overline{\mathfrak{F}(x)}\overline{w}$
$=\overline{\mathfrak{F}(D)}\overline{w}$
so $\overline{\xi\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$ as claimed.
Now, let $V_{\mathfrak{F}}(\theta_{\alpha})$ and $M_{\mathfrak{F}}(\theta_{\alpha})$ be the $\mathbb{Q}(\zeta)$-eigenspace and $\mathfrak{R}$-eigenspace, respectively, for the eigenvalue $\theta_{\alpha}$ ($\alpha\in\{1,2\}$) of the matrix $\mathfrak{F}(D)$. We prove the sequence of equalities
$\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})={\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha}),$
where $\overline{M_{\overline{\mathfrak{F}}}}$ is the vector space $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$ acted on by $\overline{\mathfrak{F}}$. In particular, $\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})$ is the $\overline{\theta_{\alpha}}$ eigenspace of $\overline{\mathfrak{F}(D)}$ in $\overline{M_{\overline{\mathfrak{F}}}}$.
We first show that $\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})={\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})$. We evidently have ${\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})\leqslant\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})$, for we may pick a basis $v_{1},\dots,v_{k}$ of $M_{\mathfrak{F}}(\theta_{\alpha})$ by [22, Theorem 7.1]. These are linearly independent in $V_{\mathfrak{F}}(\theta_{\alpha})$, for if there exists some $\mathbb{Q}(\zeta)$-linear combination of the $v_{i}$ such that $\frac{a_{1}}{b_{i}}v_{1}+\dots+\frac{a_{n}}{b_{n}}v_{n}=\frac{a_{j}}{b_{j}}v_{j}$, then setting $b=b_{1}\dots b_{n}$, we have $b(\frac{a_{1}}{b_{i}}v_{1}+\dots+\frac{a_{n}}{b_{n}}v_{n})=b\frac{a_{j}}{b_{j}}v_{j}$, a dependence relation in $M_{\mathfrak{F}}(\theta_{\alpha})$, forcing $a_{i}=0$ for each $i$. On the other hand, $\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})\leqslant{\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})$, for if we have a basis of $V_{\mathfrak{F}}(\theta_{\alpha})$, then the $\mathfrak{R}$-multiples of the basis vectors are in $M_{\mathfrak{F}}(\theta_{\alpha})$ and thus linearly independent in $M_{\mathfrak{F}}(\theta_{\alpha})$.
We now prove $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}\leqslant{\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})$. We must show first that there exists a set of basis vectors $v_{1},\dots,v_{k}$ of $M_{\mathfrak{F}}(\theta_{\alpha})$ which do not vanish in $\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$. For this, we use the Elementary Divisor Theorem (see [22, Chapter III, Theorem 7.8]), which states that there exists set of vectors $v_{1},\dots,v_{n}$ of $V_{\mathfrak{F}}(\theta_{\alpha})$ so that for some $r_{1},\dots,r_{k}\in\mathfrak{R}$, the vectors $r_{1}v_{1},\dots,r_{k}v_{k}$ form a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$. Now, if $r_{i}v_{i}$ is an eigenvector of $\mathfrak{F}(D)$, then $v_{i}$ itself must be an eigenvector of $\mathfrak{F}(D)$, for $r_{i}$ is a scalar. Therefore, $v_{i}$ is in the span of the basis, so we may choose $r_{i}$ to be 111. Therefore, $v_{1},\dots,v_{k}$ is a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$. Since $\mathfrak{p}M_{\mathfrak{F}}$ is defined as the set of $\mathfrak{p}$ multiples of the basis vectors of $M_{\mathfrak{F}}$, then $\overline{v_{i}}\neq 0$ for each $v_{i}$. Each $\overline{w}\in\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$ may be expressed as a sum of the $\overline{v_{i}}$ for we may take some pre-image $w$ of $\overline{w}$, and then $a_{1}v_{1}+\dots+a_{k}v_{k}=w$, so $\overline{a_{1}}\overline{v_{1}}+\dots+\overline{a_{k}}\overline{v_{k}}=\overline{w}$ as needed.
The equality $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}={\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})$ is obtained by the fact that if for the basis vectors $v_{1},\dots,v_{n}$ and coefficients $a_{1},\dots,a_{n}\in\mathfrak{R}$, there is a dependence $\overline{a_{1}v_{1}+\dots+a_{n}v_{n}}=0$ in $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$, then $a_{1}v_{1}+\dots+a_{n}v_{n}=b_{1}v_{1}+\dots+b_{n}v_{n}$ where $b_{i}\equiv 0\pmod{\mathfrak{p}}$. Therefore, $(a_{1}-b_{1})v_{1}+\dots+(a_{n}-b_{n})v_{n}=0$, so $a_{1}=b_{1},\dots,a_{n}=b_{n}$, and thus $a_{i}\equiv 0\pmod{\mathfrak{p}}$. So, there exists only a trivial dependence relation among the $\overline{v_{i}}$ in $\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$, forcing ${\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})\leqslant\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$, and ${\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$ by extension.
Finally, we show that $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})$, where $\overline{M_{\overline{\mathfrak{F}}}}$ is the vector space $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$ acted on by $\overline{\mathfrak{F}}$. The inequality $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}\leqslant\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})$ is straightforward, for $\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$ is a vector space contained in $\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})$. Now, ${\rm rk}\;M_{\mathfrak{F}}=\dim V_{\mathfrak{F}}$, and $\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}\leqslant{\rm rk}\;M_{\mathfrak{F}}$. Therefore,
$\dim V_{\mathfrak{F}}=\dim V_{\mathfrak{F}}(\theta_{\alpha})+\dim V_{\mathfrak{F}}(\theta_{\beta})={\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})+{\rm rk}\;M_{\mathfrak{F}}(\theta_{\beta})=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}+\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\beta})},$
and
$\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}+\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\beta})}\leqslant\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})+\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\beta})\leqslant\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}\leqslant\dim V_{\mathfrak{F}}.$
Note, the inequality $\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})+\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\beta})\leqslant\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}$ holds because $\overline{\theta}_{\alpha}\neq\overline{\theta}_{\beta}$, for only in this case is $\overline{\mathfrak{F}}$ diagonalizable. Otherwise, we have $\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})+\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\beta})>\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}$.
Our sequence of inequalities gives
$\dim V_{\mathfrak{F}}=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}+\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\beta})}\leqslant\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})+\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\beta})\leqslant\dim V_{\mathfrak{F}},$
so the inequalities are saturated, forcing
$\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}+\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\beta})}=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})+\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\beta}).$
By the above equality, $\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\beta})}\leqslant\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\beta})$ implies $\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})\leqslant\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$, so we achieve the equality $\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}$.
Therefore, we have shown that $\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})={\rm rk}\;M_{\mathfrak{F}}(\theta_{\alpha})=\dim_{L}\overline{M_{\mathfrak{F}}(\theta_{\alpha})}=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})$. By the equality $\overline{\mathfrak{F}(D)}=\overline{\xi\mathfrak{F}(D)}$ we also have $\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})=\dim_{L}\overline{M_{\overline{\xi\mathfrak{F}}}}(\theta_{\alpha})$. Therefore,
$\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{\alpha})=\dim_{L}\overline{M_{\overline{\mathfrak{F}}}}(\theta_{\alpha})=\dim_{L}\overline{M_{\overline{\xi\mathfrak{F}}}}(\theta_{\alpha})=\dim_{\mathbb{Q}(\zeta)}V_{\xi\mathfrak{F}}(\theta_{\alpha}).$
Combining this equality with Lemma 3.2,
$\xi\chi(D)$ $=(\dim_{\mathbb{Q}(\zeta)}V_{\xi\mathfrak{F}}(\theta_{1}))\theta_{1}+(\dim_{\mathbb{Q}(\zeta)}V_{\xi\mathfrak{F}}(\theta_{2}))\theta_{2}$
$=(\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{1}))\theta_{1}+(\dim_{\mathbb{Q}(\zeta)}V_{\mathfrak{F}}(\theta_{2}))\theta_{2}$
$=\chi(D),$
which proves the theorem. ∎
Corollary 3.9.
Let $\xi,\chi$ be linear characters of order $p^{a},q^{b}$ respectively such that $p$ and $q$ are coprime to $\sqrt{\Delta}$. Then, $\xi(D)=\chi(D)$. In addition, all linear characters of order coprime to $\sqrt{\Delta}$ are equivalent over $D$.
Proof.
By Theorem 3.8, $\chi(D)=\xi\chi(D)=\chi\xi(D)=\xi(D)$. Now, suppose that $\xi$ is a linear character of order coprime to $\sqrt{\Delta}$. By the fundamental theorem of abelian groups, there exist prime power order linear characters $\xi_{1},\dots,\xi_{n}$ such that $\xi_{1}\dots\xi_{n}=\xi$. Again using Theorem 3.8, we have by induction $\xi(D)=\xi_{1}\dots\xi_{n}(D)=\xi_{1}(D)$, for $\xi_{1}=\xi^{\prime}$ of prime power order. For another linear character $\chi$ of order coprime to $\sqrt{\Delta}$, we use the same technique to achieve $\chi(D)=\chi^{\prime}(D)$, for $\chi^{\prime}$ of prime power order. Since $\chi^{\prime}(D)=\xi^{\prime}(D)$, we achieve $\xi(D)=\xi^{\prime}(D)=\chi^{\prime}(D)=\chi(D)$ as claimed. ∎
In order to prove Corollary 3.10, which is useful when ruling out the possibility of certain groups containing a PDS, we introduce the set of primes $\Pi_{\alpha}$ and $\Pi_{\beta}$ with products $\pi_{\alpha}$ and $\pi_{\beta}$ defined so that $\pi_{\alpha}$ and $\pi_{\beta}$ are the largest integers dividing $v_{\alpha}$ and $v_{\beta}$, respectively, such that $\gcd(\sqrt{\Delta},\pi_{\alpha})=\gcd(\sqrt{\Delta},\pi_{\beta})=1$.
Corollary 3.10.
Suppose that $\xi\in\mathcal{L}$ has order dividing $\pi_{\alpha}$. Then, each $\xi\in\mathcal{L}$ with order coprime to $\sqrt{\Delta}$ has an order dividing $\pi_{\alpha}$.
Proof.
Let $\xi\in\mathcal{L}$ of prime order $p$ such that $p$ divides $\pi_{\alpha}$. Let $\xi\in\mathcal{L}$ have prime order $q$ coprime to $\sqrt{\Delta}$. Then, by Corollary 3.9, $\xi(D)=\chi(D)$. So, $k-\xi(D)=k-\theta_{\alpha}=k-\chi(D)$. Using Lemma 3.3, both $p$ and $q$ divide $k-\theta_{\alpha}$. Since $p,q$ divide $v$, and neither $p$ nor $q$ divide $k-\theta_{\beta}$ by Lemma 3.6, we conclude that $q$ must divide $v_{\alpha}$. Since $q$ is coprime to $\sqrt{\Delta}$, we have $q\in\Pi_{\alpha}$. It follows that all primes $q$ dividing $|\mathcal{L}|$ and coprime to $\sqrt{\Delta}$ must divide $\pi_{\alpha}$. Therefore, all $q^{a}$ dividing $|\mathcal{L}|$ must also divide $\pi_{\alpha}$. In conclusion, for every linear character $\xi$ of order $n$ coprime to $\sqrt{\Delta}$, the order $n$ is a product of prime powers $q_{1}^{a_{1}}\dots q_{m}^{a_{m}}$. Since each $q_{i}^{a_{i}}$ is coprime to $\sqrt{\Delta}$, we know $q_{i}^{a_{i}}$ divides $\pi_{\alpha}$, for each $i$, so $q_{1}^{a_{1}}\dots q_{m}^{a_{m}}=n$ divides $\pi_{\alpha}$. ∎
For this next theorem, and for the remainder of the section, let $H$ be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$, and let $N$ be the intersection of the kernels of the characters in $H$. The following theorem is a consequence of Corollary 3.9. The reader may compare Theorem 3.11 to [25, Theorem 2.7].
Theorem 3.11.
If $a\not\in N$, then for $\xi\in H$, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$.
Proof.
[…]
$\Phi(a)$ $=|C_{G}(a)||a^{G}\cap D|$
$=\sum_{\xi\in H}\xi(D)\xi(a)+\sum_{\chi\not\in H}\chi(D)\chi(a).$
By Corollary 3.9, all the nonprincipal $\xi(D)$ are equal. Therefore,
$\Phi(a)$ $=k-\xi(D)+\xi(D)\sum_{\xi\in H}\xi(a)+\sum_{\chi\not\in H}\chi(D)\chi(a)$
By orthogonality, the sum of characters in $H$ is zero, leaving
$\Phi(a)$ $=k-\xi(D)+\sum_{\chi\not\in H}\chi(D)\chi(a)$
We now show that $\sum_{\chi\not\in H}\chi(D)\chi(a)$ is zero. The abelian group $H$ acts on the set of characters by left multiplication. Denote an orbit of $\chi$ under $H$ by $\mathcal{O}$. Then it suffices to show that
0 0 $=\sum_{\tau\in\mathcal{O}}\tau(a)$
We let $H_{\chi}$ be the stabilizer of $\chi$ in $H$. Let $\xi_{1},...,\xi_{r}\in H$ be coset representatives of $H_{\chi}$ in $H$. Therefore,
$\sum_{\tau\in\mathcal{O}}\tau(a)$ $=\sum_{i=1}^{r}\xi_{i}\chi(a)$
$=\left(\sum_{i=1}^{r}\xi_{i}(a)\right)\chi(a)$
We multiply by the sum of elements in the stabilizer of $\chi$.
$\left(\sum_{\eta\in H_{\chi}}\eta(a)\right)\left(\sum_{\tau\in\mathcal{O}}\tau(a)\right)$ $=\left(\sum_{\eta\in H_{\chi}}\eta(a)\right)\left(\sum_{i=1}^{r}\xi_{i}(a)\right)\chi(a)$
$=\left(\sum_{\eta\in H_{\chi}}\sum_{i=1}^{r}\eta\xi_{i}(a)\right)\chi(a)$
$=\left(\sum_{\xi\in H}\xi(a)\right)\chi(a)$
Again, $\left(\sum_{\xi\in H}\xi(a)\right)\chi(a)=0$ by the orthogonality relations. If $\sum_{\eta\in H_{\chi}}\eta(a)\neq 0$, we are done. If it is 0, then there must be some $\eta$ such that $\eta(a)\neq 1$, since $a$ is then not in the kernel of the stabilizer. In this case, $\eta(a)\chi(a)=\chi(a)$, so $\chi(a)=0$. Either way, $\sum_{\chi\not\in H}\chi(D)\chi(a)=0$, so $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$. ∎
Lemma 3.12.
For $a\not\in N$, the greatest common divisor of $|C_{G}(a)|$ and at least one of $\pi_{\alpha},\pi_{\beta}$ is 111.
Proof.
By Theorem 3.11, $|C_{G}(a)|$ divides $k-\xi(D)$. Without loss of generality, we may suppose $\xi(D)=\theta_{\alpha}$. Then, by Lemma 3.6, $\gcd(|C_{G}(a)|,\pi_{\beta})=1$, for otherwise there exists some $p\in\Pi_{\beta}$ dividing $|C_{G}(a)|$, and so by Theorem 3.11, $p$ divides $k-\theta_{\alpha}=|C_{G}(a)||a^{G}\cap D|$. Therefore, $p$ divides both $k-\theta_{\alpha}$ and $k-\theta_{\beta}$, a contradiction. ∎
Once again, we invite the reader to compare the following Theorem 3.13 to [25, Theorem 2.8].
Theorem 3.13.
[…]
$\pi_{\beta}\equiv 1\pmod{p}.$
Proof.
[…]
$1\equiv\frac{|K|}{|N_{K}(P)|}\equiv\frac{\pi_{\alpha}\pi_{\beta}}{|N_{K}(P)|}\equiv\frac{\pi_{\alpha}}{|N_{K}(P)|}\pi_{\beta}\equiv\frac{|U|}{|N_{U}(P)|}\pi_{\beta}\equiv\pi_{\beta}\pmod{p}.$
3.3. Intersections of $D$ with certain normal subgroups
Let $G^{\prime}=[G,G]$ be the derived subgroup of $G$, and let $\mathcal{L}$ be the group of linear characters of $G$; note that $\mathcal{L}\cong G/G^{\prime}$. Recall that we defined $H$ to be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$ and $N$ to be the intersection of the kernels of the characters in $H$. In the case when $\gcd(v,\sqrt{\Delta})=1$, we see that $|\mathcal{L}|$ is coprime to $\sqrt{\Delta}$, so we can choose $H$ to be $\mathcal{L}$, in which case $N=G^{\prime}$. We note first the following result.
Proposition 3.14.
[…]
$\Phi(a)=|C_{G}(a)||a^{G}\cap D|=k-\theta_{\alpha}.$
Proof.
This follows immediately from the above discussion, Corollary 3.9, and Theorem 3.11. ∎
Going forward, we will assume that $N$ is a proper subgroup of $G$ (so that $H$ exists), $\xi(D)=\theta_{\alpha}$ for all nonprincipal linear characters $\xi$ of $H$, and we will denote the eigenvalue that is not $k$ or $\theta_{\alpha}$ by $\theta_{\beta}$. While Proposition 3.14 is quite strong, we will in fact be able to calculate $|N\cap D|$ exactly.
Theorem 3.15.
If $N<G$, then
$|N\cap D|=\frac{k-\theta_{\alpha}}{|H|}+\theta_{\alpha}.$
Proof.
First, we apply Lemma 3.4(iii) to $N$ to obtain
$\sum_{x\in N}\Phi(x)=|G|\cdot|N\cap D|.$
Next, using Proposition 3.14 and recalling that $|G|=|N|\cdot|H|$, we have
$\sum_{x\in N}\Phi(x)$ $=\sum_{g\in G}\Phi(g)-\sum_{g\notin N}\Phi(g)$
$=|G|k-\left(|G|-|N|\right)(k-\theta_{\alpha})$
$=|N|(k-\theta_{\alpha})+|G|\theta_{\alpha}$
Therefore,
$|N\cap D|$ $=\frac{1}{|G|}\left(|N|(k-\theta_{\alpha})+|G|\theta_{\alpha})\right)$
$=\frac{k-\theta_{\alpha}}{|H|}+\theta_{\alpha}.$
Next, we are able to put even more severe restrictions on the intersection of the other cosets of $N$ with $D$.
Theorem 3.16.
[…]
$|Na\cap D|=\frac{k-\theta_{\alpha}}{|H|}.$
Proof.
Note by assumption that $G^{\prime}\subseteq N$, since $G^{\prime}$ is the intersection of the kernels of all linear characters. Moreover, if $b\in a^{G}$, then
$b=a^{g}=g^{-1}ag=(g^{-1}aga^{-1})a=[g,a^{-1}]a\in G^{\prime}a,$
and so $a^{G}\subseteq G^{\prime}a\subseteq Na$, and so $Na$ will be a disjoint union of conjugacy classes. Let $a_{1}^{G},\dots,a_{s}^{G}$ be the distinct conjugacy classes in $Na$. Then, using Proposition 3.14, we have
$|Na\cap D|$ $=\sum_{i=1}^{s}|a_{i}^{G}\cap D|$
$=\sum_{i=1}^{s}\frac{k-\theta_{\alpha}}{|C_{G}(a_{i})|}$
$=\frac{k-\theta_{\alpha}}{|G|}\sum_{i=1}^{s}|a_{i}^{G}|$
$=\frac{(k-\theta_{\alpha})|N|}{|G|}$
$=\frac{k-\theta_{\alpha}}{|H|}.$
When $\gcd(v,\sqrt{\Delta})=1$, it follows that $H=\mathcal{L}$ and $N=G^{\prime}$, so we get the following corollary, which helps to explain why the sizes of conjugacy class intersections with a regular PDS take on very few values in practice in these situations (see Appendix C).
Corollary 3.17.
If $G$ contains a regular $(v,k,\lambda,\mu)$-PDS and $\gcd(v,\sqrt{\Delta})=1$, then
$|G^{\prime}\cap D|=\frac{k-\theta_{\alpha}}{|\mathcal{L}|}+\theta_{\alpha}$
and, for all $a\notin G^{\prime}$,
$|G^{\prime}a\cap D|=\frac{k-\theta_{\alpha}}{|\mathcal{L}|}.$
3.4. Example applications
We now demonstrate how these generalizations of Ott’s work can be applied in other settings.
Proposition 3.18.
There is no group acting regularly on the points of a GQ with order $(q,q^{2}-q)$ when $q$ is even.
Proof.
Let $q$ be even, and assume that a group $G$ acts regularly on the point set of a GQ with order $(q,q^{2}-q)$. This implies the existence of a $((q+1)(q^{2}-q+1),q(q^{2}-q+1),q-1,q^{2}-q+1)$-PDS; in this case, the eigenvalues $\theta_{1},\theta_{2}$ are $q-1$, $-q^{2}+q-1$, and $\sqrt{\Delta}=q^{2}$. Note that if $p|\sqrt{\Delta}$, then $p\not|\;v=(q+1)(q^{2}-q+1)$, so we may apply Theorem 3.13. Thus, we let $p$ divide $v$. Now, $k-\theta_{1}=q^{3}-q^{2}+1$ and $k-\theta_{2}=(q+1)(q^{2}-q+1)$. We show that $v_{1}=k-\theta_{1}$ and $v_{2}=\frac{k-\theta_{2}}{\mu}$. So, suppose $p|\mu$ and $p|k-\theta_{1}$. Then mod $p$,
0 0 $\equiv q^{3}-q^{2}+1$
$\equiv q(q^{2}-q+1)-q+1$
$\equiv 1-q$
$\equiv 1-q-(q^{2}-q+1)$
$\equiv q^{2}$
So $p|q$ and $p|1-q$, forcing $p=1$, so $\mu$ cannot divide $k-\theta_{1}$. Therefore, $v_{1}=q^{3}-q^{2}+1$ and $v_{2}=q+1$.
[…]
Proposition 3.19.
[…]
Proof.
[…]
Now, let $a\in G$ be an element of order divisible by 333 such that $a\not\in[G,G]$. Such an element must exist, for 333 divides $|\mathcal{L}|$. In addition, let $\xi\in\mathcal{L}$. Then, by Theorem 3.11, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$. By Lemma 3.2, $\xi(D)=\pm w$. Set $w=3m+1$ for some $m\in\mathbb{Z}$. Therefore,
$|C_{G}(a)|=\frac{k-(\pm w)}{|a^{G}\cap D|}=\frac{2w^{2}-(\pm w)}{|a^{G}\cap D|}=w\cdot\frac{2w-(\pm 1)}{|a^{G}\cap D|}=w\cdot\frac{6m+2-(\pm 1)}{|a^{G}\cap D|}.$
Since $a$ has order divisible by 333, then 333 divides $|C_{G}(a)|$, so $6m+2-(\pm 1)\equiv 0\pmod{3}$. Since $w\neq 0\pmod{3}$, then $2-(\pm 1)\equiv 0\pmod{3}$. It follows that $\xi(D)=-w$, so that 2(±1)2-($\pm$ 1)2 - ( ± 1 ) equals 333. ∎
4. Computational methods
In the previous section we established certain theoretical results on the characters of groups containing a partial difference set. We now develop related techniques to construct partial difference sets inside of a group using the group characters. In particular, for a partial difference set $D$ inside of a group $G$ with $p^{\ell}$ dividing $\sqrt{\Delta}$, we develop techniques to give the value of $|D\cap h^{G}|\pmod{p^{\ell}}$ for particular primes $p$, and calculate $|D\cap h^{G}|$ by extension.
These methods, combined with Theorem 3.11, makes it rather quick to search for groups which could contain a PDS just by utilizing the irreducible characters of the group. In some cases, simply knowing the values $|D\cap h^{G}|\pmod{p^{\ell}}$ can rule out the existence of a PDS for a certain parameter sets inside of a nonabelian group, such as in Example 4.2. In other cases, targeted information about $|D\cap h^{G}|$ can be used to quickly construct partial difference sets even inside of groups with up to 100010001000 elements, using the search method described in [2].
4.1. Modular computation of conjugacy class intersection sizes for PDSs
Now, as before, let $G$ be a group containing a $(v,k,\lambda,\mu)$-PDS $D$ with eigenvalues $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ and $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$. Let $\mathfrak{F}$ be an irreducible $G$ representation, and let $\chi\in{\rm Irr}(G)$ be afforded by $\mathfrak{F}$. Denote the eigenspace of $\mathfrak{F}(D)$ corresponding to $\theta_{i}$ by $V_{\mathfrak{F}}(\theta_{i})$. For $p$ dividing $\sqrt{\Delta}$, construct the corresponding ring $\mathfrak{R}$ given by Definition 2.3. Let $\ell$ be maximal such that $p^{\ell}$ divides $\sqrt{\Delta}$, and let $(\pi)=\mathfrak{p}$ be the unique maximal ideal containing $p$. Then, $p^{\ell}\in(\pi^{\ell})$, so set $R=\mathfrak{R}/(\pi^{\ell})$. Then, we have the following. (Compare to [32, Lemma 3.1].)
Lemma 4.1.
For $h\in G$ such that $h\neq 1$, suppose that $p^{\ell}$ does not divide $|C_{G}(h)|$. Then, over the ring $R$,
$|h^{G}\cap D||C_{G}(h)|\equiv k-\theta_{2}\pmod{(\pi^{\ell})},$
and if $h=1$, then
$|h^{G}\cap D||C_{G}(h)|\equiv k+\theta_{2}(|G|-1)\pmod{(\pi^{\ell})}.$
In particular,
$|h^{G}\cap D||C_{G}(h)|\equiv k-\theta_{2}\pmod{p^{\ell}},$
and if $h=1$, then
$|h^{G}\cap D||C_{G}(h)|\equiv k+\theta_{2}(|G|-1)\pmod{p^{\ell}}.$
Proof.
First, for a character $\chi$ afforded by a representation $\mathfrak{F}$, $\chi(1)=\dim V_{\mathfrak{F}}(\theta_{1})+\dim V_{\mathfrak{F}}(\theta_{2})$. Moreover, for $p^{\ell}$ a prime power dividing $\sqrt{\Delta}$, $\theta_{1}\equiv\theta_{2}\pmod{p^{\ell}}$. Thus, $\overline{\chi(D)}\equiv\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}\equiv(\dim V_{\mathfrak{F}}(\theta_{1})+\dim V_{\mathfrak{F}}(\theta_{2}))\theta_{2}\equiv\chi(1)\theta_{2}\pmod{p^{\ell}}$. It follows that
$\Phi\equiv k-\theta_{2}+\sum_{\chi\in{\rm Irr}(G)}(\chi(1)\theta_{2})\chi\pmod{(\pi^{\ell})}.$
Then, when $h\neq 1$,
$\Phi(h)\equiv k-\theta_{2}+\theta_{2}\sum_{\chi\in{\rm Irr}(G)}\chi(1)\chi(h)\equiv k-\theta_{2}\pmod{(\pi^{\ell})}$
by column orthogonality of the characters, and when $h=1$, then
$\Phi(1)\equiv k-\theta_{2}+\theta_{2}\sum_{\chi\in{\rm Irr}(G)}\chi(1)^{2}\equiv k-\theta_{2}+\theta_{2}|G|\equiv k+\theta_{2}(|G|-1)\pmod{(\pi^{\ell})}.$
The full equality is given by the definition $\Phi(h)=|h^{G}\cap D||C_{G}(h)|$. Finally, the last equations follow since $|h^{G}\cap D||C_{G}(h)|$ is an integer. ∎
Occasionally, Lemma 4.1 can be used to establish whether or not a PDS exists on a certain parameter set more or less by hand, occasionally aided by minor computer calculations.
Example 4.2.
There does not exist an invertible (112,75,50)(112,75,50)( 112 , 75 , 50 )-DS in a nonabelian group $G$ of order 112112112. We suppose for contradiction that there does exist such a DS in a nonabelian group $G$ of order 112112112. This parameter set has eigenvalues $\theta_{1}=5$ and $\theta_{2}=-5$. By Lemma 3.2, if $D$ exists then $\overline{\chi(D)}\equiv 0\pmod{5}$. Therefore, by Lemma 4.1, $\Phi(g)\equiv 75-5+\sum_{\chi\in{\rm Irr}(G)}(\chi(1)\theta_{2})\chi(g)\equiv 0\pmod{5}$. Finally, there is no centralizer $C_{G}(g)$ of order divisible by 555, since 555 does not divide 112112112. Therefore, for all $g\in G$, $|g^{G}\cap D|\equiv 0\pmod{5}$.
[…]
An analogous computation for each group verifies that there are no nonabelian groups of order 112112112 containing a regular (112,75,50)(112,75,50)( 112 , 75 , 50 )-DS. A direct verification may be done in GAP[15], using the following command:
ΨΨΨgps := Filtered(AllSmallGroups(112), g -> not IsAbelian(g));;
ΨΨΨList(gps, g -> 5*Sum(List(ConjugacyClasses(g), i -> Int(Size(i)/5))));
ΨΨ
The output will be a list, and every entry of the list corresponds to the largest sum of elements in a possible PDS for a given group. Note that for every group, that sum is strictly less than $k=75$.
In fact, we may upgrade this sort of technique to include PDSs more generally. For this, we first require the following Lemma.
Lemma 4.3.
Let $D$ be a regular PDS. If $p^{\ell}$ divides both $\theta_{1}$ and $\theta_{2}$, then $k\equiv 0\pmod{p^{\ell}}$. In particular, $k\equiv 0\pmod{\gcd(\theta_{1},\theta_{2})}$.
Proof.
If $D$ is a regular PDS, then $\Phi(1)=0$. Let $p^{\ell}$ divide $\theta_{1},\theta_{2}$. Let $\mathfrak{R}$ be the local ring corresponding to $p$ as before. Evaluating mod $(\pi^{\ell})$, we have
$0\equiv\Phi(1)\equiv k-\theta_{2}+\theta_{2}\sum_{\chi\in{\rm Irr}(G)}\chi(1)\chi(1)\equiv k\pmod{(\pi^{\ell})},$
by Lemma 4.1. Since $k$ is an integer, $k\equiv 0\pmod{p^{\ell}}$. ∎
Lemma 4.4.
Let $D$ be a regular $PDS$ with eigenvalues $\theta_{1},\theta_{2}$ such that $\gcd(\theta_{1},\theta_{2})\neq 1$. If $p^{\ell}$ divides $\gcd(\theta_{1},\theta_{2})$ and $p^{\ell}$ does not divide $v$, then for each conjugacy class $h^{G}$,
$|h^{G}\cap D|\equiv 0\pmod{p^{\ell}}.$
Proof.
By Lemma 4.1,
$|C_{G}(h)||h^{G}\cap D|\equiv k-\theta_{2}+\theta_{2}\sum_{\chi\in{\rm Irr}(G)}\chi(1)\chi(h)\equiv 0\pmod{(\pi^{\ell})},$
and by Lemma 4.3, $k\equiv 0\pmod{p^{\ell}}$. Since $p^{\ell}$ does not divide $v$, then $p$ does not divide $|C_{G}(h)|$, forcing $|h^{G}\cap D|\equiv 0\pmod{p^{\ell}}$. ∎
On certain parameter sets and for certain groups, Lemma 4.4 in conjunction with Theorem 3.11 allow one to compute very rapidly the structure of the class intersections of $G$ and a PDS $D$. We place here the following example to illustrate this point.
Example 4.5.
We explore the structure of a (183,112,66,72)(183,112,66,72)( 183 , 112 , 66 , 72 )-PDS $D$ in a group $G$ of order 183183183. The eigenvalues of this PDS are $\theta_{1}=4$ and $\theta_{2}=-10$. Then, $\gcd(4,-10)=2$, but 222 does not divide 183183183. Therefore, by Lemma 4.4, for every $h\in G$, we have $|h^{G}\cap D|\equiv 0\pmod{2}$. By a computation in GAP[15], there is one nonabelian group given by SmallGroup(183, 1) with two conjugacy classes $h_{1}^{G},h_{2}^{G}$ of order 616161, 202020 conjugacy classes of order 333, and one conjugacy class of order 111. If $x\in h_{1}^{G},h_{2}^{G}$, then $x\not\in G^{\prime}$ by a computation in GAP[15]. Moreover, every prime $p$ dividing 183183183 does not divide 141414, and $G$ must have a subgroup of linear characters, for it is solvable by Feit-Thompson. So, by Theorem 3.11, $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|\in\{\frac{112-4}{3},\frac{112+10}{3}\}$. Only $\frac{112-4}{3}=36$ is integral, so $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|=36$. Therefore, 11272=40112-72=40112 - 72 = 40 elements must be partitioned among 202020 conjugacy classes. The maximum size of each conjugacy class is 333, and each class must have $0\pmod{2}$ elements, so each of the remaining conjugacy classes contain 222 elements of $D$.
A PDS does indeed exist for this parameter set, which we found through a modified form of the Random Hill Climb search discussed in [2]. We will speak on this more extensively in the next section.
4.2. An algorithmic application of Lemma 4.1
Combining Lemma 4.1 and Theorem 3.11 lets us build a fast algorithm to solve for all values $|D\cap h^{G}|$ even in comparatively large groups of order up to 100010001000. We briefly describe the concept of such an algorithm, and then provide the pseudocode for an implementation.
Suppose $G$ is a fixed nonabelian group with conjugacy class representatives $h_{1},\dots,h_{r}$. We would like to search for a $(v,k,\lambda,\mu)$-PDS $D$ in $G$ with discriminant $\sqrt{\Delta}$. So, we suppose $D$ exists, and attempt to find the values $|h_{i}^{G}\cap D|$. If there exist a prime $p$ dividing $|\mathcal{L}|$ and not $\sqrt{\Delta}$, then for $h\in\{h_{1},\dots,h_{r}\}$ such that $h\not\in G^{\prime}$, we may directly compute $|h^{G}\cap D|$ as one of $\frac{k-\theta_{1}}{|C_{G}(h)|}$ or $\frac{k-\theta_{2}}{|C_{G}(h)|}$ by Theorem 3.11. Otherwise, if no such $p$ exists or $h\in G^{\prime}$, we list the prime powers $p_{1}^{\ell_{1}},\dots,p_{m}^{\ell_{m}}$ dividing $\sqrt{\Delta}$ and not dividing $|C_{G}(h)|$. If $h\neq 1$, we compute $x\equiv|C_{G}(h)|^{-1}(k-\theta_{2})\pmod{p_{i}^{\ell_{i}}}$ for each $p_{i}^{\ell_{i}}$. By Lemma 4.1, $x\equiv|h^{G}\cap D|\pmod{p_{1}^{\ell_{1}}\dots p_{m}^{\ell_{m}}}$.
We then generate every list of values $d_{1},\dots,d_{r}$ so that $d_{1}+\dots+d_{r}=k$, and either $d_{i}$ equals one of $\frac{k-\theta_{1}}{|C_{G}(h)|}$ or $\frac{k-\theta_{2}}{|C_{G}(h)|}$ if $h_{i}$ satisfies the necessary criteria, or $d_{i}\equiv|C_{G}(h)|^{-1}(k-\theta_{2})\pmod{p_{i}^{\ell_{i}}}$ for each of the $p_{i}^{\ell_{i}}$ corresponding to that $h_{i}$. We end by checking if each list $d_{1},\dots,d_{r}$ gives a valid $\Phi$ function according to the identification $\overline{\chi_{j}(D)}=d_{1}\overline{\chi_{j}(h_{1})}+\dots d_{r}\overline{\chi_{j}(h_{r})}$ for all $\chi_{j}\in{\rm Irr}(G)$.
We should mention as well the following fact, which we use as part of Algorithm CCI to determine that certain lists of conjugacy class intersection sizes are invalid and is often successful in ruling out PDSs inside of even order groups.
Lemma 4.6.
Let $G$ be a group containing a regular PDS $D$. If $h\neq h^{-1}$ and $h^{-1}\in h^{G}$, then $|h^{G}\cap D|$ is even.
Proof.
Suppose for $h\in G$ that $h\neq h^{-1}$ but $h^{-1}\in h^{G}$. Since $D=D^{(-1)}$, we can partition $h^{G}\cap D$ into inverse pairs, and so $h^{G}\cap D$ must be even. ∎
It is very easy to rule out a large number of even order groups using Lemma 4.6.
Example 4.7.
We show that there does not exist a (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 )-PDS inside of $G=$ SmallGroup(100, 3). By a computation in GAP[15], $G$ contains a conjugacy class $h^{G}$ of order 444 such that $g,g^{-1}\in h^{G}$ for all $g\in h^{G}$. Moreover, the centralizer $C_{G}(h)$ has order 252525. Since the eigenvalues corresponding to (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 ) are 888 and 3-3- 3, by Lemma 4.1,
$|h^{G}\cap D|\equiv(33-8)25^{-1}\equiv 1\pmod{11}.$
The only nonnegative integer less than or equal to 444 that is congruent to $1\pmod{11}$ is 111, so $|h^{G}\cap D|=1$. Then, by Lemma 4.6, $G$ cannot contain a (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 )-PDS.
Algorithm CCI Construct Class Intersections
[…]
4:DModIntersectionList $\leftarrow$ empty length $r$ list
5:phiLists $\leftarrow$ empty list
6:for $i=1$ to $r$ do
7:  $h\leftarrow h_{i}$
8:  primeDivsDelta $\leftarrow$ prime power divisors of $\sqrt{\Delta}$ coprime to $|C_{G}(h)|$
9:  DHModIntersection $\leftarrow N$ such that $N\equiv|C_{G}(h)|^{-1}(k-\theta_{2})\pmod{p^{\ell}}$ for each $p^{\ell}$ in primeDivsDelta
10:  if $h\not\in G^{\prime}$ then
11:   $C_{1}\leftarrow\frac{k-\theta_{1}}{|C_{G}(h)|}$
12:   $C_{2}\leftarrow\frac{k-\theta_{2}}{|C_{G}(h)|}$
13:   for $p^{\ell}$ in primeDivsDelta do
14:     if DHModIntersection $\pmod{p^{\ell}}\not\in\{C_{1}\pmod{p^{\ell}},C_{2}\pmod{p^{\ell}}\}$ then
15:      Terminate
16:     end if
17:   end for
18:   if $C_{1}$ is not integral and $C_{2}$ is not integral then
19:     Terminate
20:   else if $C_{1}$ is integral and $C_{2}$ is not integral then
21:     DModIntersectionList[i] $\leftarrow[C_{1},0]$
22:   else if $C_{1}$ is not integral and $C_{2}$ is integral then
23:     DModIntersectionList[i] $\leftarrow[C_{2},0]$
24:   else
25:     DModIntersectionList[i] $\leftarrow[C_{2},\frac{\sqrt{\Delta}}{|C_{G}(h)|}]$
26:   end if
27:  else
[…]
29:  end if
30:end for
[…]
32:Filter all $d_{1},\dots,d_{r}$ from phiLists for which the corresponding $\Phi$ is invalid.
33:Output phiLists
The phrase “Filter all $d_{1},\dots,d_{r}$ from phiLists” is rather expansive, so we make the following comment. We first check that each $\chi(D)$ generated by the $d_{1},\dots,d_{r}$ is a valid sum of eigenvalues of the corresponding possible PDS. We also verify the result against Lemma 4.6 and against Theorem 3.16. We should note as well that the full implementation of Algorithm CCI may also be found here. Now, Algorithm CCI runs quickly on most groups for which there is some prime dividing $\sqrt{\Delta}$ not dividing $v$, and it terminates especially rapidly in cases where the number of primes or the size of primes are relatively large. In fact, we were able to compute all values $|D\cap h^{G}|$ of all parameter sets for which $\gcd(v,\sqrt{\Delta})=1$ listed in Brouwer’s tables [3]. We have combined this algorithm with the Random Restart Hill Climb search given by Brady [2] as well as the linear optimization software Gurobi [19] to construct a number of partial difference sets.
Remark 4.8.
The existence of a set of conjugacy class intersection sizes that satisfies the modular constraints of Lemmas 4.1 and 4.6 is not sufficient to guarantee the existence of a PDS. For example, Algorithm CCI returns the list
[0,9,1,9,1,1,1,1,1,1,1,1,1,1,1][0,9,1,9,1,1,1,1,1,1,1,1,1,1,1][ 0 , 9 , 1 , 9 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 ]
as a potentially valid list of conjugacy class intersection sizes for a (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 )-PDS in SmallGroup(111, 1). Note that a subset of this group with these intersection sizes will evaluate on irreducible characters to values admissible by Lemma 3.2. However, by turning the conjugacy class and PDS constraints into a linear program, we are able to rule out the existence of such a PDS using the linear optimization software Gurobi [19], and hence there does not exist a (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 )-PDS.
Similarly, Algorithm CCI returns the lists
[0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][ 0 , 4 , 14 , 1 , 1 , 4 , 4 , 1 , 2 , 4 ] , [ 0 , 4 , 14 , 1 , 1 , 4 , 4 , 2 , 1 , 4 ]
as potentially valid lists of conjugacy class intersections for a (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 )-PDS in SmallGroup(96, 64) and SmallGroup(96, 227), respectively. These were also ruled out with the use of Gurobi [19], and, since these are the only valid lists of conjugacy class intersection sizes for these parameters, there does not exist a (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 )-PDS, either. The authors would be quite interested in a proof of the nonexistence of these PDSs that does not use the brute force of linear programming.
4.3. PDSs constructed using Algorithm CCI
For a number of groups, we used this method to calculate all possible sizes of intersections of the conjugacy classes of the group with a possible PDS inside this group. We then ran a restricted search for PDSs satisfying the class intersection restrictions using a modified Random Restart Hill Climb search described in [2]. Using the technique from [2], we were able to construct nineteen distinct PDSs in nonabelian groups, many of which were already explicitly recognized, such as PDSs arising from triangular graphs. We found three parameter sets, (111,44,19,16)(111,44,19,16)( 111 , 44 , 19 , 16 ), (305,76,27,16)(305,76,27,16)( 305 , 76 , 27 , 16 ), and (981,140,43,16)(981,140,43,16)( 981 , 140 , 43 , 16 ), which correspond to an infinite family of groups acting regularly on Steiner systems ${\mathrm{S}}(2,4,p^{d})$ for $p$ prime. Five others that we were able to construct computationally – in SmallGroup(57, 1), SmallGroup(155, 1), SmallGroup(301, 1), SmallGroup(737, 1), and SmallGroup(1027, 1) – have parameters corresponding to an infinite family of PDSs corresponding to Steiner triple systems. The infinite families on ${\mathrm{S}}(2,4,p)$ and ${\mathrm{S}}(2,3,p^{d})$ correspond to block-regular Steiner 222-design constructions given by Clapham in [6] and Wilson in [33]. Finally, we found a (183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 )-PDS in SmallGroup(183, 1) which is related to a ${\mathrm{S}}(2,5,61)$ Steiner system and, as it turns out, is also part of an infinite family. To the authors knowledge, the PDSs arising from [33] have not previously been recognized in the literature, and we are only aware of one paper [28] recognizing the constructions given by Clapham [6] as giving rise to PDSs.
Proposition 4.9.
Let $p$ be a prime such that $p^{d}>9$ and $p^{d}\equiv 7\pmod{12}$. Then, there exist $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$-PDSs in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
Proof.
By the results of Clapham [6, Section 4], if $p^{d}\equiv 7\pmod{12}$, there exists a Steiner triple system on $p^{d}$ elements with block-regular automorphism group isomorphic to $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$. By the discussion in Subsection 2.3, this will correspond to a PDS in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}$. ∎
Remark 4.10.
Clapham enumerated the number of isomorphism classes of such block-regular Steiner triple systems for each prime power $p^{d}\equiv 7\pmod{12}$ (see [6, Theorem 4.7]). It is well known that two Steiner triple systems are isomorphic if and only if their block-intersection graphs are isomorphic (see, e.g., [8, 26]), so the nonisomorphic Steiner triple systems will yield nonisomorphic SRGs.
In fact, Wilson [33] provided a construction of Steiner 2-designs with a prime power order $p^{d}$ number of elements (provided that $p^{d}$ is sufficiently large in relation to the size $k$ of the blocks), which to the authors’ knowledge has not been mentioned in the literature as giving rise to PDSs.
Theorem 4.11.
Let $p$ be a prime such that $p^{d}>(\frac{1}{2}k(k-1))^{k(k-1)}$. If $p^{d}\equiv k(k-1)+1\pmod{2k(k-1)}$, then there exists a $(p^{d}(p^{d}-1)/(k(k-1)),k(p^{d}-k)/(k-1),(p^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$-PDS in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/(k(k-1))}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
Proof.
Let $\alpha$ be the generator for $({\rm GF}(p^{d})^{\times},\cdot)$, and let $H=\langle\alpha^{(p^{d}-1)/k(k-1)}\rangle$. Let $B$ be a $k$ subset of $({\rm GF}(p^{d})^{\times},\cdot)$, and set =$\mathscr{B}=\{Bg:g\in H\}$. Consider the set of nonzero differences $\delta B=\{b_{i}-b_{j}:b_{i},b_{j}\in B,b_{i}\neq b_{j}\}$ evaluated over the field ${\rm GF}(p^{d})$. If $|\delta B\cap Hg|$ is constant for each coset $Hg$ in $({\rm GF}(p^{d})^{\times},\cdot)$, then by [33, Theorem 2], $\mathscr{B}$ forms the set of base blocks (defined in 2.3) for an ${\mathrm{S}}(2,k,p^{d})$. By [33, Corollary 1], there is always some set $B$ such that $\delta B$ is evenly distributed over the cosets $Hg$ whenever $p^{d}>(\frac{1}{2}k(k-1))^{k(k-1)}$.
It is straightforward to see that this construction provides a block regular construction for an ${\mathrm{S}}(2,k,p^{d})$, for each block is of the form $h+Bg$ for $h\in({\rm GF}(p^{d}),+)$ and $g\in H$. Thus, by the discussion in Subsection 2.3, these constructions correspond to a PDS in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/(k(k-1))}$, which has parameters $(p^{d}(p^{d}-1)/(k(k-1)),k(q^{d}-k)/(k-1),(q^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$. ∎
When $p^{d}\leqslant(\frac{1}{2}k(k-1))^{k(k-1)}$, Buratti [5] provides sufficient conditions for block regular ${\mathrm{S}}(2,k,p^{d})$; we mention one result here, which implies the existence of the (183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 )-PDS we found in SmallGroup(183, 1) related to a block regular ${\mathrm{S}}(2,5,61)$.
Proposition 4.12.
Let $p^{d}=20t+1$ be a prime power, let $2^{e}$ be the highest power of 222 in $t$, and suppose
$\frac{1}{2}(11+5\sqrt{5})\notin\{x^{2^{e+1}}:x\in GF(p^{d})^{\times}\}.$
Then, there exists a $(p^{d}(p^{d}-1)/20,5(p^{d}-5)/4,(p^{d}-1)/4+14,25)$-PDS in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/20}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
Proof.
Let $\alpha$ generate the multiplicative group $({\rm GF}(p^{d})^{\times},\cdot)$. By [5, Theorem 7], there always exists a block $B$ whose differences form a system of distinct representatives for the cosets of the group $H=\langle\alpha^{\frac{p-1}{20}}\rangle$. By [33, Theorem 2], the set =$\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,5,p)$. By the discussion in Subsection 2.3, this gives a $(p^{d}(p^{d}-1)/20,5(p^{d}-5)/4,(p^{d}-1)/4+14,25)$-PDS. ∎
In our own constructions, we have encountered multiple examples of block regular ${\mathrm{S}}(2,4,p)$ for $p\leqslant 6^{12}$. In fact, Fuji-Hara et al. [14] have shown via computer search that for all $p\leqslant 6^{12}$, so long as $p\neq 13$, there does exist a block $B$ satisfying the conditions in [33, Theorem 2] for a block regular ${\mathrm{S}}(2,4,p)$. Thus, we have the following proposition.
Proposition 4.13.
[…]
Proof.
[…]
4.4. Non-modular techniques
The modular techniques are useful for a number of groups and parameter sets. Nevertheless, there exists a large class of parameters $(v,k,\lambda,\mu)$ for which every prime dividing $\sqrt{\Delta}$ divides $v$. (For example, every PDS in an abelian group has such parameters.) In these cases, we frequently cannot guarantee that there is a single prime dividing $\sqrt{\Delta}$ which does not divide $|C_{G}(h)|$ for most of the elements $h\in G$. As a result, we may only compute the modular values of some of the intersections $|h^{G}\cap D|$, and it can be quite computationally intensive to guess and check all possibilities for the remaining $|h^{G}\cap D|$.
Here we provide an example computation, which shows that even if $\sqrt{\Delta}$ divides $v$, it is still possible to rule out the existence of a PDS for a particular group $G$ just through analyses of the character table of $G$ and the possible corresponding $\Phi$. Although the modular techniques of this paper do not apply in this setting, the results of Subsection 3.1 still do. Note that we are mandating that the difference sets in the following example are reversible; by doing so, we do not need to analyze concrete representations of our group $G$ such as in, e.g., [23, 29].
Example 4.14.
We will prove that there is no reversible (144,66,30)(144,66,30)( 144 , 66 , 30 )-DS $D$ in SmallGroup(144, 68). In this case, $\sqrt{\Delta}=12$, where $12^{2}=144$. Therefore, Theorem 3.8, Theorem 3.11, Theorem 3.13, and Algorithm CCI are all inapplicable.
Examining the character table for $G$ (which can be generated, for example, using GAP[15]), there are six conjugacy classes of noncentral elements of order 333. Each of these conjugacy classes has size 161616, so each element in one of these classes has a centralizer of order 999. Thus, for any noncentral element $x$ of order 333 in $G$, we have
$\Phi(x)=9|x^{G}\cap D|\equiv 0\pmod{9}.$
Moreover, if $\mathcal{L}$ denotes the set of linear characters for $G$, then $\chi\notin\mathcal{L}$ implies $\chi(x)=0$. Note that $|\mathcal{L}|=9$, and each nonprincipal character has order 333; other than the principal character, there are four conjugate pairs $\{\xi_{i},\overline{\xi_{i}}\},$ $1\leqslant i\leqslant 4$. Extracting the principal character from this sum and noting that $D$ is reversible (so $\overline{\chi(D)}=\chi(D)$ by Lemma 3.4(i)), we have
$\Phi(x)=66+\sum_{i=1}^{4}\xi_{i}(D)(\xi_{i}(x)+\overline{\xi_{i}}(x)).$
Since each $\xi_{i}(D)$ is ±6$\pm$ 6± 6 by Lemma 3.2, there are only $2^{4}=16$ possibilities in total. Checking which have $\Phi(x)$ divisible by 999, there are only 555 possibilities that work. (In each case, $\Phi(x)=36$ or $\Phi(x)=72$.)
Of the 151515 remaining nonlinear characters, there is one that equals its conjugate and seven conjugate pairs. Let $\chi$ be a nonlinear irreducible character. Note that $\theta_{1}=6$ and $\theta_{2}=-6$. By Lemma 3.2, if $\mathfrak{F}$ is the representation affording $\chi$, then $\chi(D)=6(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))$, where $-\chi(1)\leqslant(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))\leqslant\chi(1)$. So, we may conclude that $\chi(D)\in\{-18,-6,6,18\}$.
Thus, there are $4^{8}$ possibilities for the coefficients $\chi(D)$ of $\Phi$ ($\chi$ nonlinear), and, from what we have ascertained about the linear characters, there are $5\cdot 4^{8}$ possibilities for $\Phi$.
Since either $1\in D$ or $1\notin D$, we have $\Phi(1)\in\{0,144\}$. Of the $5\cdot 4^{8}$ possibilities for $\Phi$, a GAP [15] calculation shows that only 158201582015820 of these possibilites have $\Phi(1)\in\{0,144\}$.
Now, $Z(G)=\langle z\rangle$ is cyclic of order 333. Again, since $D$ is reversible, either $\{z,z^{-1}\}\subseteq D$ or $\{z,z^{-1}\}\cap D=\varnothing$. Thus, either
$\Phi(z)=\Phi(z^{-1})=0$
or
$\Phi(z)=\Phi(z^{-1})=144.$
None of the remaining possibilities for $\Phi$ satisfy either of these constraints (in fact, all the values of these class functions are $6\pmod{12}$ when evaluated at $z$), and so no such $\Phi$ can exist. Therefore, there is no reversible (144,66,30)(144,66,30)( 144 , 66 , 30 )-DS (and hence no regular (144,66,30,30)(144,66,30,30)( 144 , 66 , 30 , 30 )-PDS) in $G$.
Acknowledgements
This work was funded by the Cissy Patterson Fund at William & Mary. The authors acknowledge William & Mary Research Computing for providing computational resources and/or technical support that have contributed to the results reported within this paper. URL: https://www.wm.edu/it/rc
Appendix A Parameters Ruled using Theorem 3.13
The following tables include all parameters $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] that can be ruled out using Theorem 3.13 on either the original parameters or those of its complement, excluding those parameters with $v\in\{512,768,1024\}$, in which cases the computations are too expensive due to the number of groups on those orders. Note, we do not list both a parameter set and its complement.
The modular restriction described by Theorem 3.13 overlaps significantly with the modular restriction in [32, Corollaries 5.2, 5.3]. That is, if a parameter set is ruled out by Theorem 3.13, then it is likely ruled out by [32, Corollaries 5.2, 5.3], and vice versa. So far, we have yet to find a parameter set that Theorem 3.13 rules out, but which [32, Corollaries 5.2, 5.3] does not rule out. For the sake of comprehensiveness, we list all parameter sets which [32, Corollaries 5.2, 5.3] or Theorem 3.13 rule out.
$v$ $k$ $\lambda$ $\mu$
15 6 1 3
28 12 6 4
35 16 6 8
45 12 3 3
45 16 8 4
63 30 13 15
69 20 7 5
77 16 0 4
85 14 3 2
85 20 3 5
85 30 11 10
88 27 6 9
91 24 12 4
99 14 1 2
99 42 21 15
99 48 22 24
105 26 13 4
105 32 4 12
105 52 21 30
115 18 1 3
117 36 15 9
119 54 21 27
133 24 5 4
133 32 6 8
133 44 15 14
143 70 33 35
153 32 16 4
153 56 19 21
175 30 5 5
175 66 29 22
175 72 20 36
176 25 0 4
$v$ $k$ $\lambda$ $\mu$
176 45 18 9
176 70 18 34
176 70 24 30
176 85 48 34
189 48 12 12
195 96 46 48
208 75 30 25
208 81 24 36
209 52 15 12
209 100 45 50
217 66 15 22
217 88 39 33
221 64 24 16
225 96 51 33
231 30 9 3
231 40 20 4
231 70 21 21
231 90 33 36
232 33 2 5
232 63 14 18
232 77 36 20
232 81 30 27
235 42 9 7
235 52 9 12
236 55 18 11
245 52 3 13
245 64 18 16
245 108 39 54
247 54 21 9
249 88 27 33
255 126 61 63
259 42 5 7
$v$ $k$ $\lambda$ $\mu$
261 52 11 10
261 64 14 16
261 80 25 24
261 84 39 21
265 96 32 36
275 112 30 56
279 128 52 64
285 64 8 16
287 126 45 63
297 128 64 48
341 70 15 14
341 84 19 21
341 102 31 30
343 102 21 34
343 114 45 34
344 147 50 72
344 168 92 72
345 120 35 45
345 128 46 48
352 36 0 4
352 39 6 4
352 117 36 40
352 156 60 76
352 171 90 76
357 100 35 25
371 120 44 36
375 110 25 35
375 136 44 52
376 105 32 28
376 175 78 84
376 180 88 84
377 180 81 90
391 140 39 56
391 182 93 77
407 126 45 36
411 130 45 39
413 112 36 28
424 99 26 22
425 72 27 9
425 160 60 60
428 112 21 32
429 108 27 27
435 56 28 4
435 154 53 55
435 182 73 78
437 100 15 25
451 130 33 39
451 156 57 52
459 208 82 104
475 90 25 15
475 96 32 16
477 140 31 45
304 108 42 36
319 150 65 75
323 160 78 80
325 48 24 4
325 54 3 10
325 60 15 10
325 68 3 17
325 72 15 16
325 144 68 60
329 40 3 5
477 168 57 60
483 240 118 120
495 38 1 3
495 78 29 9
495 104 28 20
495 190 53 85
495 208 86 88
495 234 93 126
496 60 30 4
496 110 18 26
496 135 38 36
496 198 80 78
496 231 102 112
496 240 120 112
508 234 100 114
508 247 126 114
511 68 15 8
511 78 5 13
527 256 120 128
533 114 25 24
533 132 31 33
533 154 45 44
536 130 48 26
539 234 81 117
539 250 105 125
553 72 11 9
559 264 116 132
561 60 9 6
561 64 32 4
561 140 31 36
561 176 55 55
561 240 99 105
565 240 104 100
568 217 66 93
568 252 96 124
568 279 150 124
575 112 45 16
575 126 29 27
575 286 141 143
583 96 14 16
585 72 7 9
585 224 88 84
592 216 90 72
595 66 33 4
595 90 5 15
595 108 35 16
595 114 33 19
595 126 21 28
595 132 33 28
595 144 18 40
595 198 81 58
595 216 68 84
609 224 91 77
611 250 105 100
615 294 133 147
616 75 2 10
616 120 20 24
616 205 90 57
616 240 89 96
616 270 108 126
621 60 3 6
621 144 24 36
621 220 79 77
621 300 123 165
625 246 119 82
637 60 11 5
637 96 5 16
637 176 60 44
637 186 35 62
637 252 91 105
639 288 112 144
640 210 66 70
640 213 72 70
645 140 31 30
645 160 38 40
645 184 53 52
645 210 85 60
649 72 15 7
649 216 63 76
649 256 108 96
664 170 24 50
665 280 135 105
667 96 0 16
675 336 166 168
693 192 66 48
697 96 20 12
697 108 9 18
697 348 167 180
703 72 36 4
703 182 81 35
705 256 80 100
705 308 145 126
711 70 5 7
711 230 85 69
715 120 20 20
715 154 33 33
715 224 48 80
715 228 65 76
715 266 105 95
721 192 48 52
721 220 69 66
721 324 147 144
725 196 63 49
729 208 37 68
729 248 67 93
729 280 127 95
731 250 105 75
735 318 109 159
735 360 172 180
736 60 14 4
736 105 20 14
736 180 68 36
736 270 114 90
736 294 122 114
736 315 106 156
736 330 140 154
736 350 162 170
736 357 176 170
736 364 204 156
765 176 28 44
765 192 48 48
771 266 85 95
775 150 45 25
775 288 98 112
779 378 177 189
781 290 93 116
781 348 167 145
783 230 49 75
783 390 193 195
793 162 21 36
793 280 87 105
793 330 147 130
805 324 123 135
805 336 140 140
817 240 83 65
825 96 4 12
825 128 40 16
825 280 75 105
833 112 21 14
833 130 21 20
833 232 81 58
833 256 84 76
837 76 15 6
837 176 40 36
837 196 35 49
837 286 85 104
837 308 127 105
837 396 195 180
845 204 43 51
845 256 108 64
845 396 171 198
847 94 21 9
847 144 26 24
847 270 109 75
847 282 81 100
847 300 117 100
847 360 143 160
847 390 161 195
848 121 24 16
848 231 70 60
848 297 96 108
848 385 156 190
848 418 222 190
851 306 97 117
851 360 163 144
855 126 21 18
855 168 20 36
855 182 37 39
855 336 150 120
856 95 6 11
856 120 21 16
856 209 72 44
861 80 40 4
861 140 31 21
861 280 105 84
861 300 103 105
861 380 163 171
871 60 5 4
871 144 22 24
871 348 137 140
872 208 54 48
875 114 13 15
885 260 55 85
891 290 109 87
891 320 148 96
897 126 15 18
899 448 222 224
901 60 3 4
901 200 45 44
901 224 54 56
901 252 71 70
913 192 56 36
923 432 186 216
925 108 39 9
925 220 75 45
925 330 95 130
925 336 104 132
925 374 123 170
925 378 143 162
925 384 158 160
925 448 192 240
928 300 90 100
928 441 224 196
931 120 11 16
931 138 5 23
931 162 33 27
931 180 29 36
931 248 60 68
931 280 84 84
931 372 143 152
931 390 173 156
931 450 241 195
935 448 204 224
945 64 8 4
945 176 31 33
949 84 11 7
957 256 80 64
969 176 13 36
969 198 57 36
969 392 175 147
973 272 96 68
976 117 28 12
976 273 92 70
976 300 128 76
976 455 198 224
976 480 248 224
987 170 49 25
989 484 231 242
999 448 172 224
1001 136 27 17
1001 150 13 24
1001 160 12 28
1001 280 63 84
1001 300 103 84
1001 384 152 144
1001 450 207 198
1003 300 65 100
1007 486 225 243
1015 312 113 88
1016 259 42 74
1017 344 91 129
1023 336 90 120
1023 510 253 255
1035 88 44 4
1035 330 105 105
1035 440 196 180
1035 462 201 210
1035 490 217 245
1041 336 115 105
1045 234 53 52
1045 260 63 65
1045 276 83 69
1045 290 81 80
1045 324 123 90
1045 384 108 160
1045 396 143 154
1065 266 103 54
1065 364 113 130
1065 416 172 156
1067 208 42 40
1071 110 13 11
1071 320 94 96
1072 252 56 60
1072 261 80 58
1072 345 102 115
1072 391 150 138
1073 64 0 4
1075 192 44 32
1075 210 65 35
1079 238 57 51
1084 513 232 252
1084 532 270 252
1085 256 48 64
1089 320 67 105
1089 408 177 138
1105 80 15 5
1105 138 11 18
1105 208 39 39
1105 336 95 105
1105 378 135 126
1107 378 117 135
$1$ 110 9 11
1112 396 135 144
1125 308 103 77
1131 320 76 96
1135 216 45 40
1135 238 45 51
1136 280 108 56
1147 216 60 36
1155 576 286 288
1161 360 99 117
1161 380 145 114
1177 204 41 34
1177 224 36 44
1184 546 230 270
1184 585 308 270
1189 132 11 15
1189 216 35 40
1189 396 135 130
1189 588 287 294
1199 550 225 275
1208 459 150 189
1215 544 208 272
1215 574 253 287
1215 598 289 299
1216 135 6 16
1216 144 24 16
1216 162 18 22
1216 165 24 22
1216 270 38 66
1216 375 150 100
1216 405 144 130
1216 540 204 268
1216 567 246 280
1216 585 276 286
1216 594 294 286
1216 600 312 280
1225 168 35 21
1225 264 63 55
1225 384 138 112
1225 480 200 180
1225 520 255 195
1233 396 129 126
1233 532 231 228
1241 160 24 20
1241 310 81 76
1241 490 189 196
1247 126 45 9
1251 270 73 54
1261 492 171 205
1261 574 279 246
1267 186 5 31
1269 288 42 72
1271 160 48 16
1275 98 13 7
1275 98 49 4
1275 336 134 72
1275 350 85 100
1275 364 113 100
1275 490 185 190
1275 490 225 165
1275 504 228 180
1275 546 281 198
1275 572 247 264
1285 528 212 220
1288 162 36 18
1288 195 26 30
1288 195 54 25
1288 234 80 34
1288 495 206 180
1288 567 246 252
1295 646 321 323
1216 603 330 268
1221 324 99 81
1221 500 175 225
1221 610 279 330
1225 96 48 4
Appendix B Parameters for Regular PDSs Ruled out Using Algorithm CCI.
The first set of tables tables consist of all parameters $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] satisfying the requirement that $\Delta$ is divisible by some prime $p$ not dividing $v$, for which we have shown no PDS exists through direct computation using Algorithm CCI on either the original parameters or those of the complement. We have only included all parameters with $v<507$, for otherwise the computations become too expensive. Once again, we do not include complements of listed parameters.
$v$ $k$ $\lambda$ $\mu$
10 3 0 1
26 10 3 4
36 14 7 4
40 12 2 4
50 21 8 9
56 10 0 2
66 20 10 4
70 27 12 9
78 22 11 4
82 36 15 16
100 33 14 9
105 40 15 15
112 30 2 10
112 36 10 12
120 28 14 4
120 42 8 18
122 55 24 25
126 25 8 4
126 50 13 24
126 60 33 24
130 48 20 16
135 64 28 32
136 30 15 4
154 48 12 16
154 72 26 40
156 30 4 6
165 36 3 9
170 78 35 36
171 50 13 15
176 40 12 8
176 49 12 14
189 60 27 15
189 88 37 44
190 36 18 4
190 45 12 10
$v$ $k$ $\lambda$ $\mu$
190 84 33 40
190 84 38 36
190 90 45 40
196 39 2 9
196 60 23 16
196 81 42 27
204 28 2 4
204 63 22 18
208 45 8 10
210 38 19 4
220 72 22 24
220 84 38 28
222 51 20 9
226 105 48 49
231 70 21 21
238 75 20 25
243 66 9 21
246 85 20 34
246 105 36 51
246 119 64 51
260 70 15 20
266 45 0 9
273 72 21 18
273 136 65 70
276 44 22 4
276 75 10 24
276 75 18 21
276 110 52 38
276 135 78 54
280 36 8 4
280 117 44 52
286 95 24 35
286 125 60 50
288 105 52 30
288 112 36 48
$v$ $k$ $\lambda$ $\mu$
290 136 63 64
297 40 7 5
297 104 31 39
300 46 23 4
306 55 4 11
306 60 10 12
322 96 20 32
324 68 7 16
324 95 34 25
330 63 24 9
336 80 28 16
336 125 40 50
340 108 30 36
342 33 4 3
342 66 15 12
343 162 81 72
351 70 13 14
351 140 49 60
351 160 64 80
352 26 0 2
352 108 44 28
352 126 50 42
362 171 80 81
364 33 2 3
364 66 20 10
364 88 12 24
364 120 38 40
364 121 48 36
364 165 68 80
364 176 90 80
372 56 10 8
375 102 45 21
375 176 94 72
375 182 85 91
378 52 1 8
378 52 26 4
385 60 5 10
385 168 77 70
392 69 26 9
392 153 54 63
396 135 30 54
396 150 51 60
399 198 97 99
400 21 2 1
400 56 6 8
400 133 42 45
405 132 63 33
405 196 91 98
406 54 27 4
406 108 30 28
406 165 68 66
406 189 84 91
406 195 96 91
408 176 70 80
414 63 12 9
416 100 36 20
416 165 64 66
416 165 64 66
418 147 56 49
430 39 8 3
430 135 36 45
430 165 68 60
438 92 31 16
441 56 7 7
441 190 89 76
441 220 95 124
442 210 99 100
456 65 10 9
456 80 4 16
456 130 24 42
456 140 40 44
456 140 58 36
456 175 78 60
456 182 73 72
456 195 74 90
460 85 18 15
460 99 18 22
460 147 42 49
460 204 78 100
460 216 116 88
460 225 120 100
465 144 43 45
470 126 27 36
474 165 52 60
476 133 42 35
476 133 60 28
484 105 14 25
484 138 47 36
490 165 56 55
490 192 92 64
494 85 12 15
495 190 85 65
495 208 86 88
495 238 109 119
496 54 4 6
498 161 64 46
506 100 18 20
We have additionally computed all parameters $(v,k,\lambda,\mu)$ listed in Brouwer’s tables with $k<(v-1)/2$ and $507\leqslant v\leqslant 1300$ such that $\gcd(v,\Delta)=1$. Of those, the following parameter sets are ruled out by our computations using Algorithm CCI.
$v$ $k$ $\lambda$ $\mu$
528 62 31 4
530 253 120 121
532 81 30 9
546 125 40 25
606 105 4 21
610 87 32 9
626 300 143 144
630 68 1 8
636 250 95 100
658 162 51 36
666 70 35 4
676 150 23 36
676 189 62 49
690 318 121 168
696 125 10 25
700 243 90 81
714 161 40 35
714 248 97 80
$v$ $k$ $\lambda$ $\mu$
726 145 44 25
730 351 168 169
741 74 37 4
742 255 92 85
742 351 180 153
748 180 53 40
754 243 72 81
760 161 30 35
776 124 39 16
780 369 158 189
782 99 36 9
820 78 39 4
842 406 195 196
876 105 38 9
900 203 34 49
900 248 79 64
904 301 78 111
904 378 177 144
$v$ $k$ $\lambda$ $\mu$
946 243 60 63
952 351 150 117
962 465 224 225
990 86 43 4
1000 117 18 13
1026 451 212 187
1036 405 138 171
1044 175 50 25
1054 363 132 121
1086 155 4 25
1086 210 59 36
1090 528 255 256
1106 403 132 155
1116 245 70 49
1120 363 110 121
1140 469 158 217
1148 481 210 195
1155 576 286 288
1156 264 47 64
1156 315 98 81
1176 94 47 4
1178 99 0 9
1190 123 44 9
1206 490 203 196
1210 156 47 16
1212 525 230 225
1222 481 180 195
1226 595 288 289
1236 95 10 7
1236 500 205 200
1248 203 22 35
1254 203 52 29
Finally, we mention again here that the parameter sets (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 ) and (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 ) have been ruled out using a combination of Algorithm CCI and Gurobi [19] (Remark 4.8).
Appendix C Parameter sets with a valid output from Algorithm CCI
We now list all parameter sets $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] satisfying the condition that either $\Delta$ has some distinct prime factor with $v$ and $v\leqslant 507$ or $\gcd(v,\Delta)=1$ and $v\leqslant 1300$, such that there does exist a valid $\Phi$ function, or such that this function was not computable with the available computational resources. We note once again that we exclude parameter sets with $v\in\{512,768,1024\}$, as well as the parameter set (507,176,70,56)(507,176,70,56)( 507 , 176 , 70 , 56 ), for these computations are too expensive. As mentioned, we do not include complements of listed parameters – these may be easily generated from what is provided. We begin with the table of all parameters in groups which were too computationally expensive to compute.
$(v,k,\lambda,\mu)$ ID
(243, 112, 46, 56) 2
(243, 112, 46, 56) 3
(243, 112, 46, 56) 4
(243, 112, 46, 56) 5
(243, 112, 46, 56) 6
(243, 112, 46, 56) 7
(243, 112, 46, 56) 8
(243, 112, 46, 56) 9
(243, 112, 46, 56) 11
(243, 112, 46, 56) 12
(243, 112, 46, 56) 13
(243, 112, 46, 56) 14
(243, 112, 46, 56) 15
(243, 112, 46, 56) 16
(243, 112, 46, 56) 17
(243, 112, 46, 56) 18
(243, 112, 46, 56) 19
(243, 112, 46, 56) 20
(243, 112, 46, 56) 21
(243, 112, 46, 56) 22
(243, 112, 46, 56) 24
(243, 112, 46, 56) 32
(243, 112, 46, 56) 33
(243, 112, 46, 56) 34
(243, 112, 46, 56) 35
$(v,k,\lambda,\mu)$ ID
(243, 112, 46, 56) 36
(243, 112, 46, 56) 37
(243, 112, 46, 56) 38
(243, 112, 46, 56) 39
(243, 112, 46, 56) 40
(243, 112, 46, 56) 41
(243, 112, 46, 56) 42
(243, 112, 46, 56) 43
(243, 112, 46, 56) 44
(243, 112, 46, 56) 45
(243, 112, 46, 56) 46
(243, 112, 46, 56) 47
(243, 112, 46, 56) 49
(243, 112, 46, 56) 50
(243, 112, 46, 56) 51
(243, 112, 46, 56) 52
(243, 112, 46, 56) 53
(243, 112, 46, 56) 54
(243, 112, 46, 56) 55
(243, 112, 46, 56) 56
(243, 112, 46, 56) 57
(243, 112, 46, 56) 58
(243, 112, 46, 56) 59
(243, 112, 46, 56) 60
(243, 112, 46, 56) 62
$(v,k,\lambda,\mu)$ ID
(243, 112, 46, 56) 63
(243, 112, 46, 56) 64
(243, 112, 46, 56) 65
(243, 112, 46, 56) 66
(343, 162, 81, 72) 3
(343, 162, 81, 72) 4
(351, 110, 37, 33) 1
(351, 110, 37, 33) 3
(351, 110, 37, 33) 6
(351, 110, 37, 33) 13
(351, 126, 45, 45) 1
(351, 126, 45, 45) 3
(351, 126, 45, 45) 4
(351, 126, 45, 45) 5
(351, 126, 45, 45) 6
(351, 126, 45, 45) 7
(351, 126, 45, 45) 8
(351, 126, 45, 45) 13
(375, 176, 94, 72) 4
(375, 176, 94, 72) 5
(392, 153, 54, 63) 11
(392, 153, 54, 63) 39
(495, 190, 85, 65) 1
(495, 190, 85, 65) 3
[…]
$(v,k,\lambda,\mu)$ ID ? Conjugacy Class Intersections
(21,10,3,6)(21,10,3,6)( 21 , 10 , 3 , 6 ) 1 [27] [0, 3, 2, 3, 2]
(27,10,1,5)(27,10,1,5)( 27 , 10 , 1 , 5 ) 3 [21] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
4 [30] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(55,18,9,4)(55,18,9,4)( 55 , 18 , 9 , 4 ) 1 [27] [0, 4, 1, 4, 1, 4, 4]
(57,24,11,9)(57,24,11,9)( 57 , 24 , 11 , 9 ) 1 4.9 [0, 9, 1, 9, 1, 1, 1, 1, 1]
(111,44,19,16)(111,44,19,16)( 111 , 44 , 19 , 16 ) 1 4.13 [0, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(125,28,3,7)(125,28,3,7)( 125 , 28 , 3 , 7 ) 3 [21] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
4 [30] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(125,52,15,26)(125,52,15,26)( 125 , 52 , 15 , 26 ) 3 [27] [0, 2, 2, 1, 2, 2, 2, 1, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
4 ? [0, 2, 2, 1, 2, 2, 2, 1, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
(136,30,8,6)(136,30,8,6)( 136 , 30 , 8 , 6 ) 12 ? [0, 3, 3, 8, 2, 3, 3, 3, 3, 2]
(136,60,24,28)(136,60,24,28)( 136 , 60 , 24 , 28 ) 12 ? [0, 7, 10, 4, 4, 7, 7, 10, 7, 4]
(136,63,30,28)(136,63,30,28)( 136 , 63 , 30 , 28 ) 12 ? [0, 7, 7, 13, 4, 7, 7, 7, 7, 4]
(147,66,25,33)(147,66,25,33)( 147 , 66 , 25 , 33 ) 3 [2] [0, 3, 1, 3, 3, 3, 1, 1, 3, 3, 1, 1, 3, 3, 3, 1, 1, 1, 3, 3, 1, 1, 1, 3, 3, 1, 1, 1, 3, 3, 1, 1, 3, 1, 1]
4 [2] [0, 21, 3, 3, 21, 1, 3, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1]
5 [2] [0, 21, 3, 3, 21, 1, 3, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(148,63,22,30)(148,63,22,30)( 148 , 63 , 22 , 30 ) 3 ? [0, 15, 15, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2]
(148,70,36,30)(148,70,36,30)( 148 , 70 , 36 , 30 ) 3 ? [0, 15, 22, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2]
(155,42,17,9)(155,42,17,9)( 155 , 42 , 17 , 9 ) 1 4.9 [0, 9, 1, 9, 9, 1, 9, 1, 1, 1, 1]
(160,54,18,18)(160,54,18,18)( 160 , 54 , 18 , 18 ) 199 ? [0, 6, 6, 0, 0, 0, 6, 6, 6, 6, 6, 6, 6]
199 ? [0, 6, 0, 0, 6, 0, 6, 6, 6, 6, 6, 6, 6]
234 ? [0, 6, 12, 3, 3, 0, 6, 6, 6, 12]
(171,34,17,4)(171,34,17,4)( 171 , 34 , 17 , 4 ) 3 [27] [0, 4, 4, 1, 4, 4, 4, 1, 4, 4, 4]
(171,60,15,24)(171,60,15,24)( 171 , 60 , 15 , 24 ) 3 ? [0, 8, 3, 3, 8, 8, 3, 3, 8, 8, 8]
(183,52,11,16)(183,52,11,16)( 183 , 52 , 11 , 16 ) 1 ? [0, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 ) 1 4.12 [0, 25, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(205,68,15,26)(205,68,15,26)( 205 , 68 , 15 , 26 ) 1 ? [0, 13, 2, 13, 2, 13, 2, 13, 2, 2, 2, 2, 2]
(205,96,50,40)(205,96,50,40)( 205 , 96 , 50 , 40 ) 1 ? [0, 20, 2, 20, 2, 20, 2, 20, 2, 2, 2, 2, 2]
(216,90,39,36)(216,90,39,36)( 216 , 90 , 39 , 36 ) 86 ? [0, 12, 4, 4, 4, 2, 12, 12, 4, 8, 8, 12, 8]
87 ? [0, 8, 8, 4, 2, 2, 8, 16, 16, 8, 2, 8, 8]
88 ? [0, 8, 8, 4, 4, 1, 8, 8, 8, 4, 4, 1, 8, 8, 8, 8]
153 ? [0, 7, 24, 4, 8, 7, 16, 4, 16, 4]
153 ? [0, 2, 24, 4, 8, 2, 16, 9, 16, 9]
(243,112,46,56)(243,112,46,56)( 243 , 112 , 46 , 56 ) 28 ? [0, 18, 4, 0, 2, 0, 18, 12, 4, 4, 0, 2, 0, 12, 12, 4, 4, 12, 4]
28 ? [0, 12, 4, 0, 2, 0, 12, 18, 4, 4, 0, 2, 0, 12, 12, 4, 4, 18, 4]
28 ? [0, 12, 4, 6, 2, 0, 12, 12, 4, 4, 6, 2, 0, 12, 12, 4, 4, 12, 4]
28 ? [0, 12, 6, 0, 2, 0, 12, 12, 6, 6, 0, 2, 0, 12, 12, 6, 6, 12, 6]
29 ? [0, 18, 4, 0, 2, 0, 18, 12, 4, 4, 0, 2, 0, 12, 12, 4, 4, 12, 4]
29 ? [0, 12, 4, 0, 2, 0, 12, 18, 4, 4, 0, 2, 0, 12, 12, 4, 4, 18, 4]
29 ? [0, 12, 4, 6, 2, 0, 12, 12, 4, 4, 6, 2, 0, 12, 12, 4, 4, 12, 4]
29 ? [0, 12, 6, 0, 2, 0, 12, 12, 6, 6, 0, 2, 0, 12, 12, 6, 6, 12, 6]
30 ? [0, 18, 4, 0, 2, 0, 18, 12, 4, 4, 0, 2, 0, 12, 12, 4, 4, 12, 4]
30 ? [0, 12, 4, 0, 2, 0, 12, 18, 4, 4, 0, 2, 0, 12, 12, 4, 4, 18, 4]
30 ? [0, 12, 4, 6, 2, 0, 12, 12, 4, 4, 6, 2, 0, 12, 12, 4, 4, 12, 4]
30 ? [0, 12, 6, 0, 2, 0, 12, 12, 6, 6, 0, 2, 0, 12, 12, 6, 6, 12, 6]
(244,108,42,52)(244,108,42,52)( 244 , 108 , 42 , 52 ) 3 ? [0, 26, 26, 2, 26, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
(244,117,60,52)(244,117,60,52)( 244 , 117 , 60 , 52 ) 3 ? [0, 26, 35, 2, 26, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
(250,81,24,27)(250,81,24,27)( 250 , 81 , 24 , 27 ) 5 ? [0, 9, 3, 6, 0, 9, 3, 3, 6, 0, 9, 3, 0, 0, 9, 3, 0, 3, 9, 3, 3, 0]
5 ? [0, 9, 0, 0, 0, 9, 0, 3, 0, 0, 9, 0, 6, 6, 9, 0, 6, 3, 9, 3, 3, 6]
5 ? [0, 9, 0, 6, 0, 9, 0, 0, 6, 0, 9, 0, 6, 6, 9, 0, 6, 0, 9, 0, 0, 6]
5 ? [0, 9, 0, 6, 0, 9, 0, 3, 6, 0, 9, 0, 3, 3, 9, 0, 3, 3, 9, 3, 3, 3]
5 ? [0, 9, 3, 0, 0, 9, 3, 0, 0, 0, 9, 3, 6, 6, 9, 3, 6, 0, 9, 0, 0, 6]
5 ? [0, 9, 3, 0, 0, 9, 3, 3, 0, 0, 9, 3, 3, 3, 9, 3, 3, 3, 9, 3, 3, 3]
6 ? [0, 9, 3, 6, 0, 9, 3, 3, 6, 0, 9, 3, 0, 0, 9, 3, 0, 3, 9, 3, 3, 0]
6 ? [0, 9, 0, 0, 0, 9, 0, 3, 0, 0, 9, 0, 6, 6, 9, 0, 6, 3, 9, 3, 3, 6]
6 ? [0, 9, 0, 6, 0, 9, 0, 0, 6, 0, 9, 0, 6, 6, 9, 0, 6, 0, 9, 0, 0, 6]
6 ? [0, 9, 0, 6, 0, 9, 0, 3, 6, 0, 9, 0, 3, 3, 9, 0, 3, 3, 9, 3, 3, 3]
6 ? [0, 9, 3, 0, 0, 9, 3, 0, 0, 0, 9, 3, 6, 6, 9, 3, 6, 0, 9, 0, 0, 6]
6 ? [0, 9, 3, 0, 0, 9, 3, 3, 0, 0, 9, 3, 3, 3, 9, 3, 3, 3, 9, 3, 3, 3]
8 ? [0, 9, 6, 6, 0, 9, 6, 6, 6, 0, 9, 0, 0, 0, 9, 9, 0, 6, 0, 0, 0, 0]
(253,42,21,4)(253,42,21,4)( 253 , 42 , 21 , 4 ) 1 [27] [0, 4, 1, 4, 4, 4, 4, 1, 4, 4, 4, 4, 4]
(253,112,36,60)(253,112,36,60)( 253 , 112 , 36 , 60 ) 1 [4] [0, 10, 6, 10, 10, 10, 10, 6, 10, 10, 10, 10, 10]
(273,80,19,25)(273,80,19,25)( 273 , 80 , 19 , 25 ) 3 ? [0, 25, 1, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
4 ? [0, 25, 1, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(273,102,41,36)(273,102,41,36)( 273 , 102 , 41 , 36 ) 3 ? [0, 36, 1, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
4 ? [0, 36, 1, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(301,60,23,9)(301,60,23,9)( 301 , 60 , 23 , 9 ) 1 4.9 [0, 9, 1, 9, 1, 9, 1, 9, 9, 9, 1, 1, 1]
(301,108,27,45)(301,108,27,45)( 301 , 108 , 27 , 45 ) 1 ? [0, 15, 3, 15, 3, 15, 3, 15, 15, 15, 3, 3, 3]
(301,150,65,84)(301,150,65,84)( 301 , 150 , 65 , 84 ) 1 ? [0, 21, 4, 21, 4, 21, 4, 21, 21, 21, 4, 4, 4]
(305,76,27,16)(305,76,27,16)( 305 , 76 , 27 , 16 ) 1 4.13 [0, 16, 1, 16, 1, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(320,99,18,36)(320,99,18,36)( 320 , 99 , 18 , 36 ) 1635 ? [0, 12, 18, 24, 3, 0, 12, 12, 0, 6, 12]
1635 ? [0, 12, 0, 24, 0, 3, 12, 12, 6, 18, 12]
1635 ? [0, 12, 0, 24, 0, 3, 12, 12, 18, 6, 12]
1635 ? [0, 12, 0, 24, 3, 0, 12, 12, 12, 12, 12]
1635 ? [0, 12, 0, 24, 3, 0, 12, 12, 24, 0, 12]
1635 ? [0, 12, 6, 24, 0, 3, 12, 12, 6, 12, 12]
1635 ? [0, 12, 6, 24, 0, 3, 12, 12, 18, 0, 12]
1635 ? [0, 12, 6, 24, 3, 0, 12, 12, 0, 18, 12]
1635 ? [0, 12, 6, 24, 3, 0, 12, 12, 12, 6, 12]
1635 ? [0, 12, 12, 24, 0, 3, 12, 12, 6, 6, 12]
1635 ? [0, 12, 12, 24, 3, 0, 12, 12, 0, 12, 12]
1635 ? [0, 12, 12, 24, 3, 0, 12, 12, 12, 0, 12]
1635 ? [0, 12, 18, 24, 0, 3, 12, 12, 6, 0, 12]
(320,132,46,60)(320,132,46,60)( 320 , 132 , 46 , 60 ) 1635 ? [0, 16, 8, 30, 2, 4, 16, 16, 16, 8, 16]
(343,54,5,9)(343,54,5,9)( 343 , 54 , 5 , 9 ) 3 [21] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
4 [30] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(343,150,53,75)(343,150,53,75)( 343 , 150 , 53 , 75 ) 3 [27] [0, 3, 3, 1, 3, 3, 3, 1, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
4 ? [0, 3, 3, 1, 3, 3, 3, 1, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
(351,50,13,6)(351,50,13,6)( 351 , 50 , 13 , 6 ) 12 ? [0, 3, 7, 7, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
(351,50,25,4)(351,50,25,4)( 351 , 50 , 25 , 4 ) 12 [27] [0, 4, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4]
(351,110,37,33)(351,110,37,33)( 351 , 110 , 37 , 33 ) 4 ? [0, 15, 3, 1, 1, 15, 9, 3, 1, 1, 1, 1, 9, 9, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
4 ? [0, 9, 3, 1, 1, 9, 9, 3, 1, 1, 1, 1, 15, 15, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
5 ? [0, 15, 3, 1, 1, 15, 9, 3, 1, 1, 1, 1, 9, 9, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
5 ? [0, 9, 3, 1, 1, 9, 9, 3, 1, 1, 1, 1, 15, 15, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
7 ? [0, 15, 3, 1, 1, 15, 9, 3, 1, 1, 1, 1, 9, 9, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
8 ? [0, 15, 3, 1, 1, 15, 9, 3, 1, 1, 1, 1, 9, 9, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
12 ? [0, 9, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
(351,112,43,32)(351,112,43,32)( 351 , 112 , 43 , 32 ) 12 ? [0, 9, 2, 2, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
(351,126,45,45)(351,126,45,45)( 351 , 126 , 45 , 45 ) 12 ? [0, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
(351,140,73,44)(351,140,73,44)( 351 , 140 , 73 , 44 ) 12 ? [0, 11, 4, 4, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11]
(363,170,73,85)(363,170,73,85)( 363 , 170 , 73 , 85 ) 2 ? [0, 55, 3, 1, 55, 3, 1, 1, 3, 1, 1, 1, 3, 1, 3, 1, 3, 3, 1, 3, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(381,114,29,36)(381,114,29,36)( 381 , 114 , 29 , 36 ) 1 ? [0, 36, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(381,140,55,49)(381,140,55,49)( 381 , 140 , 55 , 49 ) 1 ? [0, 49, 1, 49, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
(400,156,74,52)(400,156,74,52)( 400 , 156 , 74 , 52 ) 206 ? [0, 20, 20, 10, 10, 2, 20, 4, 20, 10, 4, 20, 16]
207 ? [0, 10, 4, 4, 10, 2, 20, 20, 10, 20, 16, 16, 2, 4, 2, 16]
(405,96,18,24)(405,96,18,24)( 405 , 96 , 18 , 24 ) 15 ? [0, 18, 4, 4, 0, 2, 18, 4, 2, 0, 0, 4, 0, 2, 0, 18, 0, 0, 2, 0, 18]
15 ? [0, 18, 0, 0, 0, 0, 18, 0, 0, 0, 4, 0, 4, 0, 4, 18, 4, 4, 0, 4, 18]
15 ? [0, 18, 0, 0, 0, 2, 18, 0, 2, 0, 2, 0, 4, 2, 2, 18, 2, 2, 2, 4, 18]
15 ? [0, 18, 0, 0, 0, 4, 18, 0, 4, 0, 2, 0, 0, 4, 2, 18, 2, 2, 4, 0, 18]
15 ? [0, 18, 0, 0, 2, 0, 18, 0, 2, 2, 2, 0, 2, 0, 2, 18, 4, 4, 2, 2, 18]
15 ? [0, 18, 0, 0, 2, 0, 18, 0, 2, 2, 4, 0, 4, 0, 4, 18, 0, 0, 2, 4, 18]
15 ? [0, 18, 0, 0, 2, 4, 18, 0, 0, 2, 4, 0, 2, 4, 4, 18, 0, 0, 0, 2, 18]
15 ? [0, 18, 0, 2, 0, 4, 18, 0, 2, 0, 4, 2, 0, 4, 4, 18, 0, 0, 2, 0, 18]
15 ? [0, 18, 0, 2, 2, 2, 18, 0, 2, 2, 2, 2, 2, 2, 2, 18, 0, 0, 2, 2, 18]
(441,88,35,13)(441,88,35,13)( 441 , 88 , 35 , 13 ) 9 ? [0, 1, 1, 1, 1, 1, 7, 3, 1, 3, 7, 7, 3, 7, 3, 1, 1, 7, 3, 3, 7, 7, 3, 3, 7]
(441,152,43,57)(441,152,43,57)( 441 , 152 , 43 , 57 ) 9 ? [0, 5, 5, 1, 1, 5, 19, 7, 5, 7, 3, 11, 7, 11, 7, 1, 1, 19, 7, 7, 3, 3, 7, 7, 3]
9 ? [0, 5, 5, 1, 1, 5, 11, 7, 5, 7, 3, 19, 7, 19, 7, 1, 1, 11, 7, 7, 3, 3, 7, 7, 3]
(441,176,85,60)(441,176,85,60)( 441 , 176 , 85 , 60 ) 9 ? [0, 6, 6, 1, 1, 6, 20, 7, 6, 7, 3, 20, 7, 20, 7, 1, 1, 20, 7, 7, 3, 3, 7, 7, 3]
(441,184,87,69)(441,184,87,69)( 441 , 184 , 87 , 69 ) 9 ? [0, 3, 3, 3, 3, 3, 21, 9, 3, 9, 1, 21, 9, 21, 9, 3, 3, 21, 9, 9, 1, 1, 9, 9, 1]
(448,150,50,50)(448,150,50,50)( 448 , 150 , 50 , 50 ) 178 ? [0, 20, 5, 0, 20, 5, 5, 0, 5, 0, 5, 5, 20, 20, 20, 20]
179 ? [0, 20, 15, 0, 20, 15, 20, 20, 20, 20]
1393 ? [0, 20, 5, 5, 20, 5, 5, 5, 5, 0, 20, 0, 0, 20, 20, 20]
1394 ? [0, 20, 5, 5, 20, 5, 5, 5, 5, 0, 0, 0, 20, 20, 20, 20]
1394 ? [0, 20, 0, 0, 20, 0, 5, 5, 5, 5, 5, 5, 20, 20, 20, 20]
1394 ? [0, 20, 0, 5, 20, 0, 0, 5, 5, 5, 5, 5, 20, 20, 20, 20]
1394 ? [0, 20, 5, 5, 20, 0, 0, 5, 0, 5, 5, 5, 20, 20, 20, 20]
1394 ? [0, 20, 5, 5, 20, 0, 0, 5, 5, 5, 0, 5, 20, 20, 20, 20]
(448,162,66,54)(448,162,66,54)( 448 , 162 , 66 , 54 ) 178 ? [0, 24, 6, 6, 24, 0, 0, 0, 0, 0, 6, 0, 24, 24, 24, 24]
178 ? [0, 24, 0, 6, 24, 0, 3, 3, 3, 3, 0, 0, 24, 24, 24, 24]
179 ? [0, 24, 6, 6, 24, 6, 24, 24, 24, 24]
1393 ? [0, 24, 6, 6, 24, 6, 0, 0, 0, 0, 24, 0, 0, 24, 24, 24]
1393 ? [0, 24, 0, 0, 24, 0, 0, 0, 3, 3, 24, 6, 6, 24, 24, 24]
1393 ? [0, 24, 0, 0, 24, 0, 0, 3, 3, 3, 24, 3, 6, 24, 24, 24]
1393 ? [0, 24, 0, 0, 24, 0, 0, 3, 3, 3, 24, 6, 3, 24, 24, 24]
1393 ? [0, 24, 0, 0, 24, 0, 3, 3, 3, 3, 24, 3, 3, 24, 24, 24]
1394 ? [0, 24, 6, 6, 24, 6, 0, 0, 0, 0, 0, 0, 24, 24, 24, 24]
1394 ? [0, 24, 0, 0, 24, 0, 0, 0, 0, 6, 6, 6, 24, 24, 24, 24]
1394 ? [0, 24, 0, 0, 24, 0, 0, 0, 6, 0, 6, 6, 24, 24, 24, 24]
1394 ? [0, 24, 0, 0, 24, 0, 0, 0, 6, 6, 0, 6, 24, 24, 24, 24]
1394 ? [0, 24, 0, 0, 24, 0, 0, 3, 3, 3, 6, 3, 24, 24, 24, 24]
1394 ? [0, 24, 0, 3, 24, 0, 0, 0, 3, 6, 3, 3, 24, 24, 24, 24]
1394 ? [0, 24, 0, 3, 24, 0, 0, 3, 0, 3, 3, 6, 24, 24, 24, 24]
1394 ? [0, 24, 0, 6, 24, 0, 0, 0, 0, 0, 6, 6, 24, 24, 24, 24]
1394 ? [0, 24, 0, 6, 24, 0, 0, 3, 3, 3, 0, 3, 24, 24, 24, 24]
1394 ? [0, 24, 3, 3, 24, 0, 0, 0, 0, 0, 6, 6, 24, 24, 24, 24]
1394 ? [0, 24, 3, 3, 24, 0, 0, 3, 3, 3, 0, 3, 24, 24, 24, 24]
1394 ? [0, 24, 3, 6, 24, 0, 0, 0, 3, 0, 3, 3, 24, 24, 24, 24]
(465,58,29,4)(465,58,29,4)( 465 , 58 , 29 , 4 ) 1 [27] [ 0, 4, 4, 1, 4, 4, 4, 4, 4, 4, 1, 4, 4, 4, 4, 4, 4 ]
(465,192,72,84)(465,192,72,84)( 465 , 192 , 72 , 84 ) 1 ? [ 0, 6, 14, 6, 6, 14, 14, 14, 14, 14, 6, 14, 14, 14, 14, 14, 14 ]
(497,186,55,78)(497,186,55,78)( 497 , 186 , 55 , 78 ) 1 ? [ 0, 26, 3, 26, 3, 26, 26, 3, 26, 3, 26, 3, 3, 3, 3, 3, 3 ]
(497,240,127,105)(497,240,127,105)( 497 , 240 , 127 , 105 ) 1 ? [ 0, 35, 3, 35, 3, 35, 35, 3, 35, 3, 35, 3, 3, 3, 3, 3, 3 ]
(505,84,3,16)(505,84,3,16)( 505 , 84 , 3 , 16 ) 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16 ]
(505,120,39,25)(505,120,39,25)( 505 , 120 , 39 , 25 ) 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 25, 25, 25, 25 ]
(505,180,53,70)(505,180,53,70)( 505 , 180 , 53 , 70 ) 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 35, 35, 35, 35 ]
(505,224,108,92)(505,224,108,92)( 505 , 224 , 108 , 92 ) 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 46, 46, 46, 46 ]
(507,184,71,64)(507,184,71,64)( 507 , 184 , 71 , 64 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64]
4 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64]
5 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64]
(651,200,55,64)(651,200,55,64)( 651 , 200 , 55 , 64 ) 3 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64]
4 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64]
(651,234,89,81)(651,234,89,81)( 651 , 234 , 89 , 81 ) 3 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 81, 81]
4 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 81, 81]
(657,256,80,112)(657,256,80,112)( 657 , 256 , 80 , 112 ) 3 ? [0, 4, 4, 4, 4, 4, 4, 4, 4, 28, 28, 28, 28, 28, 28, 28, 28]
(657,328,147,180)(657,328,147,180)( 657 , 328 , 147 , 180 ) 3 ? [0, 5, 5, 5, 5, 5, 5, 5, 5, 36, 36, 36, 36, 36, 36, 36, 36]
(689,256,120,80)(689,256,120,80)( 689 , 256 , 120 , 80 ) 1 ? [0, 4, 4, 4, 4, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20]
(737,96,35,9)(737,96,35,9)( 737 , 96 , 35 , 9 ) 1 4.9 [0, 1, 1, 1, 1, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
(755,130,9,25)(755,130,9,25)( 755 , 130 , 9 , 25 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 25, 25, 25, 25]
(755,174,53,36)(755,174,53,36)( 755 , 174 , 53 , 36 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36]
(813,252,71,81)(813,252,71,81)( 813 , 252 , 71 , 81 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 81, 81]
(813,290,109,100)(813,290,109,100)( 813 , 290 , 109 , 100 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 100, 100]
(889,222,35,62)(889,222,35,62)( 889 , 222 , 35 , 62 ) 1 ? [0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 31, 31, 31, 31, 31, 31]
(889,288,112,84)(889,288,112,84)( 889 , 288 , 112 , 84 ) 1 ? [0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 42, 42, 42, 42, 42, 42]
(903,82,41,4)(903,82,41,4)( 903 , 82 , 41 , 4 ) 1 [27] [0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4]
(981,140,43,16)(981,140,43,16)( 981 , 140 , 43 , 16 ) 3 4.13 [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16]
(981,392,133,172)(981,392,133,172)( 981 , 392 , 133 , 172 ) 3 ? [0, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 43, 43, 43, 43, 43, 43, 43, 43]
(981,480,254,216)(981,480,254,216)( 981 , 480 , 254 , 216 ) 3 ? [0, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 54, 54, 54, 54, 54, 54, 54, 54]
(993,310,89,100)(993,310,89,100)( 993 , 310 , 89 , 100 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 100, 100]
(993,352,131,121)(993,352,131,121)( 993 , 352 , 131 , 121 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 121, 121]
(1027,114,41,9)(1027,114,41,9)( 1027 , 114 , 41 , 9 ) 1 4.9 [0, 1, 1, 1, 1, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
(1027,342,81,130)(1027,342,81,130)( 1027 , 342 , 81 , 130 ) 1 ? [0, 5, 5, 5, 5, 5, 5, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26]
(1027,450,225,175)(1027,450,225,175)( 1027 , 450 , 225 , 175 ) 1 ? [0, 5, 5, 5, 5, 5, 5, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35]
(1055,186,17,36)(1055,186,17,36)( 1055 , 186 , 17 , 36 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36]
(1055,238,69,49)(1055,238,69,49)( 1055 , 238 , 69 , 49 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 49, 49, 49, 49]
(1081,90,45,4)(1081,90,45,4)( 1081 , 90 , 45 , 4 ) 1 [27] [0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4]
(1081,486,177,252)(1081,486,177,252)( 1081 , 486 , 177 , 252 ) 1 ? [0, 12, 12, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21]
(1191,374,109,121)(1191,374,109,121)( 1191 , 374 , 109 , 121 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 121, 121]
(1191,420,155,144)(1191,420,155,144)( 1191 , 420 , 155 , 144 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 144, 144]
(1205,448,150,176)(1205,448,150,176)( 1205 , 448 , 150 , 176 ) 1 ? [0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 88, 88, 88, 88]

### v2

@@
 With these techniques, we are able to prove the nonexistence of PDSs in numerous instances and provide severe restrictions in cases when such PDSs may still exist.
-Furthermore, we are able to use these techniques constructively, computing several examples of PDSs in nonabelian groups not previously recognized in the literature, including an infinite family of genuinely nonabelian PDSs associated to the block-regular Steiner triple systems originally studied by Clapham [6] and related infinite families of genuinely nonabelian PDSs associated to the block-regular Steiner 222-designs first studied by Wilson [33].
+Furthermore, we are able to use these techniques constructively, computing several examples of PDSs in nonabelian groups not previously recognized in the literature, including an infinite family of genuinely nonabelian PDSs associated to the block-regular Steiner triple systems originally studied by Clapham [6] and related infinite families of genuinely nonabelian PDSs associated to the block-regular Steiner $2$-designs first studied by Wilson [33].
 1. Introduction
-If $G$ is a group, then we say that $D\subset G$ is a $(v,k,\lambda,\mu)$-partial difference set (PDS) if $G$ has order $v$, $D$ has size $k$, and if the product $gh^{-1}$ for all $g,h\in D$ with $g\neq h$ represents all nonidentity elements in $D$ exactly $\lambda$ times and all nonidentity elements in $G-D$ exactly $\mu$ times.
+If $G$ is a group, then we say that $D\subset G$ is a $(v,k,\lambda,\mu)$-partial difference set (PDS) if $G$ has order $v$, $D$ has size $k$, and if the product $gh^{-1}$ for all $g,h\in D$ with $g\neq h$ represents all nonidentity elements in $D$ exactly $\lambda$ times and all nonidentity elements in GDG-D exactly $\mu$ times.
 We refer to the tuple $(v,k,\lambda,\mu)$ as the parameters of $D$.
@@
 (1) $\chi(D)=\begin{cases}k,\text{ if $\chi$ is the trivial character,}\\ \frac{(\lambda-\mu)\pm\sqrt{\Delta}}{2},\text{ otherwise,}\end{cases}$
-[…]
+where $\Delta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(\lambda-\mu)^{2}+4(k-\mu)$ (see [24, Corollary 3.3]).
 There has been interest recently in PDSs in nonabelian groups [9, 13, 31, 32]; for a recent survey of known results related to PDSs in nonabelian groups, see [27].
@@
 Recently, Ott [25] studied $\Phi$ – which is indeed a class function on $G$ – from a character-theoretic perspective, and Ott used this to prove that GQs of order $s$ ($s+1$ points incident with each line, $s+1$ lines incident with each point), $s$ even, cannot have a group acting regularly on its point set.
-As it turns out, the ideas Ott used are applicable in situations far beyond GQs (and reversible difference sets); in fact, one could argue that the PDS setting is in fact where these techniques are most effective.
+As it turns out, the ideas Ott used are applicable in situations far beyond GQs (and reversible difference sets); in fact, one could argue that the PDS setting is where these techniques are most effective.
 The purpose of this paper is to extend Ott’s results to PDSs in nonabelian groups.
@@
 We have also used our methods to construct several PDSs.
-For $p$ a prime, we demonstrate the existence of an infinite family of PDSs in nonabelian groups related to Steiner triple systems with parameters $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$, where $p^{d}$ is greater than 999 and congruent to $7\pmod{12}$, as well as Steiner 222-designs ${\mathrm{S}}(2,k,v)$ with parameters $(p^{d}(p^{d}-1)/(k(k-1)),k(p^{d}-k)/(k-1),(p^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$ for $p^{d}$ sufficiently large and $p^{d}\equiv k(k-1)+1\pmod{2k(k-1)}$.
-(These Steiner triple systems were shown to be block-regular by Clapham [6], and the Steiner 222-designs were shown to be block regular by Wilson [33].
-It is immediate that these will be PDSs; however, PDS families in Steiner 222-designs do not seem to have been recognized thus far in the literature beyond a recent mention in [28].)
+For $p$ a prime, we demonstrate the existence of an infinite family of PDSs in nonabelian groups related to Steiner triple systems with parameters $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$, where $p^{d}$ is greater than $9$ and congruent to $7\pmod{12}$, as well as Steiner $2$-designs ${\mathrm{S}}(2,k,v)$ with parameters $(p^{d}(p^{d}-1)/(k(k-1)),k(p^{d}-k)/(k-1),(p^{d}-1)/(k-1)+(k-1)^{2}-2,k^{2})$ for $p^{d}$ sufficiently large and $p^{d}\equiv k(k-1)+1\pmod{2k(k-1)}$.
+(These Steiner triple systems were shown to be block-regular by Clapham [6], and the Steiner $2$-designs were shown to be block regular by Wilson [33].
+It is immediate that these will be PDSs; however, PDS families in Steiner $2$-designs do not seem to have been recognized thus far in the literature beyond a recent mention in [28].)
 This paper is organized as follows.
@@
 In Section 4, we introduce computational methods for the case when one of $\sqrt{\Delta}$, $v$ has factors which are relatively prime to the others that are used both to rule out the existence of numerous PDSs and also discover several new examples in nonabelian groups; we also provide an example demonstrating how the techniques from this paper can be used when $\sqrt{\Delta}$ and $v$ have the same primes as factors (such as when $\Delta=v$).
-Finally, we include appendices in which we list parameters ruled out either theoretically or computationally (Appendices A, B) and parameters for which there exists a potentially valid $\Phi$ function (Appendix C).
+Finally, we include appendices in which we list parameters that can be ruled out either theoretically or computationally by methods known already known (Appendix A) or by new methods (Appendix B) as well as a list of open cases and cases in which a PDS exists for sufficiently small $v$ (Appendix C).
 2. Preliminaries
@@
 $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta}),\;\;\;\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$
-[…]
+where $\Delta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}(\lambda-\mu)^{2}+4(k-\mu)$; we call this value the discriminant of the PDS.
+The eigenvalue $k$ is known as the Perron eigenvalue, and it has multiplicity $1$.
+The eigenvalues $\theta_{1}$ and $\theta_{2}$ occur with multiplicities $m_{1}$ and $m_{2}$ in the adjacency matrix of the associated SRG, where
 $m_{i}=\frac{1}{2}\left((v-1)+(-1)^{i}\frac{2k+(v-1)(\lambda-\mu)}{\sqrt{\Delta}}\right)$
@@
 Lemma 2.2.
-Let $\Gamma$ be a $(v,k,\lambda,\mu)$-SRG with eigenvalues $k,\theta_{1},\theta_{2}$ as before. Then,
+Let $\Gamma$ be a $(v,k,\lambda,\mu)$-SRG with eigenvalues $k,\theta_{1},\theta_{2}$. Then,
   • (i)
@@
     $\lambda-\mu=\theta_{1}+\theta_{2}$.
-[…]
+Note that for our purposes $\sqrt{\Delta}$ will always be an integer.
+Indeed, the case when $\sqrt{\Delta}$ is not an integer is quite restricted: the SRG $\Gamma$ must be a conference graph, which has parameters (v,(v1)/2,(v5)/4,(v1)/4)(v,(v-1)/2,(v-5)/4,(v-1)/4); see [16, Lemmas 10.3.2, 10.3.3].
 Furthermore, we say that the $(v,k,\lambda,\mu)$-SRG is imprimitive if either $\mu=0$ or $\mu=k$ and is primitive otherwise.
@@
     If $P$ is a point and $\ell$ a line not incident with $P$, there is exactly one point on $\ell$ collinear with $P$.
-[…]
-One particularly important family of strongly regular graphs are those related to Steiner 222-designs, which we use in Proposition 4.9, Theorem 4.11, Proposition 4.12, and Proposition 4.13.
-A Steiner 222-design, ${\mathrm{S}}(2,k,v)$, is a $v$-element set, along with a collection of $k$-subsets of this set, called blocks, such that any two elements are contained in exactly one block.
-The families of Steiner 222-designs in Proposition 4.9, Theorem 4.11, Propostion 4.12, and Proposition 4.13 are all algebraic constructions using the elements of an abelian group $G$.
-In these cases, the Steiner 222-design is represented by a distinguished set $\mathscr{B}$ of blocks, known as base blocks.
+We denote a $\rm GQ$} of order $(s,t)$ – which may not be unique – by ${\rm GQ}(s,t)$.
+We may construct an incidence graph $\Gamma$ from a GQ $S$ by identifying the points of $S$ with the vertices of $\Gamma$, and specifying that two points are adjacent if they are collinear.
+It is an easy exercise that $\Gamma$ is a $($s+1)(st+1),s(t+1),s1,t+1)$($s+1)(st+1),s(t+1),s-1,t+1)-SRG.
+See [4, Section 2.2.9] for further information on GQs.
+One particularly important family of strongly regular graphs are those related to Steiner $2$-designs, which we use in Proposition 4.9, Theorem 4.11, Proposition 4.12, and Proposition 4.13.
+A Steiner $2$-design, ${\mathrm{S}}(2,k,v)$, is a $v$-element set, along with a collection of $k$-subsets of this set, called blocks, such that any two elements are contained in exactly one block.
+The families of Steiner $2$-designs in Proposition 4.9, Theorem 4.11, Propostion 4.12, and Proposition 4.13 are all algebraic constructions using the elements of an abelian group $G$.
+In these cases, the Steiner $2$-design is represented by a distinguished set $\mathscr{B}$ of blocks, known as base blocks.
 The remaining blocks take the form $g+B$ for $B\in\mathscr{B}$ and $g\in G$.
-A Steiner 222-design has a natural representation as a strongly regular graph if one identifies the blocks of the Steiner 222-design with the vertices of the graph such that two vertices are adjacent if the intersection of their corresponding blocks has one element.
-If $v$ is the number of elements in the Steiner 222-design, then the corresponding SRG has parameters
+A Steiner $2$-design has a natural representation as a strongly regular graph if one identifies the blocks of the Steiner $2$-design with the vertices of the graph such that two vertices are adjacent if the intersection of their corresponding blocks has one element.
+If $v$ is the number of elements in the Steiner $2$-design, then the corresponding SRG has parameters
 $\left(\frac{v(v-1)}{k(k-1)},\frac{k(v-k)}{k-1},\frac{v-1}{k-1}+(k-1)^{2}-2,k^{2}\right).$
-We will speak often of one particular class of Steiner 222-designs, known as Steiner triple systems, which occur when $k=3$.
-For further information on Steiner triple systems, Steiner 222-designs, and Steiner systems more generally, see [7, Section I.4].
+We will speak often of one particular class of Steiner $2$-designs, known as Steiner triple systems, which occur when $k=3$.
+For further information on Steiner triple systems, Steiner $2$-designs, and Steiner systems more generally, see [7, Section I.4].
 2.4. Character theory and representations over local rings
@@
 For a specified prime $p$, and prime ideal $\mathfrak{p}_{0}\subseteq\mathbb{Z}[\zeta]$ such that $p\in\mathfrak{p}_{0}$, define the ring $\mathfrak{R}$ by
-=$\mathfrak{R}=\biggl{\{}\frac{x}{y}:x,y\in\mathbb{Z}[\zeta],y\not\in\mathfrak{p}_{0}\biggr{\}}.$
+$\mathfrak{R}=\biggl\{\frac{x}{y}:x,y\in\mathbb{Z}[\zeta],y\not\in\mathfrak{p}_{0}\biggr\}.$
 $\mathfrak{R}$ is known as the local ring at $\mathfrak{p}_{0}$.
@@
 Lemma 3.2.
-For $D$ a regular PDS inside $G$ and a representation $\mathfrak{F}$ affording a nonprincipal irreducible character $\chi$ with underlying vector space $V_{\mathfrak{F}}$, $\mathfrak{F}(D)$ has two eigenvalues $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ and $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$ with corresponding eigenspaces $V_{\mathfrak{F}}(\theta_{1})$ and $V_{\mathfrak{F}}(\theta_{2})$ such that $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$, and
+For $D$ a regular $(v,k,\lambda,\mu)$-PDS inside $G$ and a representation $\mathfrak{F}$ affording a nonprincipal irreducible character $\chi$ with underlying vector space $V_{\mathfrak{F}}$, $\mathfrak{F}(D)$ has two eigenvalues $\theta_{1}=\frac{1}{2}(\lambda-\mu+\sqrt{\Delta})$ and $\theta_{2}=\frac{1}{2}(\lambda-\mu-\sqrt{\Delta})$ with corresponding eigenspaces $V_{\mathfrak{F}}(\theta_{1})$ and $V_{\mathfrak{F}}(\theta_{2})$ such that $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$, and
 $\chi(D)=\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}$.
@@
 Note that $\mathfrak{F}(G)=0$, for $\mathfrak{F}$ is an irreducible representation. Therefore,
-0 0 $=\mathfrak{F}(D^{2})-(k-\mu)I-(\lambda-\mu)\mathfrak{F}(D).$
+$0$ $=\mathfrak{F}(D^{2})-(k-\mu)I-(\lambda-\mu)\mathfrak{F}(D).$
 Let $\alpha$ be an eigenvalue of $\mathfrak{F}(D)$ with corresponding eigenvector $v$. Then,
-0 0 $=\Bigl{(}\mathfrak{F}(D)^{2}-(\lambda-\mu)\mathfrak{F}(D))-(k-\mu)I\Bigr{)}v$
-0 0 $=\alpha^{2}+(\mu-\lambda)\alpha+(\mu-k).$
+$0$ $=\Bigl(\mathfrak{F}(D)^{2}-(\lambda-\mu)\mathfrak{F}(D))-(k-\mu)I\Bigr)v$
+$0$ $=\alpha^{2}+(\mu-\lambda)\alpha+(\mu-k).$
 So $\alpha=\frac{1}{2}(\lambda-\mu\pm\sqrt{\Delta})$, with $\Delta=4(k-\mu)+(\lambda-\mu)^{2}$.
@@
 So, $\theta_{1}\neq\theta_{2}$.
-Therefore, the minimal polynomial of $\mathfrak{F}(D)$ is degree 111 or 222 with distinct roots, so $\mathfrak{F}(D)$ is diagonalizable with $\theta_{1},\theta_{2}$ on the diagonal.
+Therefore, the minimal polynomial of $\mathfrak{F}(D)$ is degree $1$ or $2$ with distinct roots, so $\mathfrak{F}(D)$ is diagonalizable with $\theta_{1},\theta_{2}$ on the diagonal.
 If $V_{\mathfrak{F}}(\theta_{i})$ is the eigenspace corresponding to the eigenvectors of $\theta_{i}$, then $V_{\mathfrak{F}}=V_{\mathfrak{F}}(\theta_{1})\oplus V_{\mathfrak{F}}(\theta_{2})$ and $\chi(D)=\dim V_{\mathfrak{F}}(\theta_{1})\theta_{1}+\dim V_{\mathfrak{F}}(\theta_{2})\theta_{2}$.
-Note that if $\chi$ is a linear character, then it has one eigenspace of dimension 111, in which case $\chi(D)=\theta_{1}$ or $\chi(D)=\theta_{2}$.
+Note that if $\chi$ is a linear character, then it has one eigenspace of dimension $1$, in which case $\chi(D)=\theta_{1}$ or $\chi(D)=\theta_{2}$.
 ∎
@@
 Lemma 3.3.
-Assume ${\rm Cay}(G,D)$ is not a conference graph and $\mu>0$.
-Let $\xi\in\mathcal{L}$ be a nonprincipal linear character of order $p$ with kernel $N$.
+Assume $G$ contains a regular $(v,k,\lambda,\mu)$-PDS with nonprincipal eigenvalues $\theta_{1}$, $\theta_{2}$, ${\rm Cay}(G,D)$ is not a conference graph, and $\mu>0$.
+Let $\mathcal{L}$ be a nontrivial group of linear characters of $G$, and let $\xi\in\mathcal{L}$ be a nonprincipal linear character of order $p$ with kernel $N$.
 Then, $p|(k-\xi(D))$, $|N\cap D|=\frac{k-\xi(D)}{p}+\xi(D)$, and $|Na\cap D|=\frac{k-\xi(D)}{p}$ for every $a\not\in N$.
@@
 Then,
-0 0 $=\xi(D)-\xi(D)$
+$0$ $=\xi(D)-\xi(D)$
 $=\sum_{i=1}^{r}|h_{i}^{G}\cap D|\overline{\xi(h_{i})}-\xi(D).$
@@
 In addition, we may order the $g_{0},\dots,g_{p-1}$ so that $\overline{\xi(Ng_{i})}=\omega^{i}$.
-0 0 $=\sum_{i=0}^{p-1}|Ng_{i}\cap D|\overline{\xi(Ng_{i})}-\xi(D)$
+$0$ $=\sum_{i=0}^{p-1}|Ng_{i}\cap D|\overline{\xi(Ng_{i})}-\xi(D)$
 $=\sum_{i=1}^{p-1}|Ng_{i}\cap D|\omega^{i}+(|N\cap D|-\xi(D))\omega^{0}$
@@
   • (1)
-[…]
+    We again remark that the results in this subsection do not require that $G$ be an abelian or nonabelian group or any conditions on the factors of $|G|$ or $\sqrt{\Delta}$.
+Thus, these results may be applied even in cases when our techniques in the rest of the paper do not otherwise work; an example of such an application is given in Subsection 4.4
   • (2)
@@
   • (4)
-[…]
+    The main technique in [25] and our own Theorem 3.8 is to extract information about $D$ by analyzing $\mathfrak{F}(D)$ over fields of characteristic $p$, when $p$ does not divide $\sqrt{\Delta}=\theta_{1}-\theta_{2}$ but does divide $|G|$.
+A critical element in this proof is that $\mathfrak{F}(D)$ induces a representation on a certain field $L$ of characteristic $p$, and that the induced representation $\overline{\mathfrak{F}(D)}$ on $L$ is diagonalizable, for its minimal polynomial is $(z-\theta_{1})(z-\theta_{2})\pmod{p}$, where $\theta_{1}\not\equiv\theta_{2}\pmod{p}$.
+As a result, while the results of this subsection still hold, the techniques in the following subsections largely fail if we ever have $|G|$ sharing all of its prime divisors with $\theta_{1}-\theta_{2}$, for then the minimal polynomial of $\mathfrak{F}(D)$ for any representation $\mathfrak{F}$ is $(z-\theta_{1})(z-\theta_{2})\equiv(z-\theta_{1})^{2}\pmod{p}$, so $\mathfrak{F}(D)$ is not diagonalizable.
+Unfortunately, this case happens reasonably often, so for such parameter sets techniques involving local rings are ineffective.
+In Subsection 4.4, we provide an example computation for how character theory may still be applied to PDSs where $\sqrt{\Delta}$ does divide $v$.
 3.2. Results that hold when a prime $p$ divides $v$ but not $\sqrt{\Delta}$
@@
 Lemma 3.6.
+Let $G$ be a finite group containing a regular $(v,k,\lambda,\mu)$-PDS $D$ that has nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\sqrt{\Delta}=\theta_{1}-\theta_{2}$.
 Suppose $p$ is a prime dividing $v$ but not $\sqrt{\Delta}$.
-Assuming $\mu>0$, there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v_{1}=(k-\theta_{1})/\mu_{1}$, $v_{2}=(k-\theta_{2})/\mu_{2}$, $v=v_{1}v_{2}$, and $p$ divides only one of $v_{1}$, $v_{2}$.
+Assuming $\mu>0$, there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v_{1}=(k-\theta_{1})/\mu_{1}$, $_{2}=(k-\theta_{2})/\mu_{2}, v=v$$v_{1}v_{2}$, and $p$ divides only one of $v_{1}$, $v_{2}$.
 Proof.
@@
 Lemma 3.7.
-Let $G$ be a group with a regular PDS $D$ and a nontrivial group of linear characters $\mathcal{L}$.
-Let $P\subseteq\mathcal{L}$ be a Sylow $p$-subgroup of $\mathcal{L}$, and suppose further that $p$ and $\sqrt{\Delta}$ are coprime.
-Then, for any $\xi^{\prime}\in P$,
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\mathcal{L}$ be the group of linear characters of $G$.
+Let $P\subseteq\mathcal{L}$ be a Sylow $p$-subgroup of $\mathcal{L}$, and suppose further that $p$ and $\sqrt{\Delta}=\theta_{1}-\theta_{2}$ are coprime.
+Then, for any nontrivial $\xi^{\prime}\in P$,
 $\sum_{\xi\in P}\xi(D)=k+(|P|-1)\xi^{\prime}(D)$.
@@
 Proof.
-[…]
+First, $|P|$ divides $k-\xi(D)$ by Lemmas 3.6 and 3.3.
+Set $S\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\sum_{\xi\in P}\xi(D)$.
+Then,
 $S$ $=\sum_{\xi\in P}\xi(D)$
 $=1_{G}(D)+\sum_{\xi\in P-\{1_{G}\}}\xi(D)$
-[…]
+Fix $\theta_{\alpha}$ to be an output of at least one $\xi\in P$ on $D$, and let $r$ of the remaining $\xi$ evaluate to $\theta_{\beta}$ on $D$.
+Then there are |P|r1|P|-r-1 eigenvalues $\theta_{\alpha}$ in the sum.
+So,
 $S$ $=1_{G}(D)+r\theta_{\beta}+(|P|-r-1)\theta_{\alpha}$
@@
 $=|P||D\cap N_{p}|$
-[…]
+Recall that $|P|$ divides $k-\xi(D)$.
+Thus, evaluating $k-\theta_{\alpha}+r\theta_{\beta}+(|P|-r)\theta_{\alpha}$ modulo $|P|$ gives $0\equiv r\theta_{\beta}-r\theta_{\alpha}\equiv\pm r\sqrt{\Delta}$.
+Further, $\sqrt{\Delta}$ and $|P|$ are coprime, so $r=|P|$ or $r=0$.
+If $r=|P|$, then every $\xi$ evaluates to $\theta_{\beta}$, which contradicts our selection of $\theta_{\alpha}$ to be the output of some $\xi$, so $r=0$.
+Therefore, all the linear characters of $P$ are equivalent over $D$.
+∎
 […]
 Theorem 3.8.
-Let $\xi\in\mathcal{L}$ be a linear character of order $p^{a}$ with $p$ coprime to $\sqrt{\Delta}$, and let $\chi\in{\rm Irr}(G)$. Then, $\xi\chi(D)=\chi(D)$.
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\mathcal{L}$ be the group of linear characters of $G$.
+Let $\xi\in\mathcal{L}$ be a linear character of order $p^{a}$ with $p$ coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $\chi\in{\rm Irr}(G)$.
+Then, $\xi\chi(D)=\chi(D)$.
 The following proof is quite technical, so we provide a broad explanation of its various moving parts.
@@
 To do this, we use Theorem 2.5 to take a $\mathbb{Q}(\zeta)$-representation $\mathfrak{F}$ affording the $\mathbb{C}$-character $\chi$, use Lemma 2.4 to assume that $\mathfrak{F}$ is a $\mathfrak{R}$-representation, and construct a homomorphism, which we denote by $\overline{\phantom{x}}$, from the $\mathfrak{R}$-module underlying $\mathfrak{F}$ to an $L$-module, which is also an $L$-vector space since $L$ is a field.
-Under this homomorphism, $\overline{\xi(D)\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$, since for all $x\in G$, $\overline{\xi(x)}$ is a scalar whose order divides $p^{a}$ in $L$, so $\overline{\xi(x)}$ must evaluate to 111.
+Under this homomorphism, $\overline{\xi(D)\mathfrak{F}(D)}=\overline{\mathfrak{F}(D)}$, since for all $x\in G$, $\overline{\xi(x)}$ is a scalar whose order divides $p^{a}$ in $L$, so $\overline{\xi(x)}$ must evaluate to $1$.
 We then show that for any matrix $\mathfrak{F}(D)$ with $\mathbb{Q}(\zeta)$-eigenspace $V_{\mathfrak{F}}(\theta_{\alpha})$ and $\mathfrak{R}$-eigenspace $M_{\mathfrak{F}}(\theta_{\alpha})$, that
@@
 For further reference, see especially the exposition of Feit [12, 4.1 in Chapter 4].
-[…]
+We define the submodule $\mathfrak{p}M_{\mathfrak{F}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\{w\in M_{\mathfrak{F}}:w=qv,q\in\mathfrak{p},v\in M_{\mathfrak{F}}\}$.
+This is a closed subspace of $M_{\mathfrak{F}}$, so $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$ is well-defined.
+We therefore define the natural projection $\overline{\phantom{x}}:M_{\mathfrak{F}}\longrightarrow M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$.
+For an $\mathfrak{R}$-representation $\mathfrak{F}$, there is a corresponding representation $\overline{\mathfrak{F}}$ on $M_{\mathfrak{F}}/\mathfrak{p}M_{\mathfrak{F}}$ by $\overline{\mathfrak{F}(g)}\overline{w}=\overline{\mathfrak{F}(g)w}$.
+As before, $\overline{\xi(g)}\equiv 1\pmod{\mathfrak{p}}$, so for any vector $w\in M_{\mathfrak{F}}$, we have $\overline{\xi(g)}\overline{w}=\overline{w}$.
+Thus, for all $w\in V_{\mathfrak{F}}$,
 $\overline{\xi\mathfrak{F}(D)}\overline{w}$ $=\sum_{x\in D}\overline{\xi(x)\mathfrak{F}(x)}\overline{w}$
@@
 Now, if $r_{i}v_{i}$ is an eigenvector of $\mathfrak{F}(D)$, then $v_{i}$ itself must be an eigenvector of $\mathfrak{F}(D)$, for $r_{i}$ is a scalar.
-Therefore, $v_{i}$ is in the span of the basis, so we may choose $r_{i}$ to be 111.
+Therefore, $v_{i}$ is in the span of the basis, so we may choose $r_{i}$ to be $1$.
 Therefore, $v_{1},\dots,v_{k}$ is a basis for $M_{\mathfrak{F}}(\theta_{\alpha})$.
@@
 Corollary 3.9.
-Let $\xi,\chi$ be linear characters of order $p^{a},q^{b}$ respectively such that $p$ and $q$ are coprime to $\sqrt{\Delta}$.
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$.
+Let $\xi,\chi$ be nontrivial linear characters of $G$ of prime-power orders $p^{a},q^{b}$, respectively, such that $p$ and $q$ are coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$.
 Then, $\xi(D)=\chi(D)$.
-In addition, all linear characters of order coprime to $\sqrt{\Delta}$ are equivalent over $D$.
+Moreover, all linear characters of order coprime to $\sqrt{\Delta}$ are equivalent over $D$.
 Proof.
@@
 ∎
-In order to prove Corollary 3.10, which is useful when ruling out the possibility of certain groups containing a PDS, we introduce the set of primes $\Pi_{\alpha}$ and $\Pi_{\beta}$ with products $\pi_{\alpha}$ and $\pi_{\beta}$ defined so that $\pi_{\alpha}$ and $\pi_{\beta}$ are the largest integers dividing $v_{\alpha}$ and $v_{\beta}$, respectively, such that $\gcd(\sqrt{\Delta},\pi_{\alpha})=\gcd(\sqrt{\Delta},\pi_{\beta})=1$.
+Definition 3.10.
+Recall from Lemma 3.6 that, if there exists a prime $p$ dividing $v$ but not $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, then $k-\theta_{1}$ and $k-\theta_{2}$ are coprime, and so there exists a factorization $\mu=\mu_{1}\mu_{2}$ such that $v=v_{1}v_{2}$, where $v_{\alpha}=(k-\theta_{\alpha})\mu_{\alpha}$ for $\alpha=1,2$, and $p$ divides only one of $v_{1}$ or $v_{2}$.
+We define the set of primes $\Pi_{\alpha}$ and $\Pi_{\beta}$ ($\{\alpha,\beta\}=\{1,2\}$) with products $\pi_{\alpha}$ and $\pi_{\beta}$ defined so that $\pi_{\alpha}$ and $\pi_{\beta}$ are the largest integers dividing $v_{\alpha}$ and $v_{\beta}$, respectively, such that $\gcd(\sqrt{\Delta},\pi_{\alpha})=\gcd(\sqrt{\Delta},\pi_{\beta})=1$.
 Corollary 3.11.
-Suppose that $\xi\in\mathcal{L}$ has order dividing $\pi_{\alpha}$. Then, each $\xi\in\mathcal{L}$ with order coprime to $\sqrt{\Delta}$ has an order dividing $\pi_{\alpha}$.
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, and let $\mathcal{L}$ be the group of linear characters of $G$.
+Suppose $\pi_{\alpha}$ is defined as in Definition 3.10, and suppose that $\xi\in\mathcal{L}$ has order dividing $\pi_{\alpha}$.
+Then, each $\xi\in\mathcal{L}$ with order coprime to $\sqrt{\Delta}$ has an order dividing $\pi_{\alpha}$.
 Proof.
@@
 ∎
-For this next theorem, and for the remainder of the section, let $H$ be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$, and let $N$ be the intersection of the kernels of the characters in $H$.
-The following theorem is a consequence of Corollary 3.9.
-The reader may compare Theorem 3.11 to [25, Theorem 2.7].
+The following theorem is a consequence of Corollary 3.9. The reader may compare Theorem 3.12 to [25, Theorem 2.7].
 Theorem 3.12.
-If $a\not\in N$, then for $\xi\in H$, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$.
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$.
+If $a\not\in N$, then for nontrivial $\xi\in H$, $|C_{G}(a)||a^{G}\cap D|=k-\xi(D)$.
 Proof.
-[…]
+Let $\xi\in H$ be a linear character of order coprime to $|G|$. Let $\chi$ be an arbitrary irreducible character of $G$. We have
 $\Phi(a)$ $=|C_{G}(a)||a^{G}\cap D|$
@@
 Then it suffices to show that
-0 0 $=\sum_{\tau\in\mathcal{O}}\tau(a)$
+$0$ $=\sum_{\tau\in\mathcal{O}}\tau(a)$
 We let $H_{\chi}$ be the stabilizer of $\chi$ in $H$. Let $\xi_{1},...,\xi_{r}\in H$ be coset representatives of $H_{\chi}$ in $H$. Therefore,
@@
 Lemma 3.13.
-For $a\not\in N$, the greatest common divisor of $|C_{G}(a)|$ and at least one of $\pi_{\alpha},\pi_{\beta}$ is 111.
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, let $N$ be the intersection of the kernels of the characters in $H$, and let $\pi_{\alpha},\pi_{\beta}$ be defined as in Definition 3.10.
+For $a\not\in N$, the greatest common divisor of $|C_{G}(a)|$ and at least one of $\pi_{\alpha},\pi_{\beta}$ is $1$.
 Proof.
@@
 Proof.
-[…]
-[…]
+Since $p$ is coprime to $\sqrt{\Delta}$, and $p$ divides $|\mathcal{L}|$, we have that $H$ is nontrivial.
+Thus, $N\neq G$, so the hypotheses for Lemma 3.13 apply.
+Suppose there exists $q$ dividing $|N_{G}(P)|$ such that $q\in\Pi_{\beta}$.
+Then, there is an element $g\in N_{G}(P)$ of order $q$.
+By [18, Theorem 3.5], $P=C_{P}(g)[\langle g\rangle,P]$.
+Each element in $C_{P}(g)$ centralizes $g$, so each element $a\in C_{P}(g)$ contains $g$ in its centralizer.
+Thus, for some $m$, $\gcd(|C_{G}(a)|,\pi_{\beta})=q\cdot m$, and since $a\in C_{P}(g)$ and $a\in C_{G}(a)$, then the order of $a$ divides both $|P|$ and $|C_{G}(a)|$.
+Thus, $\gcd(|C_{G}(a)|,|P|)>1$, forcing $\gcd(|C_{G}(a)|,\pi_{\alpha})>1$.
+Therefore, by the contrapositive of Lemma 3.13, $a$ is in $N$.
+Thus, $P=C_{P}(g)[\langle g\rangle,P]\subseteq NG^{\prime}\subseteq N$ (where $G^{\prime}=[G,G]$), given that $G^{\prime}\subseteq N$, since for each $\xi\in H$, we must have $G^{\prime}\subseteq\ker\xi$.
+So, $P\subseteq N$.
+But, $p$ divides $|\mathcal{L}|$, a contradiction.
+Therefore, $\gcd(|N_{G}(P)|,\pi_{\beta})=1$.
+Now, suppose in addition that $G$ is a solvable group.
+By Hall’s Theorem, there is a subgroup $K$ of $G$ which is a $\pi_{\alpha}\pi_{\beta}$-group.
+Moreover, since $G$ is solvable, $K$ is solvable, so there exists a subgroup $U$ of $K$ which is a $\pi_{\alpha}$-group.
+Since $p\in\Pi_{\alpha}$, then $|P|$ divides $\pi_{\alpha}$.
+Therefore, $P$ is contained in some Hall subgroup, $L$.
+Any two Hall subgroups are conjugate, and thus there exists $g\in G$ such that $g^{-1}Lg=K$.
+So, $K$ contains a Sylow $p$-subgroup, $g^{-1}Pg$.
+Therefore, we are justified in assuming that $P\subseteq K$, forcing $P\subseteq U$.
+Thus by our preceding argument, $N_{K}(P)=N_{K}(U)$.
+Therefore, using Sylow III,
 $1\equiv\frac{|K|}{|N_{K}(P)|}\equiv\frac{\pi_{\alpha}\pi_{\beta}}{|N_{K}(P)|}\equiv\frac{\pi_{\alpha}}{|N_{K}(P)|}\pi_{\beta}\equiv\frac{|U|}{|N_{U}(P)|}\pi_{\beta}\equiv\pi_{\beta}\pmod{p}$.
@@
 Let $G^{\prime}=[G,G]$ be the derived subgroup of $G$, and let $\mathcal{L}$ be the group of linear characters of $G$; note that $\mathcal{L}\cong G/G^{\prime}$.
-Recall that we defined $H$ to be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$ and $N$ to be the intersection of the kernels of the characters in $H$.
+We will define $H$ to be a subgroup of linear characters with order coprime to $\sqrt{\Delta}$ and $N$ to be the intersection of the kernels of the characters in $H$.
 In the case when $\gcd(v,\sqrt{\Delta})=1$, we see that $|\mathcal{L}|$ is coprime to $\sqrt{\Delta}$, so we can choose $H$ to be $\mathcal{L}$, in which case $N=G^{\prime}$.
@@
 Proposition 3.15.
-[…]
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$.
+If $|H|\neq\{1\}$ and $\xi(D)=\theta_{\alpha}$ for some nonprincipal linear character $\xi$ in $H$, then $\xi(D)=\theta_{\alpha}$ for all nonprincipal $\xi\in H$.
+Moreover, for all $a\in G\backslash N$,
 $\Phi(a)=|C_{G}(a)||a^{G}\cap D|=k-\theta_{\alpha}.$
@@
 Theorem 3.16.
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$.
 If $N<G$, then
-$|N\cap D|=\frac{k-\theta_{\alpha}}{|H|}+\theta_{\alpha}.$
+$|N\cap D|=\frac{k-\theta_{\alpha}}{|H|}+\theta_{\alpha},$
+where $\xi(D)=\theta_{\alpha}$ for all nonprincipal $\xi\in H$.
 Proof.
-First, we apply Lemma 3.4(iii) to $N$ to obtain
+First, we apply Lemma 3.4(iii) to NN to obtain
 $\sum_{x\in N}\Phi(x)=|G|\cdot|N\cap D|.$
@@
 Theorem 3.17.
-[…]
-$|Na\cap D|=\frac{k-\theta_{\alpha}}{|H|}.$
+Let $G$ be a group with a regular $(v,k,\lambda,\mu)$-PDS $D$ with nonprincipal eigenvalues $\theta_{1}>\theta_{2}$, let $H$ be a subgroup of linear characters of $G$ with order coprime to $\sqrt{\Delta}=\theta_{1}-\theta_{2}$, and let $N$ be the intersection of the kernels of the characters in $H$.
+If $N<G$, and $a\in G\backslash N$, then
+$|Na\cap D|=\frac{k-\theta_{\alpha}}{|H|},$
+where $\xi(D)=\theta_{\alpha}$ for all nonprincipal $\xi\in H$.
 Proof.
@@
 Corollary 3.18.
-If $G$ contains a regular $(v,k,\lambda,\mu)$-PDS and $\gcd(v,\sqrt{\Delta})=1$, then
+Assume the group $\mathcal{L}$ of linear characters of $G$ is nontrivial.
+If $G$ contains a regular $(v,k,\lambda,\mu)$-PDS, $\gcd(v,\sqrt{\Delta})=1$, and $\xi(D)=\theta_{\alpha}$ for a nontrivial $\xi\in\mathcal{L}$, then
 $|G^{\prime}\cap D|=\frac{k-\theta_{\alpha}}{|\mathcal{L}|}+\theta_{\alpha}$
@@
 Proof.
-Let $q$ be even, and assume that a group $G$ acts regularly on the point set of a GQ with order $(q,q^{2}-q)$.
-This implies the existence of a $((q+1)(q^{2}-q+1),q(q^{2}-q+1),q-1,q^{2}-q+1)$-PDS; in this case, the eigenvalues $\theta_{1},\theta_{2}$ are $q-1$, $-q^{2}+q-1$, and $\sqrt{\Delta}=q^{2}$.
-Note that if $p|\sqrt{\Delta}$, then $p\not|\;v=(q+1)(q^{2}-q+1)$, so we may apply Theorem 3.13.
-Thus, we let $p$ divide $v$.
-Now, $k-\theta_{1}=q^{3}-q^{2}+1$ and $k-\theta_{2}=(q+1)(q^{2}-q+1)$.
-We show that $v_{1}=k-\theta_{1}$ and $v_{2}=\frac{k-\theta_{2}}{\mu}$.
-So, suppose $p|\mu$ and $p|k-\theta_{1}$.
-Then mod $p$,
-0 0 $\equiv q^{3}-q^{2}+1$
+[…]
+$0$ $\equiv q^{3}-q^{2}+1$
 $\equiv q(q^{2}-q+1)-q+1$
@@
 $\equiv q^{2}$
-So $p|q$ and $p|1-q$, forcing $p=1$, so $\mu$ cannot divide $k-\theta_{1}$. Therefore, $v_{1}=q^{3}-q^{2}+1$ and $v_{2}=q+1$.
-[…]
+So $p|q$ and p|1qp|1-q, forcing $p=1$, so $\mu$ cannot divide $k-\theta_{1}$. Therefore, $v_{1}=q^{3}-q^{2}+1$ and $v_{2}=q+1$.
+Note that $v$ is odd, so if $G$ acts regularly on the points of this GQ, then $G$ is solvable by Feit-Thompson.
+Therefore, if there is some linear character of order $p$, then either $v_{1}\equiv 0$ and $v_{2}\equiv 1\pmod{p}$ or $v_{1}\equiv 1$ and $v_{2}\equiv 0\pmod{p}$ by Theorem 3.14.
+If $v_{1}\equiv 0$ and $v_{2}\equiv 1$, then $q\equiv 0$, forcing $v_{1}\equiv 1$, and thus $p=1$.
+If $v_{1}\equiv 1$ and $v_{2}\equiv 0$, then $q\equiv-1$, so $v_{1}\equiv(-1)^{3}-(-1)^{2}+1\equiv-1$, forcing $1\equiv 1$ and thus $p=2$.
+Therefore, $2|q+1$, and so $q$ is odd, a contradiction.
+Therefore, no such $G$ exists.
+∎
 Proposition 3.20.
-[…]
+If there exists a $(4w^{2}-1,2w^{2},w^{2},w^{2})$-PDS $D$ in a group $G$, then $w\equiv 1\pmod{3}$, all linear characters of $G$ evaluate to w-w on $D$, every conjugacy class intersection with $D$ is a multiple of $w$, and the order of the group of linear characters is a power of $3$.
 Proof.
-[…]
-Now, let $a\in G$ be an element of order divisible by 333 such that $a\not\in[G,G]$.
-Such an element must exist, for 333 divides $|\mathcal{L}|$.
+Note that $4w^{2}-1$ is odd, so $G$ is a solvable group.
+Moreover, $\sqrt{\Delta}=2w$, $\theta_{1}=w$, and $\theta_{2}=-w$, so $v_{1}=2w-1$ and $v_{2}=2w+1$.
+Since $G$ is solvable, it has a nontrivial group of linear characters, so let $p$ be a prime dividing the order of this group.
+Then, by Theorem 3.14, either $v_{1}\equiv 1\pmod{p}$ and $v_{2}\equiv 0\pmod{p}$, or $v_{1}\equiv 0\pmod{p}$ and $v_{2}\equiv 1\pmod{p}$.
+After some algebraic manipulation, it is clear that $v_{1}\equiv 0\pmod{p}$ and $v_{2}\equiv 1\pmod{p}$ implies $p=1$, and that $v_{1}\equiv 1\pmod{p}$ and $v_{2}\equiv 0\pmod{p}$ implies $p=3$.
+Therefore, $3$ divides $2w+1$, implying $w\equiv 1\pmod{3}$.
+Additionally, by Theorem 3.14, the linear group of characters must be a power of $3$.
+Now, let $a\in G$ be an element of order divisible by 33 such that $a\not\in[G,G]$.
+Such an element must exist, for 33 divides $|\mathcal{L}|$.
 In addition, let $\xi\in\mathcal{L}$.
@@
 $|C_{G}(a)|=\frac{k-(\pm w)}{|a^{G}\cap D|}=\frac{2w^{2}-(\pm w)}{|a^{G}\cap D|}=w\cdot\frac{2w-(\pm 1)}{|a^{G}\cap D|}=w\cdot\frac{6m+2-(\pm 1)}{|a^{G}\cap D|}.$
-Since $a$ has order divisible by 333, then 333 divides $|C_{G}(a)|$, so $6m+2-(\pm 1)\equiv 0\pmod{3}$.
+Since $a$ has order divisible by $3$, then $3$ divides $|C_{G}(a)|$, so $6m+2-(\pm 1)\equiv 0\pmod{3}$.
 Since $w\neq 0\pmod{3}$, then $2-(\pm 1)\equiv 0\pmod{3}$.
-It follows that $\xi(D)=-w$, so that 2(±1)2-($\pm$ 1)2 - ( ± 1 ) equals 333.
+It follows that $\xi(D)=-w$, so that $2-(\pm 1)$ equals 33.
 ∎
@@
 In some cases, simply knowing the values $|D\cap h^{G}|\pmod{p^{\ell}}$ can rule out the existence of a PDS for a certain parameter sets inside of a nonabelian group, such as in Example 4.2.
-In other cases, targeted information about $|D\cap h^{G}|$ can be used to quickly construct partial difference sets even inside of groups with up to 100010001000 elements, using the search method described in [2].
+In other cases, targeted information about $|D\cap h^{G}|$ can be used to quickly construct partial difference sets even inside of groups with up to $1000$ elements, using the search method described in [2].
 4.1. Modular computation of conjugacy class intersection sizes for PDSs
@@
 Lemma 4.1.
-For $h\in G$ such that $h\neq 1$, suppose that $p^{\ell}$ does not divide $|C_{G}(h)|$. Then, over the ring $R$,
+Using the notation above, for $h\in G$ such that $h\neq 1$, suppose that $p^{\ell}$ does not divide $|C_{G}(h)|$. Then, over the ring $R$,
 $|h^{G}\cap D||C_{G}(h)|\equiv k-\theta_{2}\pmod{(\pi^{\ell})}$,
@@
 Example 4.2.
-There does not exist an invertible (112,75,50)(112,75,50)( 112 , 75 , 50 )-DS in a nonabelian group $G$ of order 112112112.
-We suppose for contradiction that there does exist such a DS in a nonabelian group $G$ of order 112112112.
+There does not exist an invertible $(112,75,50)$-DS in a nonabelian group $G$ of order $112$.
+We suppose for contradiction that there does exist such a DS in a nonabelian group $G$ of order $112$.
 This parameter set has eigenvalues $\theta_{1}=5$ and $\theta_{2}=-5$.
@@
 Therefore, by Lemma 4.1, $\Phi(g)\equiv 75-5+\sum_{\chi\in{\rm Irr}(G)}(\chi(1)\theta_{2})\chi(g)\equiv 0\pmod{5}$.
-Finally, there is no centralizer $C_{G}(g)$ of order divisible by 555, since 555 does not divide 112112112.
+Finally, there is no centralizer $C_{G}(g)$ of order divisible by $5$, since $5$ does not divide $112$.
 Therefore, for all $g\in G$, $|g^{G}\cap D|\equiv 0\pmod{5}$.
-[…]
-An analogous computation for each group verifies that there are no nonabelian groups of order 112112112 containing a regular (112,75,50)(112,75,50)( 112 , 75 , 50 )-DS.
+Consulting GAP[15] on all possible nonabelian groups of order $112$, we may for a given group $G$ compute the largest multiple of $5$ less than or equal to the size of each conjugacy class of $G$.
+For example, if $G=$ SmallGroup(112, 41), then it has conjugacy class sizes $[1,1,8,7,8,7,8,8,8,8,8,8,8,8,8,8]$, and the largest multiple of 5 for each conjugacy class is $[0,0,5,5,5,5,5,5,5,5,5,5,5,5,5,5]$.
+In this case, $G$ cannot contain a reversible difference set, for Sum([0,0,5,5,5,5,5,5,5,5,5,5,5,5,5,5])=70<k=75[0,0,5,5,5,5,5,5,5,5,5,5,5,5,5,5])=70<k=75.
+An analogous computation for each group verifies that there are no nonabelian groups of order $112$ containing a regular $(112,75,50)$-DS.
 A direct verification may be done in GAP[15], using the following command:
@@
 ΨΨΨList(gps, g -> 5*Sum(List(ConjugacyClasses(g), i -> Int(Size(i)/5))));
-ΨΨ
+$Ψ$
 The output will be a list, and every entry of the list corresponds to the largest sum of elements in a possible PDS for a given group. Note that for every group, that sum is strictly less than $k=75$.
@@
 Lemma 4.3.
-Let $D$ be a regular PDS. If $p^{\ell}$ divides both $\theta_{1}$ and $\theta_{2}$, then $k\equiv 0\pmod{p^{\ell}}$. In particular, $k\equiv 0\pmod{\gcd(\theta_{1},\theta_{2})}$.
+Let $G$ contain a regular $(v,k,\lambda,\mu)$-PDS with eigenvalues $\theta_{1}>\theta_{2}$ such that $p^{\ell}$ divides both $\theta_{1}$ and $\theta_{2}$, then $k\equiv 0\pmod{p^{\ell}}$.
+In particular, $k\equiv 0\pmod{\gcd(\theta_{1},\theta_{2})}$.
 Proof.
@@
 Lemma 4.4.
-Let $D$ be a regular $PDS$ with eigenvalues $\theta_{1},\theta_{2}$ such that $\gcd(\theta_{1},\theta_{2})\neq 1$.
+Let $G$ be a group containing a regular $(v,k,\lambda,\mu)$-PDS $D$ with eigenvalues $\theta_{1}>\theta_{2}$ such that $\gcd(\theta_{1},\theta_{2})\neq 1$.
 If $p^{\ell}$ divides $\gcd(\theta_{1},\theta_{2})$ and $p^{\ell}$ does not divide $v$, then for each conjugacy class $h^{G}$,
@@
 Example 4.5.
-We explore the structure of a (183,112,66,72)(183,112,66,72)( 183 , 112 , 66 , 72 )-PDS $D$ in a group $G$ of order 183183183.
+We explore the structure of a $(183,112,66,72)$-PDS $D$ in a group $G$ of order $183$.
 The eigenvalues of this PDS are $\theta_{1}=4$ and $\theta_{2}=-10$.
-Then, $\gcd(4,-10)=2$, but 222 does not divide 183183183.
+Then, $\gcd(4,-10)=2$, but $2$ does not divide $183$.
 Therefore, by Lemma 4.4, for every $h\in G$, we have $|h^{G}\cap D|\equiv 0\pmod{2}$.
-By a computation in GAP[15], there is one nonabelian group given by SmallGroup(183, 1) with two conjugacy classes $h_{1}^{G},h_{2}^{G}$ of order 616161, 202020 conjugacy classes of order 333, and one conjugacy class of order 111.
+By a computation in GAP[15], there is one nonabelian group given by SmallGroup(183, 1) with two conjugacy classes $h_{1}^{G},h_{2}^{G}$ of order $61$, 2020 conjugacy classes of order $3$, and one conjugacy class of order $1$.
 If $x\in h_{1}^{G},h_{2}^{G}$, then $x\not\in G^{\prime}$ by a computation in GAP[15].
-Moreover, every prime $p$ dividing 183183183 does not divide 141414, and $G$ must have a subgroup of linear characters, for it is solvable by Feit-Thompson.
+Moreover, every prime $p$ dividing $183$ does not divide $14$, and $G$ must have a subgroup of linear characters, for it is solvable by Feit-Thompson.
 So, by Theorem 3.12, $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|\in\{\frac{112-4}{3},\frac{112+10}{3}\}$.
 Only $\frac{112-4}{3}=36$ is integral, so $|h_{1}^{G}\cap D|,|h_{2}^{G}\cap D|=36$.
-Therefore, 11272=40112-72=40112 - 72 = 40 elements must be partitioned among 202020 conjugacy classes.
-The maximum size of each conjugacy class is 333, and each class must have $0\pmod{2}$ elements, so each of the remaining conjugacy classes contain 222 elements of $D$.
+Therefore, 11272=40112-72=40 elements must be partitioned among $20$ conjugacy classes.
+The maximum size of each conjugacy class is $3$, and each class must have $0\pmod{2}$ elements, so each of the remaining conjugacy classes contain $2$ elements of $D$.
 A PDS does indeed exist for this parameter set, which we found through a modified form of the Random Hill Climb search discussed in [2]. We will speak on this more extensively in the next section.
 4.2. An algorithmic application of Lemma 4.1
-Combining Lemma 4.1 and Theorem 3.11 lets us build a fast algorithm to solve for all values $|D\cap h^{G}|$ even in comparatively large groups of order up to 100010001000.
+Combining Lemma 4.1 and Theorem 3.12 lets us build a fast algorithm to solve for all values $|D\cap h^{G}|$ even in comparatively large groups of order up to $1000$.
 We briefly describe the concept of such an algorithm, and then provide the pseudocode for an implementation.
@@
 Example 4.7.
-We show that there does not exist a (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 )-PDS inside of $G=$ SmallGroup(100, 3).
-By a computation in GAP[15], $G$ contains a conjugacy class $h^{G}$ of order 444 such that $g,g^{-1}\in h^{G}$ for all $g\in h^{G}$.
-Moreover, the centralizer $C_{G}(h)$ has order 252525.
-Since the eigenvalues corresponding to (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 ) are 888 and 3-3- 3, by Lemma 4.1,
+We show that there does not exist a $(100,33,14,9)$-PDS inside of $G=$ SmallGroup(100, 3).
+By a computation in GAP[15], $G$ contains a conjugacy class $h^{G}$ of order $4$ such that $g,g^{-1}\in h^{G}$ for all $g\in h^{G}$.
+Moreover, the centralizer $C_{G}(h)$ has order $25$.
+Since the eigenvalues corresponding to $(100,33,14,9)$ are $8$ and 3-3, by Lemma 4.1,
 $|h^{G}\cap D|\equiv(33-8)25^{-1}\equiv 1\pmod{11}$.
-The only nonnegative integer less than or equal to 444 that is congruent to $1\pmod{11}$ is 111, so $|h^{G}\cap D|=1$.
-Then, by Lemma 4.6, $G$ cannot contain a (100,33,14,9)(100,33,14,9)( 100 , 33 , 14 , 9 )-PDS.
+The only nonnegative integer less than or equal to $4$ that is congruent to $1\pmod{11}$ is $1$, so $|h^{G}\cap D|=1$. Then, by Lemma 4.6, $G$ cannot contain a $(100,33,14,9)$-PDS.
 Algorithm CCI Construct Class Intersections
@@
 27:  else
-[…]
+28:   DModIntersectionList[i] $\leftarrow[N,\text{Product(primeDivsDelta)}]$
 29:  end if
@@
 For example, Algorithm CCI returns the list
-[0,9,1,9,1,1,1,1,1,1,1,1,1,1,1][0,9,1,9,1,1,1,1,1,1,1,1,1,1,1][ 0 , 9 , 1 , 9 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 , 1 ]
-as a potentially valid list of conjugacy class intersection sizes for a (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 )-PDS in SmallGroup(111, 1).
+$[0,9,1,9,1,1,1,1,1,1,1,1,1,1,1]$
+as a potentially valid list of conjugacy class intersection sizes for a $(111,30,5,9)$-PDS in SmallGroup(111, 1).
 Note that a subset of this group with these intersection sizes will evaluate on irreducible characters to values admissible by Lemma 3.2.
-However, by turning the conjugacy class and PDS constraints into a linear program, we are able to rule out the existence of such a PDS using the linear optimization software Gurobi [19], and hence there does not exist a (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 )-PDS.
+However, by turning the conjugacy class and PDS constraints into a linear program, we are able to rule out the existence of such a PDS using the linear optimization software Gurobi [19], and hence there does not exist a $(111,30,5,9)$-PDS.
 Similarly, Algorithm CCI returns the lists
-[0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][ 0 , 4 , 14 , 1 , 1 , 4 , 4 , 1 , 2 , 4 ] , [ 0 , 4 , 14 , 1 , 1 , 4 , 4 , 2 , 1 , 4 ]
-as potentially valid lists of conjugacy class intersections for a (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 )-PDS in SmallGroup(96, 64) and SmallGroup(96, 227), respectively.
-These were also ruled out with the use of Gurobi [19], and, since these are the only valid lists of conjugacy class intersection sizes for these parameters, there does not exist a (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 )-PDS, either.
+[0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4][0,4,14,1,1,4,4,1,2,4],[0,4,14,1,1,4,4,2,1,4]
+as potentially valid lists of conjugacy class intersections for a $(96,35,10,14)$-PDS in SmallGroup(96, 64) and SmallGroup(96, 227), respectively.
+These were also ruled out with the use of Gurobi [19], and, since these are the only valid lists of conjugacy class intersection sizes for these parameters, there does not exist a $(96,35,10,14)$-PDS, either.
 The authors would be quite interested in a proof of the nonexistence of these PDSs that does not use the brute force of linear programming.
@@
 Using the technique from [2], we were able to construct nineteen distinct PDSs in nonabelian groups, many of which were already explicitly recognized, such as PDSs arising from triangular graphs.
-We found three parameter sets, (111,44,19,16)(111,44,19,16)( 111 , 44 , 19 , 16 ), (305,76,27,16)(305,76,27,16)( 305 , 76 , 27 , 16 ), and (981,140,43,16)(981,140,43,16)( 981 , 140 , 43 , 16 ), which correspond to an infinite family of groups acting regularly on Steiner systems ${\mathrm{S}}(2,4,p^{d})$ for $p$ prime.
+We found three parameter sets, $(111,44,19,16)$, $(305,76,27,16)$, and $(981,140,43,16)$, which correspond to an infinite family of groups acting regularly on Steiner systems ${\mathrm{S}}(2,4,p^{d})$ for $p$ prime.
 Five others that we were able to construct computationally – in SmallGroup(57, 1), SmallGroup(155, 1), SmallGroup(301, 1), SmallGroup(737, 1), and SmallGroup(1027, 1) – have parameters corresponding to an infinite family of PDSs corresponding to Steiner triple systems.
-The infinite families on ${\mathrm{S}}(2,4,p)$ and ${\mathrm{S}}(2,3,p^{d})$ correspond to block-regular Steiner 222-design constructions given by Clapham in [6] and Wilson in [33].
-Finally, we found a (183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 )-PDS in SmallGroup(183, 1) which is related to a ${\mathrm{S}}(2,5,61)$ Steiner system and, as it turns out, is also part of an infinite family.
+The infinite families on ${\mathrm{S}}(2,4,p)$ and ${\mathrm{S}}(2,3,p^{d})$ correspond to block-regular Steiner $2$-design constructions given by Clapham in [6] and Wilson in [33].
+Finally, we found a $(183,70,29,25)$-PDS in SmallGroup(183, 1) which is related to a ${\mathrm{S}}(2,5,61)$ Steiner system and, as it turns out, is also part of an infinite family.
 To the authors knowledge, the PDSs arising from [33] have not previously been recognized in the literature, and we are only aware of one paper [28] recognizing the constructions given by Clapham [6] as giving rise to PDSs.
@@
 Let $p$ be a prime such that $p^{d}>9$ and $p^{d}\equiv 7\pmod{12}$.
-Then, there exist $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$-PDSs in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
+Then, there exists a $(p^{d}(p^{d}-1)/6,3(p^{d}-3)/2,(p^{d}+3)/2,9)$-PDS in the nonabelian group $C_{p^{d}}\rtimes C_{(p^{d}-1)/6}\leqslant({\rm GF}(p^{d}),+)\rtimes({\rm GF}(p^{d})^{\times},\cdot)$.
 Proof.
@@
 Let $\alpha$ be the generator for $({\rm GF}(p^{d})^{\times},\cdot)$, and let $H=\langle\alpha^{(p^{d}-1)/k(k-1)}\rangle$.
-Let $B$ be a $k$ subset of $({\rm GF}(p^{d})^{\times},\cdot)$, and set =$\mathscr{B}=\{Bg:g\in H\}$.
+Let $B$ be a $k$ subset of $({\rm GF}(p^{d})^{\times},\cdot)$, and set $\mathscr{B}=\{Bg:g\in H\}$.
 Consider the set of nonzero differences $\delta B=\{b_{i}-b_{j}:b_{i},b_{j}\in B,b_{i}\neq b_{j}\}$ evaluated over the field ${\rm GF}(p^{d})$.
@@
 ∎
-When $p^{d}\leqslant(\frac{1}{2}k(k-1))^{k(k-1)}$, Buratti [5] provides sufficient conditions for block regular ${\mathrm{S}}(2,k,p^{d})$; we mention one result here, which implies the existence of the (183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 )-PDS we found in SmallGroup(183, 1) related to a block regular ${\mathrm{S}}(2,5,61)$.
+When $p^{d}\leqslant(\frac{1}{2}k(k-1))^{k(k-1)}$, Buratti [5] provides sufficient conditions for block regular ${\mathrm{S}}(2,k,p^{d})$; we mention one result here, which implies the existence of the $(183,70,29,25)$-PDS we found in SmallGroup(183, 1) related to a block regular ${\mathrm{S}}(2,5,61)$.
 Proposition 4.12.
-Let $p^{d}=20t+1$ be a prime power, let $2^{e}$ be the highest power of 222 in $t$, and suppose
+Let $p^{d}=20t+1$ be a prime power, let $2^{e}$ be the highest power of $2$ in $t$, and suppose
 $\frac{1}{2}(11+5\sqrt{5})\notin\{x^{2^{e+1}}:x\in GF(p^{d})^{\times}\}.$
@@
 By [5, Theorem 7], there always exists a block $B$ whose differences form a system of distinct representatives for the cosets of the group $H=\langle\alpha^{\frac{p-1}{20}}\rangle$.
-By [33, Theorem 2], the set =$\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,5,p)$.
+By [33, Theorem 2], the set $\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,5,p)$.
 By the discussion in Subsection 2.3, this gives a $(p^{d}(p^{d}-1)/20,5(p^{d}-5)/4,(p^{d}-1)/4+14,25)$-PDS.
@@
 Proposition 4.13.
-[…]
+If $p\equiv 13\pmod{2k(k-1)}$ and $p\neq 13$, then there exists a (p(p1)/12,4(p4)/3,(p+20)/3,16)(p(p-1)/12,4(p-4)/3,(p+20)/3,16)-PDS inside the nonabelian group $C_{p}\rtimes C_{(p-1)/12}\leqslant({\rm GF}(p),+)\rtimes({\rm GF}(p)^{\times},\cdot)$.
 Proof.
-[…]
+Let $\alpha$ generate the multiplicative group $({\rm GF}(p)^{\times},\cdot)$.
+By [14, Theorem 3.9], there always exists a block $B$ whose differences form a system of distinct representatives for the cosets of the group $H=\langle\alpha^{\frac{p-1}{12}}\rangle$.
+By [33, Theorem 2], the set $\mathscr{B}=\{Bg:g\in H\}$ forms the set of base blocks for a block regular ${\mathrm{S}}(2,4,p)$.
+By the discussion in Subsection 2.3, this gives a (p(p1)/12,4(p4)/3,(p+20)/3,16)(p(p-1)/12,4(p-4)/3,(p+20)/3,16)-PDS.
+∎
 4.4. Non-modular techniques
@@
 Example 4.14.
-We will prove that there is no reversible (144,66,30)(144,66,30)( 144 , 66 , 30 )-DS $D$ in SmallGroup(144, 68).
+We will prove that there is no reversible $(144,66,30)$-DS $D$ in SmallGroup(144, 68).
 In this case, $\sqrt{\Delta}=12$, where $12^{2}=144$.
 Therefore, Theorem 3.8, Theorem 3.12, Theorem 3.14, and Algorithm CCI are all inapplicable.
-Examining the character table for $G$ (which can be generated, for example, using GAP[15]), there are six conjugacy classes of noncentral elements of order 333.
-Each of these conjugacy classes has size 161616, so each element in one of these classes has a centralizer of order 999.
-Thus, for any noncentral element $x$ of order 333 in $G$, we have
+Examining the character table for $G$ (which can be generated, for example, using GAP[15]), there are six conjugacy classes of noncentral elements of order $3$.
+Each of these conjugacy classes has size $16$, so each element in one of these classes has a centralizer of order $9$.
+Thus, for any noncentral element $x$ of order $3$ in $G$, we have
 $\Phi(x)=9|x^{G}\cap D|\equiv 0\pmod{9}$.
 Moreover, if $\mathcal{L}$ denotes the set of linear characters for $G$, then $\chi\notin\mathcal{L}$ implies $\chi(x)=0$.
-Note that $|\mathcal{L}|=9$, and each nonprincipal character has order 333; other than the principal character, there are four conjugate pairs $\{\xi_{i},\overline{\xi_{i}}\},$ $1\leqslant i\leqslant 4$.
+Note that $|\mathcal{L}|=9$, and each nonprincipal character has order $3$; other than the principal character, there are four conjugate pairs $\{\xi_{i},\overline{\xi_{i}}\},$ $1\leqslant i\leqslant 4$.
 Extracting the principal character from this sum and noting that $D$ is reversible (so $\overline{\chi(D)}=\chi(D)$ by Lemma 3.4(i)), we have
 $\Phi(x)=66+\sum_{i=1}^{4}\xi_{i}(D)(\xi_{i}(x)+\overline{\xi_{i}}(x)).$
-Since each $\xi_{i}(D)$ is ±6$\pm$ 6± 6 by Lemma 3.2, there are only $2^{4}=16$ possibilities in total.
-Checking which have $\Phi(x)$ divisible by 999, there are only 555 possibilities that work.
+Since each $\xi_{i}(D)$ is $\pm 6$ by Lemma 3.2, there are only $2^{4}=16$ possibilities in total.
+Checking which have $\Phi(x)$ divisible by $9$, there are only $5$ possibilities that work.
 (In each case, $\Phi(x)=36$ or $\Phi(x)=72$.)
-Of the 151515 remaining nonlinear characters, there is one that equals its conjugate and seven conjugate pairs.
+Of the $15$ remaining nonlinear characters, there is one that equals its conjugate and seven conjugate pairs.
 Let $\chi$ be a nonlinear irreducible character.
 Note that $\theta_{1}=6$ and $\theta_{2}=-6$.
-By Lemma 3.2, if $\mathfrak{F}$ is the representation affording $\chi$, then $\chi(D)=6(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))$, where $-\chi(1)\leqslant(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))\leqslant\chi(1)$.
+By Lemma 3.2, if $\mathfrak{F}$ is the representation affording $\chi$, then $\chi(D)=6(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))$, where $\chi(1)\leqslant(\dim V_{\mathfrak{F}}(6)-\dim V_{\mathfrak{F}}(-6))\leqslant\chi(1)$.
 So, we may conclude that $\chi(D)\in\{-18,-6,6,18\}$.
@@
 Since either $1\in D$ or $1\notin D$, we have $\Phi(1)\in\{0,144\}$.
-Of the $5\cdot 4^{8}$ possibilities for $\Phi$, a GAP [15] calculation shows that only 158201582015820 of these possibilites have $\Phi(1)\in\{0,144\}$.
-Now, $Z(G)=\langle z\rangle$ is cyclic of order 333. Again, since $D$ is reversible, either $\{z,z^{-1}\}\subseteq D$ or $\{z,z^{-1}\}\cap D=\varnothing$. Thus, either
+Of the $5\cdot 4^{8}$ possibilities for $\Phi$, a GAP [15] calculation shows that only $15820$ of these possibilites have $\Phi(1)\in\{0,144\}$.
+Now, $Z(G)=\langle z\rangle$ is cyclic of order $3$. Again, since $D$ is reversible, either $\{z,z^{-1}\}\subseteq D$ or $\{z,z^{-1}\}\cap D=\varnothing$. Thus, either
 $\Phi(z)=\Phi(z^{-1})=0$
@@
 None of the remaining possibilities for $\Phi$ satisfy either of these constraints (in fact, all the values of these class functions are $6\pmod{12}$ when evaluated at $z$), and so no such $\Phi$ can exist.
-Therefore, there is no reversible (144,66,30)(144,66,30)( 144 , 66 , 30 )-DS (and hence no regular (144,66,30,30)(144,66,30,30)( 144 , 66 , 30 , 30 )-PDS) in $G$.
+Therefore, there is no reversible $(144,66,30)$-DS (and hence no regular $(144,66,30,30)$-PDS) in $G$.
 Acknowledgements
@@
 URL: https://www.wm.edu/it/rc
-Appendix A Parameters Ruled using Theorem 3.13
-The following tables include all parameters $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] that can be ruled out using Theorem 3.13 on either the original parameters or those of its complement, excluding those parameters with $v\in\{512,768,1024\}$, in which cases the computations are too expensive due to the number of groups on those orders.
+Appendix A Parameters already ruled out by [32]
+The following tables include all parameters $(v,k,\lambda,\mu)$ with k<(v1)/2k<(v-1)/2 listed in Brouwer’s tables [3] that can be ruled out (entirely, in any group) using either Theorem 3.14 or [32, Corollaries 5.2, 5.3] on either the original parameters or those of its complement, excluding those parameters with $v\in\{512,768,1024\}$, in which cases the computations are too expensive due to the number of groups on those orders.
 Note, we do not list both a parameter set and its complement.
@@
 175 72 20 36
+$v$ $k$ $\lambda$ $\mu$
 176 25 0 4
-$v$ $k$ $\lambda$ $\mu$
 176 45 18 9
@@
 249 88 27 33
+$v$ $k$ $\lambda$ $\mu$
 255 126 61 63
 259 42 5 7
-$v$ $k$ $\lambda$ $\mu$
 261 52 11 10
@@
 1107 378 117 135
-$1$ 110 9 11
+$11$ 110 9 11
 1112 396 135 144
@@
 Appendix B Parameters for regular PDSs ruled out using Algorithm CCI.
-The first set of tables tables consist of all parameters $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] satisfying the requirement that $\Delta$ is divisible by some prime $p$ not dividing $v$, for which we have shown no PDS exists through direct computation using Algorithm CCI on either the original parameters or those of the complement.
+The first set of tables tables consist of all parameters $(v,k,\lambda,\mu)$ with k<(v1)/2k<(v-1)/2 listed in Brouwer’s tables [3] satisfying the requirement that $\Delta$ is divisible by some prime $p$ not dividing $v$, for which we have shown no PDS exists in any group through direct computation using Algorithm CCI on either the original parameters or those of the complement.
 We have only included all parameters with $v<507$, for otherwise the computations become too expensive.
@@
 10 3 0 1
-26 10 3 4
-36 14 7 4
 40 12 2 4
@@
 70 27 12 9
-78 22 11 4
-82 36 15 16
 100 33 14 9
@@
 126 25 8 4
-126 50 13 24
-126 60 33 24
 130 48 20 16
 135 64 28 32
-136 30 15 4
 154 48 12 16
@@
 165 36 3 9
-170 78 35 36
 171 50 13 15
@@
 190 45 12 10
-$v$ $k$ $\lambda$ $\mu$
 190 84 33 40
@@
 196 39 2 9
-196 60 23 16
 196 81 42 27
@@
 208 45 8 10
+$v$ $k$ $\lambda$ $\mu$
 210 38 19 4
@@
 288 112 36 48
-$v$ $k$ $\lambda$ $\mu$
-290 136 63 64
 297 40 7 5
 297 104 31 39
-300 46 23 4
 306 55 4 11
@@
 322 96 20 32
-324 68 7 16
-324 95 34 25
 330 63 24 9
@@
 340 108 30 36
+$v$ $k$ $\lambda$ $\mu$
 342 33 4 3
@@
 364 165 68 80
-364 176 90 80
 372 56 10 8
@@
 375 182 85 91
-378 52 1 8
 378 52 26 4
@@
 405 196 91 98
-406 54 27 4
 406 108 30 28
@@
 430 165 68 60
-438 92 31 16
 441 56 7 7
@@
 441 220 95 124
-442 210 99 100
 456 65 10 9
@@
 470 126 27 36
+470 126 27 36
 474 165 52 60
@@
 484 105 14 25
-484 138 47 36
 490 165 56 55
@@
 506 100 18 20
-We have additionally computed all parameters $(v,k,\lambda,\mu)$ listed in Brouwer’s tables with $k<(v-1)/2$ and $507\leqslant v\leqslant 1300$ such that $\gcd(v,\Delta)=1$.
+We have additionally computed all parameters $(v,k,\lambda,\mu)$ listed in Brouwer’s tables with k<(v1)/2k<(v-1)/2 and $507\leqslant v\leqslant 1300$ such that $\gcd(v,\Delta)=1$.
 Of those, the following parameter sets are ruled out by our computations using Algorithm CCI.
@@
 610 87 32 9
-626 300 143 144
 630 68 1 8
@@
 658 162 51 36
-666 70 35 4
-676 150 23 36
 676 189 62 49
@@
 714 248 97 80
-$v$ $k$ $\lambda$ $\mu$
 726 145 44 25
@@
 742 255 92 85
+$v$ $k$ $\lambda$ $\mu$
 742 351 180 153
@@
 760 161 30 35
-776 124 39 16
 780 369 158 189
 782 99 36 9
-820 78 39 4
-842 406 195 196
 876 105 38 9
 900 203 34 49
-900 248 79 64
 904 301 78 111
-904 378 177 144
-$v$ $k$ $\lambda$ $\mu$
 946 243 60 63
@@
 1044 175 50 25
+1044 175 50 25
 1054 363 132 121
+$v$ $k$ $\lambda$ $\mu$
 1086 155 4 25
-1086 210 59 36
-1090 528 255 256
 1106 403 132 155
@@
 1155 576 286 288
-1156 264 47 64
 1156 315 98 81
-1176 94 47 4
 1178 99 0 9
@@
 1206 490 203 196
-1210 156 47 16
-1212 525 230 225
+$12$ 525 230 225
 1222 481 180 195
@@
 1254 203 52 29
-Finally, we mention again here that the parameter sets (111,30,5,9)(111,30,5,9)( 111 , 30 , 5 , 9 ) and (96,35,10,14)(96,35,10,14)( 96 , 35 , 10 , 14 ) have been ruled out using a combination of Algorithm CCI and Gurobi [19] (Remark 4.8).
+Finally, we mention again here that the parameter sets $(111,30,5,9)$ and $(96,35,10,14)$ have been ruled out using a combination of Algorithm CCI and Gurobi [19] (Remark 4.8).
 Appendix C Parameter sets with a valid output from Algorithm CCI
-We now list all parameter sets $(v,k,\lambda,\mu)$ with $k<(v-1)/2$ listed in Brouwer’s tables [3] satisfying the condition that either $\Delta$ has some distinct prime factor with $v$ and $v\leqslant 507$ or $\gcd(v,\Delta)=1$ and $v\leqslant 1300$, such that there does exist a valid $\Phi$ function, or such that this function was not computable with the available computational resources.
-We note once again that we exclude parameter sets with $v\in\{512,768,1024\}$, as well as the parameter set (507,176,70,56)(507,176,70,56)( 507 , 176 , 70 , 56 ), for these computations are too expensive.
+We now list all parameter sets $(v,k,\lambda,\mu)$ with k<(v1)/2k<(v-1)/2 listed in Brouwer’s tables [3] satisfying the condition that either $\Delta$ has some distinct prime factor with $v$ and $v\leqslant 507$ or $\gcd(v,\Delta)=1$ and $v\leqslant 1300$, such that there does exist a valid $\Phi$ function, or such that this function was not computable with the available computational resources.
+We note once again that we exclude parameter sets with $v\in\{512,768,1024\}$, as well as the parameter set $(507,176,70,56)$, for these computations are too expensive.
 As mentioned, we do not include complements of listed parameters – these may be easily generated from what is provided.
@@
 $(v,k,\lambda,\mu)$ ID
-(243, 112, 46, 56) 2
-(243, 112, 46, 56) 3
-(243, 112, 46, 56) 4
-(243, 112, 46, 56) 5
-(243, 112, 46, 56) 6
-(243, 112, 46, 56) 7
-(243, 112, 46, 56) 8
-(243, 112, 46, 56) 9
-(243, 112, 46, 56) 11
-(243, 112, 46, 56) 12
-(243, 112, 46, 56) 13
-(243, 112, 46, 56) 14
-(243, 112, 46, 56) 15
-(243, 112, 46, 56) 16
-(243, 112, 46, 56) 17
-(243, 112, 46, 56) 18
-(243, 112, 46, 56) 19
-(243, 112, 46, 56) 20
-(243, 112, 46, 56) 21
-(243, 112, 46, 56) 22
-(243, 112, 46, 56) 24
-(243, 112, 46, 56) 32
-(243, 112, 46, 56) 33
-(243, 112, 46, 56) 34
-(243, 112, 46, 56) 35
-$(v,k,\lambda,\mu)$ ID
-(243, 112, 46, 56) 36
-(243, 112, 46, 56) 37
-(243, 112, 46, 56) 38
-(243, 112, 46, 56) 39
-(243, 112, 46, 56) 40
-(243, 112, 46, 56) 41
-(243, 112, 46, 56) 42
-(243, 112, 46, 56) 43
-(243, 112, 46, 56) 44
-(243, 112, 46, 56) 45
-(243, 112, 46, 56) 46
-(243, 112, 46, 56) 47
-(243, 112, 46, 56) 49
-(243, 112, 46, 56) 50
-(243, 112, 46, 56) 51
-(243, 112, 46, 56) 52
-(243, 112, 46, 56) 53
-(243, 112, 46, 56) 54
-(243, 112, 46, 56) 55
-(243, 112, 46, 56) 56
-(243, 112, 46, 56) 57
-(243, 112, 46, 56) 58
-(243, 112, 46, 56) 59
-(243, 112, 46, 56) 60
-(243, 112, 46, 56) 62
-$(v,k,\lambda,\mu)$ ID
-(243, 112, 46, 56) 63
-(243, 112, 46, 56) 64
-(243, 112, 46, 56) 65
-(243, 112, 46, 56) 66
 $(343,162,81,72)$ 3
 $(343,162,81,72)$ 4
-(351, 110, 37, 33) 1
-(351, 110, 37, 33) 3
-(351, 110, 37, 33) 6
-(351, 110, 37, 33) 13
-(351, 126, 45, 45) 1
-(351, 126, 45, 45) 3
-(351, 126, 45, 45) 4
-(351, 126, 45, 45) 5
-(351, 126, 45, 45) 6
-(351, 126, 45, 45) 7
-(351, 126, 45, 45) 8
-(351, 126, 45, 45) 13
 $(375,176,94,72)$ 4
+$(v,k,\lambda,\mu)$ ID
 $(375,176,94,72)$ 5
@@
 $(392,153,54,63)$ 39
+$(v,k,\lambda,\mu)$ ID
 $(495,190,85,65)$ 1
 $(495,190,85,65)$ 3
-[…]
+We now provide all successful computations.
+We have organized the table as follows.
+First by the parameter set $(v,k,\lambda,\mu)$, then by the column labeled ID, which provides the GAP [15] group ID for which a valid $\Phi$ function has been constructed, then by the column labeled ?, which describes if a PDS for that group on that parameter set has been constructed (in which case a reference is given to its construction, either in this paper or elsewhere) or has not been either constructed or ruled out (?), in which the existence is still an open question.
+Finally, we provide the conjugacy class intersections of the possible PDS inside of the group.
+The intersections are given in the form of a list $a_{1},\dots,a_{n}$], such that, if $G$ is the group, then the $i$th entry of $a_{1},\dots,a_{n}$] corresponds to the intersection of the $i$th conjugacy class in ConjugacyClasses(CharacterTable($G$)) with the possible PDS.
+(Note that ConjugacyClasses($G$) and ConjugacyClasses(CharacterTable($G$)) may have different conjugacy class orderings.) We have listed all conjugacy class intersections up to the automorphisms of $G$.
+That is, if a conjugacy class intersection is omitted for a given group, then an automorphically equivalent intersection is listed.
+We also provide a full .txt file containing all of these parameters in this GitHub repository.
+We also mention that the parameter set $(243,112,46,56)$ has over 400 valid, automorphically distinct intersections.
+To avoid cluttering the paper, these have been placed in a separate file at the same GitHub repository located here.
 $(v,k,\lambda,\mu)$ ID ? Conjugacy Class Intersections
-(21,10,3,6)(21,10,3,6)( 21 , 10 , 3 , 6 ) 1 [27] [0, 3, 2, 3, 2]
-(27,10,1,5)(27,10,1,5)( 27 , 10 , 1 , 5 ) 3 [21] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
+$(21,10,3,6)$ 1 [27] [ 0, 3, 2, 3, 2 ]
+$(26,10,3,4)$ 1 ? [ 0, 4, 1, 1, 1, 1, 1, 1 ]
+$(27,10,1,5)$ 3 [21] [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 4 [30] [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
-(55,18,9,4)(55,18,9,4)( 55 , 18 , 9 , 4 ) 1 [27] [0, 4, 1, 4, 1, 4, 4]
-(57,24,11,9)(57,24,11,9)( 57 , 24 , 11 , 9 ) 1 4.9 [0, 9, 1, 9, 1, 1, 1, 1, 1]
-(111,44,19,16)(111,44,19,16)( 111 , 44 , 19 , 16 ) 1 4.13 [0, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(125,28,3,7)(125,28,3,7)( 125 , 28 , 3 , 7 ) 3 [21] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
+$(36,14,7,4)$ 9 ? [ 0, 4, 4, 1, 1, 4 ]
+$(55,18,9,4)$ 1 [27] [ 0, 4, 1, 4, 1, 4, 4 ]
+$(57,24,11,9)$ 1 4.9 [ 0, 9, 1, 9, 1, 1, 1, 1, 1 ]
+$(78,22,11,4)$ 1 ? [ 0, 4, 4, 1, 4, 4, 1, 4 ]
+$(82,36,15,16)$ 1 ? [ 0, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(111,44,19,16)$ 1 4.13 [ 0, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(125,28,3,7)$ 3 [21] [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 4 [30] [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
-(125,52,15,26)(125,52,15,26)( 125 , 52 , 15 , 26 ) 3 [27] [0, 2, 2, 1, 2, 2, 2, 1, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
+$(125,52,15,26)$ 3 [27] [ 0, 2, 2, 1, 2, 2, 2, 1, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2 ]
 4 ? [ 0, 2, 2, 1, 2, 2, 2, 1, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 1, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2 ]
-(136,30,8,6)(136,30,8,6)( 136 , 30 , 8 , 6 ) 12 ? [0, 3, 3, 8, 2, 3, 3, 3, 3, 2]
-(136,60,24,28)(136,60,24,28)( 136 , 60 , 24 , 28 ) 12 ? [0, 7, 10, 4, 4, 7, 7, 10, 7, 4]
-(136,63,30,28)(136,63,30,28)( 136 , 63 , 30 , 28 ) 12 ? [0, 7, 7, 13, 4, 7, 7, 7, 7, 4]
-(147,66,25,33)(147,66,25,33)( 147 , 66 , 25 , 33 ) 3 [2] [0, 3, 1, 3, 3, 3, 1, 1, 3, 3, 1, 1, 3, 3, 3, 1, 1, 1, 3, 3, 1, 1, 1, 3, 3, 1, 1, 1, 3, 3, 1, 1, 3, 1, 1]
+$(126,50,13,24)$ 9 ? [ 0, 8, 1, 1, 3, 8, 1, 7, 3, 8, 7, 3 ]
+$(126,60,33,24)$ 9 ? [ 0, 8, 6, 1, 3, 8, 6, 7, 3, 8, 7, 3 ]
+$(136,30,8,6)$ 12 ? [ 0, 3, 3, 8, 2, 3, 3, 3, 3, 2 ]
+$(136,30,15,4)$ 12 ? [ 0, 4, 4, 4, 1, 4, 4, 4, 4, 1 ]
+$(136,60,24,28)$ 12 ? [ 0, 7, 10, 4, 4, 7, 7, 10, 7, 4 ]
+$(136,63,30,28)$ 12 ? [ 0, 7, 7, 13, 4, 7, 7, 7, 7, 4 ]
+$(147,66,25,33)$ 3 [2] [ 0, 3, 1, 3, 3, 3, 1, 1, 3, 3, 1, 1, 3, 3, 3, 1, 1, 1, 3, 3, 1, 1, 1, 3, 3, 1, 1, 1, 3, 3, 1, 1, 3, 1, 1 ]
 4 [2] [ 0, 21, 3, 3, 21, 1, 3, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 5 [2] [ 0, 21, 3, 3, 21, 1, 3, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
-(148,63,22,30)(148,63,22,30)( 148 , 63 , 22 , 30 ) 3 ? [0, 15, 15, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2]
-(148,70,36,30)(148,70,36,30)( 148 , 70 , 36 , 30 ) 3 ? [0, 15, 22, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2]
-(155,42,17,9)(155,42,17,9)( 155 , 42 , 17 , 9 ) 1 4.9 [0, 9, 1, 9, 9, 1, 9, 1, 1, 1, 1]
-(160,54,18,18)(160,54,18,18)( 160 , 54 , 18 , 18 ) 199 ? [0, 6, 6, 0, 0, 0, 6, 6, 6, 6, 6, 6, 6]
+$(148,63,22,30)$ 3 ? [ 0, 15, 15, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2 ]
+$(148,70,36,30)$ 3 ? [ 0, 15, 8, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2 ]
+3 ? [ 0, 15, 22, 2, 15, 2, 2, 2, 2, 2, 2, 2, 2 ]
+$(155,42,17,9)$ 1 4.9 [ 0, 9, 1, 9, 9, 1, 9, 1, 1, 1, 1 ]
+$(160,54,18,18)$ 199 ? [ 0, 6, 0, 0, 6, 0, 6, 6, 6, 6, 6, 6, 6 ]
+199 ? [ 0, 6, 6, 0, 0, 0, 6, 6, 6, 6, 6, 6, 6 ]
 199 ? [ 0, 6, 0, 3, 3, 0, 6, 6, 6, 6, 6, 6, 6 ]
 234 ? [ 0, 6, 12, 3, 3, 0, 6, 6, 6, 12 ]
-(171,34,17,4)(171,34,17,4)( 171 , 34 , 17 , 4 ) 3 [27] [0, 4, 4, 1, 4, 4, 4, 1, 4, 4, 4]
-(171,60,15,24)(171,60,15,24)( 171 , 60 , 15 , 24 ) 3 ? [0, 8, 3, 3, 8, 8, 3, 3, 8, 8, 8]
-(183,52,11,16)(183,52,11,16)( 183 , 52 , 11 , 16 ) 1 ? [0, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(183,70,29,25)(183,70,29,25)( 183 , 70 , 29 , 25 ) 1 4.12 [0, 25, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(205,68,15,26)(205,68,15,26)( 205 , 68 , 15 , 26 ) 1 ? [0, 13, 2, 13, 2, 13, 2, 13, 2, 2, 2, 2, 2]
-(205,96,50,40)(205,96,50,40)( 205 , 96 , 50 , 40 ) 1 ? [0, 20, 2, 20, 2, 20, 2, 20, 2, 2, 2, 2, 2]
-(216,90,39,36)(216,90,39,36)( 216 , 90 , 39 , 36 ) 86 ? [0, 12, 4, 4, 4, 2, 12, 12, 4, 8, 8, 12, 8]
+$(170,78,35,36)$ 3 ? [ 0, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(171,34,17,4)$ 3 [27] [ 0, 4, 4, 1, 4, 4, 4, 1, 4, 4, 4 ]
+$(171,60,15,24)$ 3 ? [ 0, 8, 3, 3, 8, 8, 3, 3, 8, 8, 8 ]
+$(183,52,11,16)$ 1 ? [ 0, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(183,70,29,25)$ 1 4.12 [ 0, 25, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(196,60,23,16)$ 8 ? [ 0, 16, 16, 1, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(205,68,15,26)$ 1 ? [ 0, 13, 2, 13, 2, 13, 2, 13, 2, 2, 2, 2, 2 ]
+$(205,96,50,40)$ 1 ? [ 0, 20, 2, 20, 2, 20, 2, 20, 2, 2, 2, 2, 2 ]
+$(216,90,39,36)$ 86 ? [ 0, 12, 4, 4, 4, 2, 12, 12, 4, 8, 8, 12, 8 ]
 87 ? [ 0, 8, 8, 4, 2, 2, 8, 16, 16, 8, 2, 8, 8 ]
@@
 153 ? [ 0, 2, 24, 4, 8, 2, 16, 9, 16, 9 ]
-(243,112,46,56)(243,112,46,56)( 243 , 112 , 46 , 56 ) 28 ? [0, 18, 4, 0, 2, 0, 18, 12, 4, 4, 0, 2, 0, 12, 12, 4, 4, 12, 4]
-28 ? [0, 12, 4, 0, 2, 0, 12, 18, 4, 4, 0, 2, 0, 12, 12, 4, 4, 18, 4]
-28 ? [0, 12, 4, 6, 2, 0, 12, 12, 4, 4, 6, 2, 0, 12, 12, 4, 4, 12, 4]
-28 ? [0, 12, 6, 0, 2, 0, 12, 12, 6, 6, 0, 2, 0, 12, 12, 6, 6, 12, 6]
-29 ? [0, 18, 4, 0, 2, 0, 18, 12, 4, 4, 0, 2, 0, 12, 12, 4, 4, 12, 4]
-29 ? [0, 12, 4, 0, 2, 0, 12, 18, 4, 4, 0, 2, 0, 12, 12, 4, 4, 18, 4]
-29 ? [0, 12, 4, 6, 2, 0, 12, 12, 4, 4, 6, 2, 0, 12, 12, 4, 4, 12, 4]
-29 ? [0, 12, 6, 0, 2, 0, 12, 12, 6, 6, 0, 2, 0, 12, 12, 6, 6, 12, 6]
-30 ? [0, 18, 4, 0, 2, 0, 18, 12, 4, 4, 0, 2, 0, 12, 12, 4, 4, 12, 4]
-30 ? [0, 12, 4, 0, 2, 0, 12, 18, 4, 4, 0, 2, 0, 12, 12, 4, 4, 18, 4]
-30 ? [0, 12, 4, 6, 2, 0, 12, 12, 4, 4, 6, 2, 0, 12, 12, 4, 4, 12, 4]
-30 ? [0, 12, 6, 0, 2, 0, 12, 12, 6, 6, 0, 2, 0, 12, 12, 6, 6, 12, 6]
-(244,108,42,52)(244,108,42,52)( 244 , 108 , 42 , 52 ) 3 ? [0, 26, 26, 2, 26, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
-(244,117,60,52)(244,117,60,52)( 244 , 117 , 60 , 52 ) 3 ? [0, 26, 35, 2, 26, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2]
-(250,81,24,27)(250,81,24,27)( 250 , 81 , 24 , 27 ) 5 ? [0, 9, 3, 6, 0, 9, 3, 3, 6, 0, 9, 3, 0, 0, 9, 3, 0, 3, 9, 3, 3, 0]
+$(244,108,42,52)$ 3 ? [ 0, 26, 26, 2, 26, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2 ]
+$(244,117,60,52)$ 3 ? [ 0, 26, 35, 2, 26, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2 ]
+$(250,81,24,27)$ 5 ? [ 0, 9, 3, 0, 0, 9, 3, 3, 0, 0, 9, 3, 3, 3, 9, 3, 3, 3, 9, 3, 3, 3 ]
 5 ? [ 0, 9, 3, 0, 0, 9, 3, 0, 0, 0, 9, 3, 6, 6, 9, 3, 6, 0, 9, 0, 0, 6 ]
@@
 8 ? [ 0, 9, 6, 6, 0, 9, 6, 6, 6, 0, 9, 0, 0, 0, 9, 9, 0, 6, 0, 0, 0, 0 ]
-(253,42,21,4)(253,42,21,4)( 253 , 42 , 21 , 4 ) 1 [27] [0, 4, 1, 4, 4, 4, 4, 1, 4, 4, 4, 4, 4]
-(253,112,36,60)(253,112,36,60)( 253 , 112 , 36 , 60 ) 1 [4] [0, 10, 6, 10, 10, 10, 10, 6, 10, 10, 10, 10, 10]
-(273,80,19,25)(273,80,19,25)( 273 , 80 , 19 , 25 ) 3 ? [0, 25, 1, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
+$(253,42,21,4)$ 1 [27] [ 0, 4, 1, 4, 4, 4, 4, 1, 4, 4, 4, 4, 4 ]
+$(253,112,36,60)$ 1 [4] [ 0, 10, 6, 10, 10, 10, 10, 6, 10, 10, 10, 10, 10 ]
+$(273,80,19,25)$ 3 ? [ 0, 25, 1, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 4 ? [ 0, 25, 1, 1, 25, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
-(273,102,41,36)(273,102,41,36)( 273 , 102 , 41 , 36 ) 3 ? [0, 36, 1, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
+$(273,102,41,36)$ 3 ? [ 0, 36, 1, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 4 ? [ 0, 36, 1, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
-(301,60,23,9)(301,60,23,9)( 301 , 60 , 23 , 9 ) 1 4.9 [0, 9, 1, 9, 1, 9, 1, 9, 9, 9, 1, 1, 1]
-(301,108,27,45)(301,108,27,45)( 301 , 108 , 27 , 45 ) 1 ? [0, 15, 3, 15, 3, 15, 3, 15, 15, 15, 3, 3, 3]
-(301,150,65,84)(301,150,65,84)( 301 , 150 , 65 , 84 ) 1 ? [0, 21, 4, 21, 4, 21, 4, 21, 21, 21, 4, 4, 4]
-(305,76,27,16)(305,76,27,16)( 305 , 76 , 27 , 16 ) 1 4.13 [0, 16, 1, 16, 1, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(320,99,18,36)(320,99,18,36)( 320 , 99 , 18 , 36 ) 1635 ? [0, 12, 18, 24, 3, 0, 12, 12, 0, 6, 12]
+$(290,136,63,64)$ 3 ?
+[ 0, 64, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(300,46,23,4)$ 23 ? [ 0, 12, 4, 8, 1, 1, 12, 8 ]
+24 ? [ 0, 4, 4, 4, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(301,60,23,9)$ 1 4.9 [ 0, 9, 1, 9, 1, 9, 1, 9, 9, 9, 1, 1, 1 ]
+$(301,108,27,45)$ 1 ? [ 0, 15, 3, 15, 3, 15, 3, 15, 15, 15, 3, 3, 3 ]
+$(301,150,65,84)$ 1 ? [ 0, 21, 4, 21, 4, 21, 4, 21, 21, 21, 4, 4, 4 ]
+$(305,76,27,16)$ 1 4.13 [ 0, 16, 1, 16, 1, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(320,99,18,36)$ 1635 ? [ 0, 12, 6, 24, 3, 0, 12, 12, 0, 18, 12 ]
 1635 ? [ 0, 12, 6, 24, 3, 0, 12, 12, 12, 6, 12 ]
@@
 1635 ? [ 0, 12, 18, 24, 0, 3, 12, 12, 6, 0, 12 ]
-(320,132,46,60)(320,132,46,60)( 320 , 132 , 46 , 60 ) 1635 ? [0, 16, 8, 30, 2, 4, 16, 16, 16, 8, 16]
-(343,54,5,9)(343,54,5,9)( 343 , 54 , 5 , 9 ) 3 [21] [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
+$(320,132,46,60)$ 1635 ? [ 0, 16, 8, 30, 2, 4, 16, 16, 16, 8, 16 ]
+1635 ? [ 0, 16, 8, 30, 2, 4, 16, 16, 5, 19, 16 ]
+$(324,68,7,16)$ 35 ? [ 0, 16, 16, 1, 1, 1, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+164 ? [ 0, 16, 16, 1, 1, 1, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(324,95,34,25)$ 160 ? [ 0, 9, 2, 11, 9, 9, 23, 3, 9, 9, 1, 9, 1 ]
+$(343,54,5,9)$ 3 [21] [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 4 [30] [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
-(343,150,53,75)(343,150,53,75)( 343 , 150 , 53 , 75 ) 3 [27] [0, 3, 3, 1, 3, 3, 3, 1, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
+$(343,150,53,75)$ 3 [27] [ 0, 3, 3, 1, 3, 3, 3, 1, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 ]
 4 ? [ 0, 3, 3, 1, 3, 3, 3, 1, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 1, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 ]
-(351,50,13,6)(351,50,13,6)( 351 , 50 , 13 , 6 ) 12 ? [0, 3, 7, 7, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
-(351,50,25,4)(351,50,25,4)( 351 , 50 , 25 , 4 ) 12 [27] [0, 4, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4]
-(351,110,37,33)(351,110,37,33)( 351 , 110 , 37 , 33 ) 4 ?
-[0, 15, 3, 1, 1, 15, 9, 3, 1, 1, 1, 1, 9, 9, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
+$(351,50,13,6)$ 12 ? [ 0, 3, 7, 7, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3 ]
+$(351,50,25,4)$ 12 [27] [ 0, 4, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(351,110,37,33)$ 3 ?
+[ 0, 9, 1, 1, 1, 9, 3, 3, 1, 1, 1, 1, 1, 1, 3, 3, 3, 3, 3, 1, 1, 1, 1, 1, 1, 1, 3, 3, 3, 3, 3, 1, 1, 1, 1, 1, 1, 1, 3, 3, 3, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+3 ? [ 0, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 5, 7, 5, 3, 7, 1, 1, 1, 1, 1, 1, 1, 3, 5, 1, 5, 3, 1, 1, 1, 1, 1, 1, 1, 3, 5, 5, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+4 ? [ 0, 15, 3, 1, 1, 15, 9, 3, 1, 1, 1, 1, 9, 9, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
 4 ? [ 0, 9, 3, 1, 1, 9, 9, 3, 1, 1, 1, 1, 15, 15, 1, 1, 1, 1, 1, 9, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
@@
 12 ? [ 0, 9, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9 ]
-(351,112,43,32)(351,112,43,32)( 351 , 112 , 43 , 32 ) 12 ? [0, 9, 2, 2, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
-(351,126,45,45)(351,126,45,45)( 351 , 126 , 45 , 45 ) 12 ? [0, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
-(351,140,73,44)(351,140,73,44)( 351 , 140 , 73 , 44 ) 12 ? [0, 11, 4, 4, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11]
-(363,170,73,85)(363,170,73,85)( 363 , 170 , 73 , 85 ) 2 ? [0, 55, 3, 1, 55, 3, 1, 1, 3, 1, 1, 1, 3, 1, 3, 1, 3, 3, 1, 3, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(381,114,29,36)(381,114,29,36)( 381 , 114 , 29 , 36 ) 1 ? [0, 36, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(381,140,55,49)(381,140,55,49)( 381 , 140 , 55 , 49 ) 1 ? [0, 49, 1, 49, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
-(400,156,74,52)(400,156,74,52)( 400 , 156 , 74 , 52 ) 206 ? [0, 20, 20, 10, 10, 2, 20, 4, 20, 10, 4, 20, 16]
+13 ? [ 0, 9, 1, 1, 1, 9, 3, 3, 1, 1, 1, 1, 1, 1, 3, 3, 3, 3, 3, 1, 1, 1, 1, 1, 1, 1, 3, 3, 3, 3, 3, 1, 1, 1, 1, 1, 1, 1, 3, 3, 3, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+13 ? [ 0, 1, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 7, 5, 7, 3, 5, 1, 1, 1, 1, 1, 1, 1, 1, 5, 3, 3, 5, 1, 1, 1, 1, 1, 1, 1, 5, 3, 5, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(351,112,43,32)$ 12 ? [ 0, 9, 2, 2, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9 ]
+$(351,126,45,45)$ 12 ? [ 0, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9 ]
+$(351,140,73,44)$ 12 ? [ 0, 11, 4, 4, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11, 11 ]
+$(363,170,73,85)$ 2 ? [ 0, 55, 3, 1, 55, 3, 1, 1, 3, 1, 1, 1, 3, 1, 3, 1, 3, 3, 1, 3, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 3, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(364,176,90,80)$ 6 ? [ 0, 40, 12, 1, 2, 40, 13, 1, 2, 2, 13, 1, 2, 2, 13, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2 ]
+$(378,52,1,8)$ 18 ? [ 0, 8, 1, 1, 1, 2, 8, 1, 7, 1, 1, 8, 7, 1, 1, 1, 1, 1, 1 ]
+19 ? [ 0, 8, 1, 1, 1, 2, 8, 1, 7, 1, 1, 8, 7, 1, 1, 1, 1, 1, 1 ]
+22 ? [ 0, 8, 1, 1, 2, 1, 8, 1, 7, 1, 1, 8, 7, 1, 1, 1, 1, 1, 1 ]
+$(381,114,29,36)$ 1 ? [ 0, 36, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(381,140,55,49)$ 1 ? [ 0, 49, 1, 49, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(400,156,74,52)$ 206 ? [ 0, 20, 20, 10, 10, 2, 20, 4, 20, 10, 4, 20, 16 ]
 207 ? [ 0, 10, 4, 4, 10, 2, 20, 20, 10, 20, 16, 16, 2, 4, 2, 16 ]
-(405,96,18,24)(405,96,18,24)( 405 , 96 , 18 , 24 ) 15 ? [0, 18, 4, 4, 0, 2, 18, 4, 2, 0, 0, 4, 0, 2, 0, 18, 0, 0, 2, 0, 18]
+$(405,96,18,24)$ 15 ? [ 0, 18, 4, 4, 0, 2, 18, 4, 2, 0, 0, 4, 0, 2, 0, 18, 0, 0, 2, 0, 18 ]
 15 ? [ 0, 18, 0, 0, 0, 0, 18, 0, 0, 0, 4, 0, 4, 0, 4, 18, 4, 4, 0, 4, 18 ]
@@
 15 ? [ 0, 18, 0, 0, 2, 0, 18, 0, 2, 2, 4, 0, 4, 0, 4, 18, 0, 0, 2, 4, 18 ]
-(441,88,35,13)(441,88,35,13)( 441 , 88 , 35 , 13 ) 9 ? [0, 1, 1, 1, 1, 1, 7, 3, 1, 3, 7, 7, 3, 7, 3, 1, 1, 7, 3, 3, 7, 7, 3, 3, 7]
-(441,152,43,57)(441,152,43,57)( 441 , 152 , 43 , 57 ) 9 ? [0, 5, 5, 1, 1, 5, 19, 7, 5, 7, 3, 11, 7, 11, 7, 1, 1, 19, 7, 7, 3, 3, 7, 7, 3]
+$(406,54,27,4)$ 1 ? [ 0, 4, 4, 1, 4, 4, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(438,92,31,16)$ 1 ? [ 0, 16, 16, 1, 16, 16, 1, 16, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(441,88,35,13)$ 9 ? [ 0, 1, 1, 1, 1, 1, 7, 3, 1, 3, 7, 7, 3, 7, 3, 1, 1, 7, 3, 3, 7, 7, 3, 3, 7 ]
+$(441,152,43,57)$ 9 ? [ 0, 5, 5, 1, 1, 5, 19, 7, 5, 7, 3, 11, 7, 11, 7, 1, 1, 19, 7, 7, 3, 3, 7, 7, 3 ]
 9 ? [ 0, 5, 5, 1, 1, 5, 11, 7, 5, 7, 3, 19, 7, 19, 7, 1, 1, 11, 7, 7, 3, 3, 7, 7, 3 ]
-(441,176,85,60)(441,176,85,60)( 441 , 176 , 85 , 60 ) 9 ? [0, 6, 6, 1, 1, 6, 20, 7, 6, 7, 3, 20, 7, 20, 7, 1, 1, 20, 7, 7, 3, 3, 7, 7, 3]
-(441,184,87,69)(441,184,87,69)( 441 , 184 , 87 , 69 ) 9 ? [0, 3, 3, 3, 3, 3, 21, 9, 3, 9, 1, 21, 9, 21, 9, 3, 3, 21, 9, 9, 1, 1, 9, 9, 1]
-(448,150,50,50)(448,150,50,50)( 448 , 150 , 50 , 50 ) 178 ? [0, 20, 5, 0, 20, 5, 5, 0, 5, 0, 5, 5, 20, 20, 20, 20]
+$(441,176,85,60)$ 9 ? [ 0, 6, 6, 1, 1, 6, 20, 7, 6, 7, 3, 20, 7, 20, 7, 1, 1, 20, 7, 7, 3, 3, 7, 7, 3 ]
+$(441,184,87,69)$ 9 ? [ 0, 3, 3, 3, 3, 3, 21, 9, 3, 9, 1, 21, 9, 21, 9, 3, 3, 21, 9, 9, 1, 1, 9, 9, 1 ]
+$(442,210,99,100)$ 3 ?
+[ 0, 100, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(448,150,50,50)$ 178 ? [ 0, 20, 5, 0, 20, 5, 5, 0, 5, 0, 5, 5, 20, 20, 20, 20 ]
 179 ? [ 0, 20, 15, 0, 20, 15, 20, 20, 20, 20 ]
@@
 1394 ? [ 0, 20, 5, 5, 20, 0, 0, 5, 0, 5, 5, 5, 20, 20, 20, 20 ]
-(448,162,66,54)(448,162,66,54)( 448 , 162 , 66 , 54 ) 178 ? [0, 24, 6, 6, 24, 0, 0, 0, 0, 0, 6, 0, 24, 24, 24, 24]
+$(448,162,66,54)$ 178 ? [ 0, 24, 6, 6, 24, 0, 0, 0, 0, 0, 6, 0, 24, 24, 24, 24 ]
 178 ? [ 0, 24, 0, 6, 24, 0, 3, 3, 3, 3, 0, 0, 24, 24, 24, 24 ]
@@
 1394 ? [ 0, 24, 0, 0, 24, 0, 0, 0, 6, 0, 6, 6, 24, 24, 24, 24 ]
-(465,58,29,4)(465,58,29,4)( 465 , 58 , 29 , 4 ) 1 [27] [ 0, 4, 4, 1, 4, 4, 4, 4, 4, 4, 1, 4, 4, 4, 4, 4, 4 ]
-(465,192,72,84)(465,192,72,84)( 465 , 192 , 72 , 84 ) 1 ? [ 0, 6, 14, 6, 6, 14, 14, 14, 14, 14, 6, 14, 14, 14, 14, 14, 14 ]
-(497,186,55,78)(497,186,55,78)( 497 , 186 , 55 , 78 ) 1 ? [ 0, 26, 3, 26, 3, 26, 26, 3, 26, 3, 26, 3, 3, 3, 3, 3, 3 ]
-(497,240,127,105)(497,240,127,105)( 497 , 240 , 127 , 105 ) 1 ? [ 0, 35, 3, 35, 3, 35, 35, 3, 35, 3, 35, 3, 3, 3, 3, 3, 3 ]
-(505,84,3,16)(505,84,3,16)( 505 , 84 , 3 , 16 ) 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16 ]
-(505,120,39,25)(505,120,39,25)( 505 , 120 , 39 , 25 ) 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 25, 25, 25, 25 ]
-(505,180,53,70)(505,180,53,70)( 505 , 180 , 53 , 70 ) 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 35, 35, 35, 35 ]
-(505,224,108,92)(505,224,108,92)( 505 , 224 , 108 , 92 ) 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 46, 46, 46, 46 ]
-(507,184,71,64)(507,184,71,64)( 507 , 184 , 71 , 64 ) 1 ?
+$(465,58,29,4)$ 1 [27] [ 0, 4, 4, 1, 4, 4, 4, 4, 4, 4, 1, 4, 4, 4, 4, 4, 4 ]
+$(465,192,72,84)$ 1 ? [ 0, 6, 14, 6, 6, 14, 14, 14, 14, 14, 6, 14, 14, 14, 14, 14, 14 ]
+$(484,138,47,36)$ 8 ? [ 0, 36, 36, 1, 1, 36, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 ]
+$(497,186,55,78)$ 1 ? [ 0, 26, 3, 26, 3, 26, 26, 3, 26, 3, 26, 3, 3, 3, 3, 3, 3 ]
+$(497,240,127,105)$ 1 ? [ 0, 35, 3, 35, 3, 35, 35, 3, 35, 3, 35, 3, 3, 3, 3, 3, 3 ]
+$(505,84,3,16)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16 ]
+$(505,120,39,25)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 25, 25, 25, 25 ]
+$(505,180,53,70)$ 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 35, 35, 35, 35 ]
+$(505,224,108,92)$ 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 46, 46, 46, 46 ]
+$(507,184,71,64)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64 ]
@@
 5 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64 ]
-(651,200,55,64)(651,200,55,64)( 651 , 200 , 55 , 64 ) 3 ?
+$(626,300,143,144)$ 1 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 144 ]
+$(651,200,55,64)$ 3 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64 ]
@@
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64 ]
-(651,234,89,81)(651,234,89,81)( 651 , 234 , 89 , 81 ) 3 ?
+$(651,234,89,81)$ 3 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 81, 81 ]
@@
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 81, 81 ]
-(657,256,80,112)(657,256,80,112)( 657 , 256 , 80 , 112 ) 3 ? [0, 4, 4, 4, 4, 4, 4, 4, 4, 28, 28, 28, 28, 28, 28, 28, 28]
-(657,328,147,180)(657,328,147,180)( 657 , 328 , 147 , 180 ) 3 ? [0, 5, 5, 5, 5, 5, 5, 5, 5, 36, 36, 36, 36, 36, 36, 36, 36]
-(689,256,120,80)(689,256,120,80)( 689 , 256 , 120 , 80 ) 1 ? [0, 4, 4, 4, 4, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20]
-(737,96,35,9)(737,96,35,9)( 737 , 96 , 35 , 9 ) 1 4.9 [0, 1, 1, 1, 1, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
-(755,130,9,25)(755,130,9,25)( 755 , 130 , 9 , 25 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 25, 25, 25, 25]
-(755,174,53,36)(755,174,53,36)( 755 , 174 , 53 , 36 ) 1 ? [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36]
-(813,252,71,81)(813,252,71,81)( 813 , 252 , 71 , 81 ) 1 ?
+$(657,256,80,112)$ 3 ? [ 0, 4, 4, 4, 4, 4, 4, 4, 4, 28, 28, 28, 28, 28, 28, 28, 28 ]
+$(657,328,147,180)$ 3 ? [ 0, 5, 5, 5, 5, 5, 5, 5, 5, 36, 36, 36, 36, 36, 36, 36, 36 ]
+$(666,70,35,4)$ 7 ? [ 0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(676,150,23,36)$ 3 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36 ]
+10 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36 ]
+11 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36 ]
+$(689,256,120,80)$ 1 ? [ 0, 4, 4, 4, 4, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20, 20 ]
+$(737,96,35,9)$ 1 4.9 [ 0, 1, 1, 1, 1, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9 ]
+$(755,130,9,25)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 25, 25, 25, 25 ]
+$(755,174,53,36)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36 ]
+$(776,124,39,16)$ 12 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16 ]
+$(813,252,71,81)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 81, 81 ]
-(813,290,109,100)(813,290,109,100)( 813 , 290 , 109 , 100 ) 1 ?
+$(813,290,109,100)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 100, 100 ]
-(889,222,35,62)(889,222,35,62)( 889 , 222 , 35 , 62 ) 1 ? [0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 31, 31, 31, 31, 31, 31]
-(889,288,112,84)(889,288,112,84)( 889 , 288 , 112 , 84 ) 1 ? [0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 42, 42, 42, 42, 42, 42]
-(903,82,41,4)(903,82,41,4)( 903 , 82 , 41 , 4 ) 1 [27] [0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4]
-(981,140,43,16)(981,140,43,16)( 981 , 140 , 43 , 16 ) 3 4.13 [0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16]
-(981,392,133,172)(981,392,133,172)( 981 , 392 , 133 , 172 ) 3 ? [0, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 43, 43, 43, 43, 43, 43, 43, 43]
-(981,480,254,216)(981,480,254,216)( 981 , 480 , 254 , 216 ) 3 ? [0, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 54, 54, 54, 54, 54, 54, 54, 54]
-(993,310,89,100)(993,310,89,100)( 993 , 310 , 89 , 100 ) 1 ?
+$(820,78,39,4)$ 7 ? [ 0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(842,406,195,196)$ 1 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 196 ]
+$(889,222,35,62)$ 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 31, 31, 31, 31, 31, 31 ]
+$(889,288,112,84)$ 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 42, 42, 42, 42, 42, 42 ]
+$(900,248,79,64)$ 33 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64, 64 ]
+119 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64, 64 ]
+120 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64, 64 ]
+$(903,82,41,4)$ 1 [27] [ 0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(904,378,177,144)$ 12 ? [ 0, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 48, 48, 48, 48, 48, 48, 48 ]
+$(981,140,43,16)$ 3 4.13 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16 ]
+$(981,392,133,172)$ 3 ? [ 0, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 43, 43, 43, 43, 43, 43, 43, 43 ]
+$(981,480,254,216)$ 3 ? [ 0, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 54, 54, 54, 54, 54, 54, 54, 54 ]
+$(993,310,89,100)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 100, 100 ]
-(993,352,131,121)(993,352,131,121)( 993 , 352 , 131 , 121 ) 1 ?
+$(993,352,131,121)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 121, 121 ]
-(1027,114,41,9)(1027,114,41,9)( 1027 , 114 , 41 , 9 ) 1 4.9 [0, 1, 1, 1, 1, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9]
-(1027,342,81,130)(1027,342,81,130)( 1027 , 342 , 81 , 130 ) 1 ? [0, 5, 5, 5, 5, 5, 5, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26]
-(1027,450,225,175)(1027,450,225,175)( 1027 , 450 , 225 , 175 ) 1 ? [0, 5, 5, 5, 5, 5, 5, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35]
-(1055,186,17,36)(1055,186,17,36)( 1055 , 186 , 17 , 36 ) 1 ?
-[0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36]
-(1055,238,69,49)(1055,238,69,49)( 1055 , 238 , 69 , 49 ) 1 ?
-[0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 49, 49, 49, 49]
-(1081,90,45,4)(1081,90,45,4)( 1081 , 90 , 45 , 4 ) 1 [27] [0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4]
-(1081,486,177,252)(1081,486,177,252)( 1081 , 486 , 177 , 252 ) 1 ? [0, 12, 12, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21]
-(1191,374,109,121)(1191,374,109,121)( 1191 , 374 , 109 , 121 ) 1 ?
+$(1027,114,41,9)$ 1 4.9 [ 0, 1, 1, 1, 1, 1, 1, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9, 9 ]
+$(1027,342,81,130)$ 1 ? [ 0, 5, 5, 5, 5, 5, 5, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26, 26 ]
+$(1027,450,225,175)$ 1 ? [ 0, 5, 5, 5, 5, 5, 5, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35, 35 ]
+$(1055,186,17,36)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36 ]
+$(1055,238,69,49)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 49, 49, 49, 49 ]
+$(1081,90,45,4)$ 1 [27] [ 0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+$(1081,486,177,252)$ 1 ? [ 0, 12, 12, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21, 21 ]
+$(1086,210,59,36)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 36, 36, 36, 36, 36 ]
+$(1090,528,255,256)$ 3 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 256 ]
+$(1156,264,47,64)$ 3 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64, 64 ]
+10 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64, 64 ]
+11 ?
+[ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 64, 64, 64 ]
+$(1176,94,47,4)$ 213 ? [ 0, 1, 1, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4, 4 ]
+215 ? [ 0, 1, 1, 4, 24, 16, 16, 16, 16 ]
+218 ? [ 0, 1, 1, 4, 4, 4, 4, 4, 8, 8, 8, 8, 8, 8, 8, 8, 8 ]
+$(1191,374,109,121)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 121, 121 ]
-(1191,420,155,144)(1191,420,155,144)( 1191 , 420 , 155 , 144 ) 1 ?
+$(1191,420,155,144)$ 1 ?
 [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 144, 144 ]
-(1205,448,150,176)(1205,448,150,176)( 1205 , 448 , 150 , 176 ) 1 ?
-[0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 88, 88, 88, 88]
+$(1205,448,150,176)$ 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 88, 88, 88, 88 ]
+$(1205,516,235,210)$ 1 ? [ 0, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 2, 105, 105, 105, 105 ]
+$(1210,156,47,16)$ 1 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16, 16 ]
+9 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16, 16 ]
+10 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16, 16 ]
+11 ? [ 0, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 16, 16, 16, 16, 16, 16, 16, 16, 16 ]